Computing Science Technical report No. 145

A Permuted Index for TEX and L^{A}TEX

Bill Cheswick ches@research.att.com

Last revision: 12 August 1991

Contents

1 Introduction 3 2 Permuted Index 5 3 Primitive and Plain TEX commands 95 4 L^{A}TEX Commands 107 2

Chapter 1

Introduction

This Index is designed to help TEX and L^{A}TEX users find the right command among the one thousand-odd commands documented. It is a permuted, or keyword-in-context index. Each command appears under each significant keyword in its definition. For example, \eject has the following definition:

     force a page break

The following two entries appear in the Index, alphabetized by the keywords “page” and “break”:

force a page  break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \eject^{P}
force a  page break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \eject^{P}

The superscript “P” at the end of the line means this is defined in the plain TEX macros. The full list is:

*  TEX primitive
(none) plain TEX
L  L^{A}TEX command
3  new TEX version 3 command

If a definition is too long to fit on one side of a definition, it is wrapped around. The command definitions appear after the permuted index. There are two sections, one for TEX primitives and plain TEX, and the other for L^{A}TEX commands.

There may be some difficulty thinking of the right keyword used in a definition. I have tried to be consistent. Here are a few words I chose (in bold) and some likely synonyms: create: make

define: create, initialize, set

select, construct: use

space: distance, glue, separation

test: if

The Proceedings of the summer 1989 TEX Users Group meeting contain a description of how this Index was prepared. The Proceedings were published in the TUGboat 10, no. 4.

I would like to thank Barbara Beeton, Michael Downes, John Hobby, Howard Trickey, and many others for definitions and suggestions. Doug McIlroy had some helpful suggestions about editing the Index. I’d also like to thank my wife Lorette, who endured my discovery of how much work an index requires.

Chapter 2

Permuted Index

10 point bold font. . . . . . . . . . . . . . . . . . . . . . . . . .  \tenbf
10 point circle font. . . . . . . . . . . . . . . . . . . . . . . . . \tencirc^{L}
10 point L^{A}TEX symbol font. . . . . . . . . . . . . . . . .  \tenly^{L}
10 point line font. . . . . . . . . . . . . . . . . . . . . . . . . . .  \tenln^{L}
10 point math extension symbol font. . . . . . .  \tenex
10 point math italic font. . . . . . . . . . . . . . . . . . .  \teni
10 point math italic font. . . . . . . . . . . . . . . . . . .  \tenmi^{L}
10 point math symbol font. . . . . . . . . . . . . . . . .  \tensy
10 point Roman font. . . . . . . . . . . . . . . . . . . . . . .  \tenrm
10 point sans serif font. . . . . . . . . . . . . . . . . . . . .  \tensf^{L}
10 point slanted font. . . . . . . . . . . . . . . . . . . . . . .  \tensl
10 point text italic font. . . . . . . . . . . . . . . . . . . . .  \tenit
10 point typewriter font. . . . . . . . . . . . . . . . . . . .  \tentt
10 point wide circle font. . . . . . . . . . . . . . . . . . . \tencircw^{L}
10 point wide line font. . . . . . . . . . . . . . . . . . . . . . \tenlnw^{L}
ratio for variable delimiters times  1000. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \delimiterfactor^{*}
magnification ratio times 1000. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mag^{*}
11 point bold extended font. . . . . . . . . . . . . . . .  \elvbf^{L}
11 point italic font. . . . . . . . . . . . . . . . . . . . . . . . .  \elvit^{L}
11 point L^{A}TEX symbol font. . . . . . . . . . . . . . . . .  \elvly^{L}
use  11 point magnification. . . . . . . . . . . . . . . . . \magstephalf
11 point math italic font. . . . . . . . . . . . . . . . . . .  \elvmi^{L}
11 point math symbol font. . . . . . . . . . . . . . . . .  \elvsy^{L}
11 point Roman font. . . . . . . . . . . . . . . . . . . . . . .  \elvrm^{L}
11 point sans serif font. . . . . . . . . . . . . . . . . . . . .  \elvsf^{L}
11 point slanted font. . . . . . . . . . . . . . . . . . . . . . .  \elvsl^{L}
11 point typewriter font. . . . . . . . . . . . . . . . . . . .  \elvtt^{L}
12 point bold extended font. . . . . . . . . . . . . . . .  \twlbf^{L}
12 point italic font. . . . . . . . . . . . . . . . . . . . . . . . .  \twlit^{L}
12 point L^{A}TEX symbol font. . . . . . . . . . . . . . . . .  \twlly^{L}
12 point math italic font. . . . . . . . . . . . . . . . . . .  \twlmi^{L}
12 point math symbol font. . . . . . . . . . . . . . . . .  \twlsy^{L}
12 point Roman font. . . . . . . . . . . . . . . . . . . . . . .  \twlrm^{L}
12 point sans serif font. . . . . . . . . . . . . . . . . . . . .  \twlsf^{L}
12 point slanted font. . . . . . . . . . . . . . . . . . . . . . .  \twlsl^{L}
12 point typewriter font. . . . . . . . . . . . . . . . . . . .  \twltt^{L}
14 point bold extended font. . . . . . . . . . . . . . . .  \frtnbf^{L}
14 point L^{A}TEX symbol font. . . . . . . . . . . . . . . . . \frtnly^{L}

14 point math italic font. . . . . . . . . . . . . . . . . . .  \frtnmi^{L}
14 point math symbol font. . . . . . . . . . . . . . . . .  \frtnsy^{L}
14 point Roman font. . . . . . . . . . . . . . . . . . . . . . . \frtnrm^{L}
1.5-line left math delimiter. . . . . . . . . . . . . . . . .  \Bigl
1.5-line math delimiter size. . . . . . . . . . . . . . . . .  \Big
1.5-line middle math delimiter. . . . . . . . . . . . . .  \Bigm
1.5-line right math delimiter. . . . . . . . . . . . . . . .  \Bigr
17 point bold extended font. . . . . . . . . . . . . . . .  \svtnbf^{L}
17 point L^{A}TEX symbol font. . . . . . . . . . . . . . . . . \svtnly^{L}
17 point math italic font. . . . . . . . . . . . . . . . . . .  \svtnmi^{L}
17 point math symbol font. . . . . . . . . . . . . . . . .  \svtnsy^{L}
17 point Roman font. . . . . . . . . . . . . . . . . . . . . . . \svtnrm^{L}
1-line left math delimiter. . . . . . . . . . . . . . . . . . .  \bigl
1-line math delimiter size. . . . . . . . . . . . . . . . . . .  \big
1-line middle math delimiter. . . . . . . . . . . . . . . .  \bigm
1-line right math delimiter. . . . . . . . . . . . . . . . . .  \bigr
an unoriented skip amount with stretch of 1fill. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fill^{L}
20 point bold extended font. . . . . . . . . . . . . . . .  \twtybf^{L}
20 point L^{A}TEX symbol font. . . . . . . . . . . . . . . . . \twtyly^{L}
20 point math italic font. . . . . . . . . . . . . . . . . . .  \twtymi^{L}
20 point math symbol font. . . . . . . . . . . . . . . . .  \twtysy^{L}
20 point Roman font. . . . . . . . . . . . . . . . . . . . . . . \twtyrm^{L}
2.5-line left math delimiter. . . . . . . . . . . . . . . . .  \Biggl
2.5-line math delimiter size. . . . . . . . . . . . . . . . .  \Bigg
2.5-line middle math delimiter. . . . . . . . . . . . . .  \Biggm
2.5-line right math delimiter. . . . . . . . . . . . . . . .  \Biggr
2-line left math delimiter. . . . . . . . . . . . . . . . . . .  \biggl
2-line math delimiter size. . . . . . . . . . . . . . . . . . .  \bigg
2-line middle math delimiter. . . . . . . . . . . . . . . .  \biggm
2-line right math delimiter. . . . . . . . . . . . . . . . . .  \biggr
5 point boldface Roman font. . . . . . . . . . . . . . .  \fivebf
5 point L^{A}TEX symbol font. . . . . . . . . . . . . . . . . .  \fivly^{L}
5 point math italic font. . . . . . . . . . . . . . . . . . . . .  \fivei
5 point math italic font. . . . . . . . . . . . . . . . . . . . .  \fivmi^{L}
5 point math symbol font. . . . . . . . . . . . . . . . . . . \fivesy
5 point math symbol font. . . . . . . . . . . . . . . . . . .  \fivsy^{L}
5 point Roman font. . . . . . . . . . . . . . . . . . . . . . . .  \fiverm
5 point Roman font. . . . . . . . . . . . . . . . . . . . . . . .  \fivrm^{L}
6 point L^{A}TEX symbol font. . . . . . . . . . . . . . . . . .  \sixly^{L}
6 point math italic font. . . . . . . . . . . . . . . . . . . . .  \sixmi^{L}
6 point math symbol font. . . . . . . . . . . . . . . . . . .  \sixsy^{L}
6 point Roman font. . . . . . . . . . . . . . . . . . . . . . . .  \sixrm^{L}
7 point bold Roman font. . . . . . . . . . . . . . . . . . . \sevenbf
7 point italic font. . . . . . . . . . . . . . . . . . . . . . . . . . .  \sevit^{L}
7 point L^{A}TEX symbol font. . . . . . . . . . . . . . . . . .  \sevly^{L}
7 point math italic font. . . . . . . . . . . . . . . . . . . . . \seveni
7 point math italic font. . . . . . . . . . . . . . . . . . . . .  \sevmi^{L}
7 point math symbol font. . . . . . . . . . . . . . . . . . . \sevensy
7 point math symbol font. . . . . . . . . . . . . . . . . . .  \sevsy^{L}
7 point Roman font. . . . . . . . . . . . . . . . . . . . . . . . \sevenrm
7 point Roman font. . . . . . . . . . . . . . . . . . . . . . . .  \sevrm^{L}
8 point italic font. . . . . . . . . . . . . . . . . . . . . . . . . . .  \egtit^{L}
8 point L^{A}TEX symbol font. . . . . . . . . . . . . . . . . .  \egtly^{L}

8 point math italic font. . . . . . . . . . . . . . . . . . . . .  \egtmi^{L}
8 point math symbol font. . . . . . . . . . . . . . . . . . .  \egtsy^{L}
8 point Roman font. . . . . . . . . . . . . . . . . . . . . . . .  \egtrm^{L}
9 point bold extended font. . . . . . . . . . . . . . . . .  \ninbf^{L}
9 point italic font. . . . . . . . . . . . . . . . . . . . . . . . . . .  \ninit^{L}
9 point L^{A}TEX symbol font. . . . . . . . . . . . . . . . . .  \ninly^{L}
9 point math italic font. . . . . . . . . . . . . . . . . . . . .  \ninmi^{L}
9 point math symbol font. . . . . . . . . . . . . . . . . . .  \ninsy^{L}
9 point Roman font. . . . . . . . . . . . . . . . . . . . . . . .  \ninrm^{L}
9 point typewriter font. . . . . . . . . . . . . . . . . . . . .  \nintt^{L}
Scandinavian letter:  a with circle (å). . . . . . . . . . . . . . . . . . . . . . . . . . . .  \aa
Scandinavian letter: capital  A with circle (Å). . . . . . . . . . . . . . . . . . . . . . . . . . .  \AA
place math limits  above and below math operators. . . . . . . . . . .  \limits^{*}
stack one equation  above another. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \stackrel^{L}
extra space  above displays. . . . . . . . . . . . . . . . . . . \abovedisplayskip^{*}
lines.  extra space  above displays following short \abovedisplayshortskip^{*}
math  accent: acute (x́). . . . . . . . . . . . . . . . . . . . . . . . . . .  \acute
math  accent: bar (x̄). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bar
math  accent: bar under (x̄). . . . . . . . . . . . . . . . . . . . . . .  \b
math  accent: breve (x̆). . . . . . . . . . . . . . . . . . . . . . . . . . .  \breve
cedilla  accent (ç). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \c
math  accent: check (x̌). . . . . . . . . . . . . . . . . . . . . . . . . . .  \check
check  accent (x̌). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \v
math  accent: dot (ẋ). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \dot
math  accent: double dot (ẍ). . . . . . . . . . . . . . . . . . . . . .  \ddot
dot under  accent (\begin{matrix}x \\ . \end{matrix}). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \d
acute  accent (é). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \'
grave  accent (è). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \`
math  accent: grave (x̀). . . . . . . . . . . . . . . . . . . . . . . . . . .  \grave
math  accent: hat (x̂). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \hat
tilde  accent (ñ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \~
circumflex  accent (ô). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \^
macron or bar  accent (ō). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \=
put an  accent over the next character. . . . . . . . . . . . . . \accent^{*}
place an  accent over the next math field. . . . . . . . . . \mathaccent^{*}
diæresis or umlaut  accent (ẍ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \"
long Hungarian umlaut  accent (ő). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \H
tie-after  accent (oo). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \t
math  accent: tilde (x̃). . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tilde
breve  accent (x̆). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \u
math  accent: vector (x⃗). . . . . . . . . . . . . . . . . . . . . . . . . .  \vec
math  accent: wide hat (x̂). . . . . . . . . . . . . . . . . . . . . . . . \widehat
math  accent: wide tilde (x̃). . . . . . . . . . . . . . . . . . . . . \widetilde
dot accent (ẋ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \.
default horizontal kern character to position  accents. . . . . . . . . . . . . . . . . . . . . . . . . . . . \defaultskewchar^{*}
shift super accents. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \skew
horizontal kern to position  accents. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \skewchar^{*}
produce TEX accents in tabbing environment. . . . . . . . . . . . . \a^{L}
accept looser line and page breaks. . . . . . . . . . \sloppy^{L}
set @ alphabetic to  access internal commands. . . . . . . . . . . . . \makeatletter^{L}
category code for active characters. . . . . . . . . . . . . . . . . . . . . . . . . . . \active
acute accent (é). . . . . . . . . . . . . . . . . . . . . . . . . . . .  \'
math accent:  acute (x́). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \acute

add a box to the vertical list shifted left… \moveleft^{*}
add a box to the vertical list shifted right.  \moveright^{*}
or tabular environment.  add extra space before a column in array \extracolsep^{L}
add extra vertical space. . . . . . . . . . . . . . . . . . \addvspace^{L}
add footnote to title page. . . . . . . . . . . . . . . . . .  \thanks^{L}
add horizontal space. . . . . . . . . . . . . . . . . . . . . . . .  \hskip^{*}
un-box an hbox and  add it to the horizontal list. . . . . . . . . . . . . . . . . \unhbox^{*}
un-box a copy of an hbox and  add it to the horizontal list. . . . . . . . . . . . . . . . . \unhcopy^{*}
un-box a vbox and add it to the vertical list. . . . . . . . . . . . . . . . . . . \unvbox^{*}
un-box a copy of a vbox and  add it to the vertical list. . . . . . . . . . . . . . . . . . . \unvcopy^{*}
tables. add text to table contents, figures, or \addtocontents^{L}
add vertical space. . . . . . . . . . . . . . . . . . . . . . . . . .  \vskip^{*}
paragraph. amount added to badness of every line in a . . . . . \linepenalty^{*}
remove a kern just  added to the current list. . . . . . . . . . . . . . . . . . . . \unkern^{*}
remove a penalty just  added to the current list. . . . . . . . . . . . . . . . . . \unpenalty^{*}
remove a skip just  added to the current list. . . . . . . . . . . . . . . . . . . . \unskip^{*}
extra space  added to top of page. . . . . . . . . . . . . . . . . . . . . \topmargin^{L}
additional authors on title page. . . . . . . . . . . . .  \and^{L}
between lines. additional penalty for page break \interlinepenalty^{*}
table.  adds an entry to the specified list or \addcontentsline^{L}
demerits for  adjacent incompatible lines. . . . . . . . . . . . . \adjdemerits^{*}
advance \pageno by one. . . . . . . . . . . . . \advancepageno
ligature digraph symbol  ae (æ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ae
ligature digraph symbol capital  AE (Æ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \AE
penalty for page break just  after a display. . . . . . . . . . . . . . . . . \postdisplaypenalty^{*}
permit a line break after a slash. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \slash
penalty for line break  after binary operation. . . . . . . . . . . . . . . . . \binoppenalty^{*}
penalty for line break  after discretionary hyphen. . . . . . . . . . . \hyphenpenalty^{*}
pause  after each line is read from a file. . . . . . . . . . . . \pausing^{*}
tokens to insert  after every \cr or nonredundant \crcr. . . . . \everycr^{*}
penalty for line break  after explicit hyphen. . . . . . . . . . . . . . \exhyphenpenalty^{*}
penalty if page break  after first line of paragraph. . . . . . . . . . . . . \clubpenalty^{*}
create end-of-sentence space after following punctuation. . . . . . . . . . . . . . . . . \@^{L}
penalty if page break  after hyphenated line. . . . . . . . . . . . . . . . \brokenpenalty^{*}
badness tolerance  after hyphenation. . . . . . . . . . . . . . . . . . . . . . . . \tolerance^{*}
a word. minimum number of characters after hyphenation at the end of \righthyphenposition^{3}
place superscripts and subscripts  after math operators. . . . . . . . . . . . . . . . . . . . . . . \nolimits^{*}
penalty for line break  after math relation. . . . . . . . . . . . . . . . . . . . . . \relpenalty^{*}
suppress special spacing  after punctuation. . . . . . . . . . . . . . . . . . . . \frenchspacing
enable special spacing  after punctuation. . . . . . . . . . . . . . . . \nonfrenchspacing
hanging indentation changes  after specified number of lines. . . . . . . . . . . . \hangafter^{*}
extra space  after subscript or superscript. . . . . . . . . . . \scriptspace^{*}
insert a token  after the current group is completed. . . . . \aftergroup^{*}
insert a token  after the next assignment command. \afterassignment^{*}
math symbol: aleph (ℵ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \aleph
align a stack of equations. . . . . . . . . . . . . . . . . . . \eqalign
numbers.  align a stack of equations with equation . . \eqalignno
numbers.  align a stack of equations with left equation \leqalignno
align zero or more columns. . . . . . . . . . . . . . . . . \valign^{*}
space between  aligned tab entries. . . . . . . . . . . . . . . . . . . . . . . . . . \tabskip^{*}
create an  aligned table. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \halign^{*}
create a multicolumn entry in an  aligned table. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \span^{*}
end a line in \halign  aligned text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \cr^{*}

equivalent to \cr, end of  aligned text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \endline
display a stack of formulas without  alignment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \displaylines
special Plain TEX space used in  alignment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \hideskip
span several columns in an  alignment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \multispan
permit an  alignment entry to stick out of its column.  \hidewidth
suppress the template in the  alignment preamble for this entry. . . . . . . . . . .  \omit^{*}
allocate a new box register. . . . . . . . . . . . . . . . .  \newbox
allocate a new count register. . . . . . . . . . . . . . . \newcount
allocate a new dimension register. . . . . . . . . . \newdimen
allocate a new input file. . . . . . . . . . . . . . . . . . . . \newread
allocate a new insert register. . . . . . . . . . . . . \newinsert
allocate a new math skip register. . . . . . . . . \newmuskip
allocate a new output file. . . . . . . . . . . . . . . . . . \newwrite
allocate a new skip register. . . . . . . . . . . . . . . . . \newskip
allocate a new token register. . . . . . . . . . . . . . . \newtoks
most recently  allocated register number. . . . . . . . \allocationnumber
allow a line break. . . . . . . . . . . . . . . . . . . . . . . \allowbreak
math Greek letter: alpha (α). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \alpha
set @  alphabetic to access internal commands. \makeatletter^{L}
alternative to an \if. . . . . . . . . . . . . . . . . . . . . . .  \else^{*}
test  always false. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \iffalse^{*}
test  always succeeds. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \iftrue^{*}
math operator:  amalgamated sum, co-product (⨿). . . . . . . . . .  \amalg
vertical skip a large  amount. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigskip^{L}
vertical skip a medium  amount. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \medskip^{L}
vertical skip a small  amount. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \smallskip^{L}
a paragraph.  amount added to badness of every line in \linepenalty^{*}
fail to span included material.  amount by which delimiters can \delimitershortfall^{*}
split off a specified  amount from a vbox. . . . . . . . . . . . . . . . . . . . . . . . \vsplit^{*}
an index. amount of extra space between entries in \indexspace^{L}
amount of fil space in current page.  \pagefilstretch^{*}
amount of fill space in current page. \pagefillstretch^{*}
amount of filll space in current page. \pagefilllstretch^{*}
amount of glue shrinkage in current page. \pageshrink^{*}
amount of glue stretch in current page.  \pagestretch^{*}
generate a short  amount of verbatim text. . . . . . . . . . . . . . . . . . . .  \verb^{L}
an unoriented skip  amount with stretch of 1fill. . . . . . . . . . . . . .  \fill^{L}
ampersand. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \&
math delimiter: left  angle bracket (⟨). . . . . . . . . . . . . . . . . . . . . . . . . . .  \langle
math delimiter: right  angle bracket (⟩). . . . . . . . . . . . . . . . . . . . . . . . . . .  \rangle
angle symbol (^{}). . . . . . . . . . . . . . . . . . . . . . . . . . .  \angle
stack one equation above  another. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \stackrel^{L}
the following macro must not be called from another macro. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \outer^{*}
current list.  append a discretionary item to the . . \discretionary^{*}
width of rules  appended to overfull boxes. . . . . . . . . . . . \overfullrule^{*}
set sectional units to  appendix style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \appendix^{L}
math relation: approximately equal (≈). . . . . . . . . . . . . . . . . . . \approx
display counter as  Arabic numerals. . . . . . . . . . . . . . . . . . . . . . . . . . . . \arabic^{L}
math function:  arc cosine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \arccos
math function: arc sine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \arcsin
math function:  arc tangent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \arctan
math function: arg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \arg
protect fragile commands and moving  arguments. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \protect^{L}

perform  arithmetic on a register. . . . . . . . . . . . . . . . . . . . . \advance^{*}
vertical space  around a float in the middle of a page. … \intextsep^{L}
define kerning  around math in text. . . . . . . . . . . . . . . . . . \mathsurround^{*}
put a frame  around some text. . . . . . . . . . . . . . . . . . . . . . . . . . .  \frame^{L}
horizontal line in  array and tabular environments. . . . . . . . . . . .  \hline^{L}
vertical line in  array and tabular environments. . . . . . . . . . . .  \vline^{L}
column separation in  array environment. . . . . . . . . . . . . . . . . . . . . \arraycolsep^{L}
begin array environment. . . . . . . . . . . . . . . . . . . . . . . . . . \array^{L}
end  array environment. . . . . . . . . . . . . . . . . . . . . . . . . \endarray^{L}
space between rows of  array or tabular environment. . . . . . . . . . \arraystretch^{L}
multicolumn line in array or tabular environment. . . . . . . . . . . . . . . \cline^{L}
double rule separation in  array or tabular environment. . . . . . . . \doublerulesep^{L}
add extra space before a column in  array or tabular environment. . . . . . . . . . . \extracolsep^{L}
multicolumn entry in  array or tabular environment. . . . . . . . . . . \multicolumn^{L}
set width of  array rules. . . . . . . . . . . . . . . . . . . . . . . . . \arrayrulewidth^{L}
fill a space with a left  arrow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leftarrowfill
fill a space with a right  arrow. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \rightarrowfill
math symbol: down  arrow (↓). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \downarrow
math symbol: downward double  arrow (⇓). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \Downarrow
math symbol: hook left  arrow (↩). . . . . . . . . . . . . . . . . . . . . . . . . . . \hookleftarrow
math symbol: hook right  arrow (↪). . . . . . . . . . . . . . . . . . . . . . . . . \hookrightarrow
math symbol: left  arrow (←). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leftarrow
math symbol: left double  arrow (⇐). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \Leftarrow
math symbol: left-right  arrow (↔). . . . . . . . . . . . . . . . . . . . . . . . . . \leftrightarrow
math symbol: left-right double  arrow (⇔). . . . . . . . . . . . . . . . . . . . . . . . . . \Leftrightarrow
math symbol: long left  arrow (⟵). . . . . . . . . . . . . . . . . . . . . . . . . . \longleftarrow
math symbol: long left double  arrow (⟸). . . . . . . . . . . . . . . . . . . . . . . . . . \Longleftarrow
math symbol: long left and right  arrow (⟷). . . . . . . . . . . . . . . . . . . \longleftrightarrow
math symbol: long left-right double  arrow (⟺). . . . . . . . . . . . . . . . . . . \Longleftrightarrow
math symbol: long right  arrow (⟶). . . . . . . . . . . . . . . . . . . . . . . . \longrightarrow
math symbol: long right double  arrow (⟹). . . . . . . . . . . . . . . . . . . . . . . . \Longrightarrow
math symbol: northeast  arrow (↗). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \nearrow
math symbol: northwest  arrow (↖). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \nwarrow
left  arrow over a math formula. . . . . . . . . . . \overleftarrow
right  arrow over a math formula. . . . . . . . . \overrightarrow
math symbol: right  arrow (→). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \rightarrow
math symbol: right double  arrow (⇒). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \Rightarrow
math symbol: southeast  arrow (↘). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \searrow
math symbol: southwest  arrow (↙). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \swarrow
math delimiter: upward  arrow (↑). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \uparrow
math symbol: upward double  arrow (⇑). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \Uparrow
math delimiter: up-and-down  arrow (↕). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \updownarrow
math symbol: up-and-down double  arrow (⇕). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \Updownarrow
Plain TEX command to piece together long arrows. internal . . . . . . . . . . . . . . . . . . . . . . . . \joinrel
help message to display if user  asks for help. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \errhelp^{*}
insert a token after the next  assignment command. . . . . . . . . . . . . . \afterassignment^{*}
sequence and continues.  assigns the second token to a control . . . . . \futurelet^{*}
command.  associate a counter with an item-type … \usecounter^{L}
math operator:  asterisk (∗). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ast
math relation:  asymptote (≍). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \asymp
badness tolerance before hyphenation is  attempted. . . . . . . . . . . . . . . . . . . . . . . . . . . . \pretolerance^{*}
title page  author. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \author^{L}
additional  authors on title page. . . . . . . . . . . . . . . . . . . . . . .  \and^{L}

suppress writing all  auxiliary files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \nofiles^{L}
select largest available font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \Huge^{L}
math delimiter:  backslash (\). . . . . . . . . . . . . . . . . . . . . . . . . . . . . \backslash
extra space in  badly-stretched lines. . . . . . . . . . . . . \emergencystretch^{3}
badness of a box. . . . . . . . . . . . . . . . . . . . . . . . . . . \badness^{3}
amount added to  badness of every line in a paragraph. … \linepenalty^{*}
badness tolerance after hyphenation. . . . . . \tolerance^{*}
attempted.  badness tolerance before hyphenation is \pretolerance^{*}
write  balanced error message to the terminal. . \errmessage^{*}
include  balanced text in DVI file for post-processing. \special^{*}
list. place balanced text into a mark item on the current \mark^{*}
translate  balanced text to lower-case. . . . . . . . . . . . . . . \lowercase^{*}
write  balanced text to terminal. . . . . . . . . . . . . . . . . . . \message^{*}
translate  balanced text to upper-case. . . . . . . . . . . . . . . \uppercase^{*}
macron or  bar accent (ō). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \=
math accent:  bar (x̄). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bar
math symbol: double vertical  bar (‖). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \Arrowvert
math symbol: vertical  bar (|). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \arrowvert
math symbol: double relation  bar (=). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Relbar
math symbol: relation  bar (−). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \relbar
math symbol: vertical  bar (∥). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \|
math accent:  bar under (x̄). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \b
math delimiter: double vertical  bar (∥). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Vert
math delimiter: vertical  bar (|). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vert
create a box with vertical mode material with the  baseline at the top. . . . . . . . . . . . . . . . . . . . . . . . .  \vtop^{*}
multiple of normal  \baselineskip. . . . . . . . . . . . . . . . . . . \baselinestretch^{L}

normal  \baselineskip. . . . . . . . . . . . . . . . \normalbaselineskip

extra space between lines if  \baselineskip isn’t enough. . . . . . . . . . . . . . . \lineskip^{*}
begin a floating figure. . . . . . . . . . . . . . . . . . . . . . \figure^{L}
begin a group. . . . . . . . . . . . . . . . . . . . . . . . . . . \begingroup^{*}
begin a group. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bgroup
begin a loop. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \loop
begin a new math list with a left delimiter. .  \left^{*}
begin a tabbed line in an inner environment. \tabalign
begin a tabbed line in an outer environment.  \+
begin and cross reference an equation. . . . . . \equation^{L}
begin array environment. . . . . . . . . . . . . . . . . . . .  \array^{L}
begin centering environment. . . . . . . . . . . . . . . . \center^{L}
begin comment. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  %
begin display math mode. . . . . . . . . . . . . . . . . . .  \[^{L}
command name.  begin expanding tokens to construct a . . . . . . \csname^{*}
begin flush left environment. . . . . . . . . . . . . . \flushleft^{L}
begin flush right environment. . . . . . . . . . . . \flushright^{L}
begin math mode. . . . . . . . . . . . . . . . . . . . . . . . . . .  \(^{L}
begin picture environment. . . . . . . . . . . . . . . . . . \picture^{L}
environment. begin theorem with special format in math \proclaim
penalty at the  beginning of a list. . . . . . . . . . . . . . . . \@beginparpenalty^{L}
mark the  beginning of a major subdivision. . . . . . \beginsection
beginning of an environment. . . . . . . . . . . . . . . .  \begin^{L}
positive if hyphenating words  beginning with capital letters. . . . . . . . . . . . . . . \uchyph^{*}
tokens to insert when display math  begins. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \everydisplay^{*}
tokens to insert when an hbox  begins. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \everyhbox^{*}

tokens to insert when the job  begins. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \everyjob^{*}
tokens to insert when math in text  begins. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \everymath^{*}
tokens to insert when a paragraph  begins. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \everypar^{*}
tokens to insert when a vbox  begins. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \everyvbox^{*}
break a page unless there is a better \filbreak  below. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \filbreak
extra space just  below displays. . . . . . . . . . . . . . . . . . . \belowdisplayskip^{*}
lines.  extra space just  below displays following short \belowdisplayshortskip^{*}
place math limits above and  below math operators. . . . . . . . . . . . . . . . . . . . . .  \limits^{*}
math Greek letter:  beta (β). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \beta
\framebox.  space  between a box and its contents in \fbox and  \fboxsep^{L}
space  between a label and text of a list item. . . . . \labelsep^{L}
space  between aligned tab entries. . . . . . . . . . . . . . . . . \tabskip^{*}
space  between columns in double column text… \columnsep^{L}
space  between double-column floats. . . . . . . . . . \dblfloatsep^{L}
rule  between double-column floats and text. . \dblfigrule^{L}
text. space between double-column floats and . \dbltextfloatsep^{L}
amount of extra space  between entries in an index. . . . . . . . . . . . . . \indexspace^{L}
paragraph. space between first list item and preceding . . . . . . . \topsep^{L}
space  between floats. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \floatsep^{L}
space  between floats and the text. . . . . . . . . . . \textfloatsep^{L}
normal space  between lines. . . . . . . . . . . . . . . . . . . . . . . . . \baselineskip^{*}
additional penalty for page break  between lines. . . . . . . . . . . . . . . . . . . . . \interlinepenalty^{*}
turn off extra space  between lines. . . . . . . . . . . . . . . . . . . . . \offinterlineskip
extra space  between lines if \baselineskip isn’t enough. \lineskip^{*}
minimum space  between lines in a vertical list. . . . . . . . \lineskiplimit^{*}
penalty for breaking  between lines of a display. \interdisplaylinepenalty
footnote. penalty for breaking between lines of a . . . . . . \interfootnotelinepenalty
penalty  between list items. . . . . . . . . . . . . . . . . . . . . \@itempenalty^{L}
space  between main text and footnotes. . . . . . . . . . . \footins^{L}
distance  between marginal note and text. . . . . . . \marginparsep^{L}
vertical space  between marginal notes. . . . . . . . . . . . . . \marginparpush^{L}
space  between page header and text. . . . . . . . . . . . . . \headsep^{L}
define space  between paragraphs. . . . . . . . . . . . . . . . . . . . . . . . \parskip^{*}
environment.  space  between paragraphs within an item in a list  \parsep^{L}
environment.  space  between rows of array or tabular . . . . . . \arraystretch^{L}
space  between sentences. . . . . . . . . . . . . . . . . . . . . . . \xspaceskip^{*}
space  between successive items in a list environment. \itemsep^{L}
distance left by \' command  between tabbing fields. . . . . . . . . . . . . . . . . . . \tabbingsep^{L}
place rule  between text and bottom floats. . . . . . . . . . \botfigrule^{L}
space  between text and page footer. . . . . . . . . . . . . . \footskip^{L}
place rule  between top floats and text. . . . . . . . . . . . . . \topfigrule^{L}
space  between words. . . . . . . . . . . . . . . . . . . . . . . . . . . . \spaceskip^{*}
specify  bibliographic style for BibTEX… \bibliographystyle^{L}
create a  bibliography entry. . . . . . . . . . . . . . . . . . . . . . . . . . \bibitem^{L}
end  bibliography environment. . . . . . \endthebibliography^{L}
enable BibTEX  bibliography interface. . . . . . . . . . . . . . . . . \bibliography^{L}
specify bibliographic style for  BibTEX. . . . . . . . . . . . . . . . . . . . . . . . . \bibliographystyle^{L}
enable  BibTEX bibliography interface. . . . . . . . . \bibliography^{L}
reference a  BibTEX item without citation. . . . . . . . . . . . . . . \nocite^{L}
big font size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \big^{L}
bigger font size. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Big^{L}
biggest font size. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \BIG^{L}
\bigskip space. . . . . . . . . . . . . . . . . . . . . . \bigskipamount

define a  binary math operator. . . . . . . . . . . . . . . . . . . . . . \mathbin^{*}
math operator: binary modulo (mod). . . . . . . . . . . . . . . . . . . . . . \bmod
penalty for line break after  binary operation. . . . . . . . . . . . . . . . . . . . . . \binoppenalty^{*}
math  binary operator: join or V (∨). . . . . . . . . . . . . .  \vee
math  binary operator: meet or wedge (∧). . . . . . . .  \wedge
math  binary operator: wreath product (≀). . . . . . . .  \wr
a  blank space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \space
end of a \loop  body. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \repeat
produce a vertical list of the  body of a page. . . . . . . . . . . . . . . . . . . . . . . . \pagecontents
select  bold extended font. . . . . . . . . . . . . . . . . . . . . . . . .  \bf
11 point  bold extended font. . . . . . . . . . . . . . . . . . . . . . . . .  \elvbf^{L}
14 point  bold extended font. . . . . . . . . . . . . . . . . . . . . . . . .  \frtnbf^{L}
9 point  bold extended font. . . . . . . . . . . . . . . . . . . . . . . . .  \ninbf^{L}
17 point  bold extended font. . . . . . . . . . . . . . . . . . . . . . . . .  \svtnbf^{L}
12 point  bold extended font. . . . . . . . . . . . . . . . . . . . . . . . .  \twlbf^{L}
20 point  bold extended font. . . . . . . . . . . . . . . . . . . . . . . . .  \twtybf^{L}
10 point  bold font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tenbf
bold font family. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bffam
bold font style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bf^{L}
select  bold math italic and symbol fonts. . . . . . . . . . \boldmath^{L}
unselect  bold math italic and symbol fonts. . . . . . . \unboldmath^{L}
7 point  bold Roman font. . . . . . . . . . . . . . . . . . . . . . . . . . . \sevenbf
5 point  boldface Roman font. . . . . . . . . . . . . . . . . . . . . . . \fivebf
value of  \botmark just before current page was boxed. \topmark^{*}
math symbol: bottom (⊥). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bot
place rule between text and  bottom floats. . . . . . . . . . . . . . . . . . . . . . . . . . . . \botfigrule^{L}
fraction of column for  bottom floats. . . . . . . . . . . . . . . . . . . . . . . \bottomfraction^{L}
suppress  bottom justification of page text. . . . . . \raggedbottom
justify page  bottoms to the same height. . . . . . . . . . . . \flushbottom^{L}
justify page  bottoms to the same height. . . . . . . . . . . \normalbottom
justify page  bottoms to their natural height. . . . . . . \raggedbottom^{L}
upper  bound on output routine calls. . . . . . . . \maxdeadcycles^{*}
math relation:  bowtie (▷◁). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bowtie
badness of a  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \badness^{3}
use a  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \box^{*}
use a copy of a  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \copy^{*}
depth of a  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \dp^{*}
create and frame a  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \fbox^{L}
thickness of rule surrounding framed  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fboxrule^{L}
height of a  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ht^{*}
test for a horizontal  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ifhbox^{*}
test for a vertical  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ifvbox^{*}
test for an empty  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ifvoid^{*}
create a  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \mbox^{L}
suppress interline space before next vertical  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \nointerlineskip
put a paragraph in a  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \parbox^{L}
create and name a  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sbox^{L}
show a  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \showbox^{*}
mark text last encountered in a split  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \splitbotmark^{*}
mark text first encountered in a split  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \splitfirstmark^{*}
maximum depth of boxes in a split  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \splitmaxdepth^{*}
space at top of a split  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \splittopskip^{*}
use a named  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \usebox^{L}

width of a  box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \wd^{*}
raise a  box a distance. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \raisebox^{L}
lower a  box a given distance. . . . . . . . . . . . . . . . . . . . . . . .  \lower^{*}
raise a  box a given distance. . . . . . . . . . . . . . . . . . . . . . . .  \raise^{*}
space between a  box and its contents in \fbox and \framebox. \fboxsep^{L}
math symbol: box (□). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \Box^{L}
box containing a strut. . . . . . . . . . . . . . . . . . . . . \strutbox
box containing current tabs. . . . . . . . . . . . . . . .  \tabs
maximum  box depth shown. . . . . . . . . . . . . . . . . . . . . . \showboxdepth^{*}
create a  box in dashes, with positioning. . . . . . . . . . . . . \dashbox^{L}
create a box of current line width. . . . . . . . . . . . . . . . . . . \line
put page contents in a  box of the proper height. . . . . . . . . . . . . . . . . . . \pagebody
fetch last  box off the current list. . . . . . . . . . . . . . . . . . . . . . \lastbox^{*}
depth of the last  box on the vertical list. . . . . . . . . . . . . . . . . . . . \prevdepth^{*}
surround a space with a repeated  box or rule. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \cleaders^{*}
fill a space with a repeated  box or rule. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leaders^{*}
fill a space with an evenly distributed  box or rule. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \xleaders^{*}
allocate a new  box register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newbox
define a new  box register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newsavebox^{L}
store an hbox or vbox in a  box register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \setbox^{*}
send a  box to the DVI file. . . . . . . . . . . . . . . . . . . . . . . . . \shipout^{*}
add a  box to the vertical list shifted left. . . . . . . . . . \moveleft^{*}
add a  box to the vertical list shifted right. . . . . . . \moveright^{*}
create a  box with a single column of items. . . . . . . \shortstack^{L}
create a  box with an underline. . . . . . . . . . . . . . . . . . . . . \underbar
create a box with horizontal mode material. . . . . . . . . \hbox^{*}
create and frame a  box, with positioning. . . . . . . . . . . . . . . . . . . . . . \framebox^{L}
create a  box, with positioning. . . . . . . . . . . . . . . . . . . . . . . \makebox^{L}
create and name a  box, with positioning. . . . . . . . . . . . . . . . . . . . . . . \savebox^{L}
create an underlined  box with text in it. . . . . . . . . . . . . . . . . . . . . . . . \underline^{L}
create a zero-width  box with text to the left. . . . . . . . . . . . . . . . . . . .  \llap
create a zero-width  box with text to the right. . . . . . . . . . . . . . . . . .  \rlap
constructs a  box with the page foot. . . . . . . . . . . . . . . . \makefootline
constructs a  box with the page header. . . . . . . . . . . . . \makeheadline
create a  box with vertical mode material. . . . . . . . . . . .  \vbox^{*}
baseline at the top.  create a  box with vertical mode material with the …  \vtop^{*}
mark text last encountered on page just  boxed. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \botmark^{*}
mark text first encountered on page just  boxed. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \firstmark^{*}
value of \botmark just before current page was  boxed. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \topmark^{*}
maximum  boxed items shown at a given depth. \showboxbreadth^{*}
width of rules appended to overfull  boxes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \overfullrule^{*}
maximum depth of  boxes in a split box. . . . . . . . . . . . . . . . . . \splitmaxdepth^{*}
maximum depth of  boxes on explicit pages. . . . . . . . . . . . . . . . . \boxmaxdepth^{*}
maximum depth of  boxes on main pages. . . . . . . . . . . . . . . . . . . . . . . \maxdepth^{*}
show  boxes that are shipped out. . . . . . . . . . . \tracingoutput^{*}
lower left piece of a horizontal  brace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \braceld
upper left piece of a horizontal  brace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bracelu
lower right piece of a horizontal  brace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bracerd
upper right piece of a horizontal  brace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \braceru
fill a space with a downward  brace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \downbracefill
fill a space with an upward  brace. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \upbracefill
piece of a vertical  brace (|). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bracevert

generate a matrix with a left  brace delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \cases
math delimiter: left curly  brace ({). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lbrace
horizontal  brace over a math formula. . . . . . . . . . . . . . . . \overbrace
math delimiter: right curly  brace (}). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \rbrace
close  brace symbol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \}
open  brace symbol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \{
horizontal  brace under a math formula. . . . . . . . . . . . . \underbrace
fraction without a rule with  braces. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \brace
math delimiter: left angle  bracket (⟨). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \langle
math delimiter: left  bracket ([). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lbrack
math delimiter: left ceiling bracket (⌈). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lceil
math delimiter: left floor  bracket (⌊). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lfloor
math delimiter: right angle bracket (⟩). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \rangle
math delimiter: right  bracket (]). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \rbrack
math delimiter: right ceiling  bracket (⌉). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \rceil
fraction without a rule with  brackets. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \brack
allow a line  break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \allowbreak
large vertical space or a good page  break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigbreak
large vertical  break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigskip
force a line  break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \break
force a page  break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \eject
good page  break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \goodbreak
encourage a line  break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \linebreak^{L}
medium vertical space or a good page  break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \medbreak
prohibit a line or page  break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \nobreak
suppress a line  break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \nolinebreak^{L}
suppress a page  break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \nopagebreak^{L}
penalty at the current page  break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \outputpenalty^{*}
encourage a page  break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pagebreak^{L}
specify penalty for a line or page  break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \penalty^{*}
small vertical space and a good page  break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \smallbreak
\filbreak below.  break a page unless there is a better . . . . . . . \filbreak
permit a line break after a slash. . . . . . . . . . . . . . . . . . . . . . . . . . \slash
penalty for line  break after binary operation. . . . . . . . . . \binoppenalty^{*}
penalty for line  break after discretionary hyphen. . . . . \hyphenpenalty^{*}
penalty for line  break after explicit hyphen. . . . . . . . \exhyphenpenalty^{*}
penalty if page  break after first line of paragraph. . . . . . \clubpenalty^{*}
penalty if page  break after hyphenated line. . . . . . . . . . \brokenpenalty^{*}
penalty for line  break after math relation. . . . . . . . . . . . . . . . \relpenalty^{*}
additional penalty for page  break between lines. . . . . . . . . . . . . . \interlinepenalty^{*}
force a line break in a paragraph. . . . . . . . . . . . . . . . . . . . . . . \\^{L}
force a line  break in a paragraph. . . . . . . . . . . . . . . . . . . . . . . \newline^{L}
penalty for page  break just after a display. . . . . . . \postdisplaypenalty^{*}
penalty for page  break just before a display. . . . . . \predisplaypenalty^{*}
inhibit a page  break over a region. . . . . . . . . . . . . . . . . . . . . . . . \samepage^{L}
large breakable horizontal skip. . . . . . . . . . . . . . . . . . . \qquad
breakable medium horizontal skip. . . . . . . . . .  \quad
breakable small horizontal skip. . . . . . . . . . . . . \enskip
display.  penalty for  breaking between lines of a \interdisplaylinepenalty
a footnote.  penalty for  breaking between lines of \interfootnotelinepenalty
require strict line and page  breaks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \fussy^{L}
accept looser line and page  breaks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \sloppy^{L}
verbatim line  breaks in input text. . . . . . . . . . . . . . . . . . . . . . \obeylines

breve accent (x̆). . . . . . . . . . . . . . . . . . . . . . . . . . . .  \u
math accent: breve (x̆). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \breve
British pound symbol (£). . . . . . . . . . . . . . . . . .  \pounds^{L}
demerits for a penultimate  broken line. . . . . . . . . . . . . . . . . . . \finalhyphendemerits^{*}
demerits for consecutive  broken lines. . . . . . . . . . . . . . . . . \doublehyphendemerits^{*}
math operator:  bullet (•). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bullet
show page-break  calculations. . . . . . . . . . . . . . . . . . . . . . . . . . . \tracingpages^{*}
show line-break  calculations. . . . . . . . . . . . . . . . . . . . . \tracingparagraphs^{*}
math mode calligraphic letters font. . . . . . . . . . . . . . . . . . . . . \cal
upper bound on output routine  calls. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \maxdeadcycles^{*}
number of output routine  calls since last \shipout. . . . . . . . . . . . . . . . \deadcycles^{*}
amount by which delimiters  can fail to span included material. \delimitershortfall^{*}
environment. cancel effect of one \+ command in tabbing \<^{L}
cancel infinitely stretchable horizontal space.  \hfilneg^{*}
cancel infinitely stretchable vertical space… \vfilneg^{*}
math operator: large  cap (⋂). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigcap
math operator: intersection or  cap (∩). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \cap
math operator: square  cap (⊓). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sqcap
Scandinavian letter:  capital A with circle (Å). . . . . . . . . . . . . . . . . . .  \AA
ligature digraph symbol  capital AE (Æ). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \AE
math Greek letter:  capital delta (Δ). . . . . . . . . . . . . . . . . . . . . . . . . . .  \Delta
math Greek letter:  capital gamma (Γ). . . . . . . . . . . . . . . . . . . . . . . . .  \Gamma
math Greek letter:  capital lambda (Λ). . . . . . . . . . . . . . . . . . . . . . . . . \Lambda
positive if hyphenating words beginning with  capital letters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \uchyph^{*}
Norwegian letter:  capital O with slash (Ø). . . . . . . . . . . . . . . . . . . .  \O
ligature digraph symbol  capital OE (Œ). . . . . . . . . . . . . . . . . . . . . . . . . . . .  \OE
math Greek letter:  capital omega (Ω). . . . . . . . . . . . . . . . . . . . . . . . . .  \Omega
math Greek letter:  capital phi (Φ). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Phi
math Greek letter:  capital pi (Π). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Pi
math Greek letter:  capital psi (Ψ). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Psi
math Greek letter:  capital sigma (Σ). . . . . . . . . . . . . . . . . . . . . . . . . . .  \Sigma
math Greek letter:  capital theta (Θ). . . . . . . . . . . . . . . . . . . . . . . . . . .  \Theta
math Greek letter:  capital upsilon (Υ). . . . . . . . . . . . . . . . . . . . . . . . . \Upsilon
math Greek letter:  capital xi (Ξ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Xi
select small  caps font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sc^{L}
produce a numbered  caption. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \caption^{L}
place a footnote in a  caption or other vertical list. . . . . . . . . . . . . . \vfootnote
defines a  carriage return as \\. . . . . . . . . . . . . . . . . . . . . . . \obeycr^{L}
restores a  carriage return to its usual meaning. . . . . . \restorecr^{L}
separate cases in an \ifcase. . . . . . . . . . . . . . . . . . . . . . . . \or^{*}
category code for active characters. . . . . . . . .  \active
test category codes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ifcat^{*}
set a character’s processing  category type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \catcode^{*}
set a math character’s processing  category type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathcode^{*}
cedilla accent (ç). . . . . . . . . . . . . . . . . . . . . . . . . . .  \c
math delimiter: left  ceiling bracket (⌈). . . . . . . . . . . . . . . . . . . . . . . . . .  \lceil
math delimiter: right  ceiling bracket (⌉). . . . . . . . . . . . . . . . . . . . . . . . . .  \rceil
declaration to  center lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \centering^{L}
center text in a line. . . . . . . . . . . . . . . . . . . . . \centerline
math operator:  centered dot (·). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \cdot
mode.  centered dot with special spacing in math . .  \cdotp
three  centered dots in math mode (⋯). . . . . . . . . . .  \cdots
create a  centered vbox in a math list. . . . . . . . . . . . . . . . \vcenter^{*}

internal Plain TEX space used for  centering. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \centering
begin  centering environment. . . . . . . . . . . . . . . . . . . . . . \center^{L}
end  centering environment. . . . . . . . . . . . . . . . . . . . \endcenter^{L}
hanging indentation  changes after specified number of lines. … \hangafter^{*}
start a  chapter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \chapter^{L}
math symbol: maps to  char ( ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mapstochar
put an accent over the next  character. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \accent^{*}
word delimiter  character. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \boundarychar^{3}
define a name for a  character. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \chardef^{*}
test the next  character. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \@ifnextchar^{L}
lower-case code for a  character. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lccode^{*}
define a name for a math  character. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathchardef^{*}
suppress expansion of a  character. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \noexpand^{*}
space character. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . ␣^{*}
upper-case code for a  character. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \uccode^{*}
define a  character as a delimiter. . . . . . . . . . . . . . . . . . . . . \delcode^{*}
specify a character by its numeric code. . . . . . . . . . . . . . . \char^{*}
specify a math  character code. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathchar^{*}
characters.  change the  character code for a set of special . . . . . . . \dospecials
compare two  character codes. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \if^{*}
hyphenation  character for this font. . . . . . . . . . . . . . . . . . . \hyphenchar^{*}
tokens. escape character in the output of control sequence \escapechar^{*}
input line.  character placed at the right end of an . \endlinechar^{*}
a write statement.  character that starts a new output line in \newlinechar^{*}
default horizontal kern  character to position accents. . . . . . \defaultskewchar^{*}
expand a control sequence into  character tokens. . . . . . . . . . . . . . . . . . . . . . . . . . . . \string^{*}
category code for active  characters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \active
change the character code for a set of special  characters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \dospecials
a font with no  characters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \nullfont^{*}
the end of a word.  minimum number of  characters after hyphenation at \righthyphenposition^{3}
the start of a word.  minimum number of  characters before hyphenation at \lefthyphenposition^{3}
show  characters not in the font. . . . . . . . \tracinglostchars^{*}
set a  character’s processing category type. . . . . . . . \catcode^{*}
set a math  character’s processing category type. . . . . . . \mathcode^{*}
set a  character’s space factor. . . . . . . . . . . . . . . . . . . . . \sfcode^{*}
set a  character’s spacing factor. . . . . . . . . . . . . . \spacefactor^{*}
check accent (x̌). . . . . . . . . . . . . . . . . . . . . . . . . . . .  \v
math accent:  check (x̌). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \check
math Greek letter:  chi (χ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \chi
Scandinavian letter: a with  circle (å). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \aa
Scandinavian letter: capital A with  circle (Å). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \AA
math operator: large  circle (◯). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigcirc
math operator:  circle (∘). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \circ
math operator:  circle dot (⊙). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \odot
10 point  circle font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tencirc^{L}
10 point wide  circle font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tencircw^{L}
circle in a picture environment. . . . . . . . . . . . .  \circle^{L}
math operator:  circle minus (⊖). . . . . . . . . . . . . . . . . . . . . . . . . . . . \ominus
math operator: direct sum, circle plus (⊕). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \oplus
math operator:  circle slash (⊘). . . . . . . . . . . . . . . . . . . . . . . . . . . . . \oslash
math operator: tensor product, circle times (⊗). . . . . . . . . . . . . . . . . . . . . . . . . . . . \otimes
math operator: large  circle with dot (⨀). . . . . . . . . . . . . . . . . . . . . . . . . \bigodot
math operator: large  circle with plus (⨁). . . . . . . . . . . . . . . . . . . . . . . \bigoplus

math operator: large  circle with times (⨂). . . . . . . . . . . . . . . . . . . . . \bigotimes
thick lines for lines and  circles. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \thicklines^{L}
thin lines for lines and  circles. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \thinlines^{L}
circumflex accent (ô). . . . . . . . . . . . . . . . . . . . . . .  \^
reference a BibTEX item without  citation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \nocite^{L}
generate an in-text  citation of a reference. . . . . . . . . . . . . . . . . . . . . .  \cite^{L}
insertion  class for footnote inserts. . . . . . . . . . . . . . . . . . . . \footins
insertion  class for inserts at the top of a page. . . . . . . . \topins
end of an \if clause. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fi^{*}
select math spacing of a  closing delimiter for the next item. . . . . . . . \mathclose^{*}
math symbol:  club suit (♣). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \clubsuit
specify a character by its numeric code. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \char^{*}
specify a math character  code. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathchar^{*}
lower-case  code for a character. . . . . . . . . . . . . . . . . . . . . . . . \lccode^{*}
upper-case  code for a character. . . . . . . . . . . . . . . . . . . . . . . . \uccode^{*}
change the character  code for a set of special characters. . . . . . . \dospecials
category  code for active characters. . . . . . . . . . . . . . . . . . . \active
compare two character  codes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \if^{*}
test category  codes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ifcat^{*}
colon in a math formula. . . . . . . . . . . . . . . . . . . .  \colon
width of a  column. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \columnwidth^{L}
permit an alignment entry to stick out of its  column. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \hidewidth
fraction of  column for bottom floats. . . . . . . . . . . \bottomfraction^{L}
fraction of  column for top floats. . . . . . . . . . . . . . . . . . . \topfraction^{L}
add extra space before a  column in array or tabular environment. \extracolsep^{L}
put text flush right in a column in tabbing environment. . . . . . . . . . . . . \'^{L}
create a box with a single  column of items. . . . . . . . . . . . . . . . . . . . . . . . . \shortstack^{L}
column separation in array environment. \arraycolsep^{L}
space between columns in double  column text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \columnsep^{L}
size of  column that must contain text. . . . . . . . \textfraction^{L}
generate a matrix labeled on rows and  columns. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bordermatrix
width of rule separating double  columns. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \columnseprule^{L}
align zero or more  columns. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \valign^{*}
half the width separating  columns in a tabular environment. . . . . . . . \tabcolsep^{L}
span several  columns in an alignment. . . . . . . . . . . . . . . . . . \multispan
space between  columns in double column text. . . . . . . . . . . \columnsep^{L}
increment the value of a length  command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \addtolength^{L}
insert a token after the next assignment  command. . . . . . . . . . . . . . . . . . . . . . . . . \afterassignment^{*}
define a new  command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newcommand^{L}
define a new length  command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newlength^{L}
includes the section number in a \contentsline  command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \numberline^{L}
redefine a  command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \renewcommand^{L}
set a length  command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \setlength^{L}
define a new \if command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newif
associate a counter with an item-type  command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \usecounter^{L}
distance left by \'  command between tabbing fields. . . . . . . . \tabbingsep^{L}
execute a  command from the terminal. . . . . . . . . . . . . . . . \typein^{L}
cancel effect of one \+  command in tabbing environment. . . . . . . . . .  \<^{L}
begin expanding tokens to construct a  command name. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \csname^{*}
internal Plain TEX  command to piece together long arrows. . . . . \joinrel
set a length  command to width of text. . . . . . . . . . . . . . . \settowidth^{L}
enable only specific \include  commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \includeonly^{L}
set @ alphabetic to access internal  commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . \makeatletter^{L}

set @ non-alphabetic to hide internal  commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \makeatother^{L}
turn on all debugging  commands. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tracingall
protect fragile  commands and moving arguments. . . . . . . . . . \protect^{L}
show  commands before they are executed. \tracingcommands^{*}
begin comment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . %
compare tokens. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ifx^{*}
compare two character codes. . . . . . . . . . . . . . .  \if^{*}
insert a token after the current group is  completed. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \aftergroup^{*}
math relation: congruent (≅). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \cong
demerits for  consecutive broken lines. . . . . \doublehyphendemerits^{*}
math symbol: Planck’s  constant or h-bar (h̄). . . . . . . . . . . . . . . . . . . . . . .  \hbar
begin expanding tokens to  construct a command name. . . . . . . . . . . . . . . .  \csname^{*}
constructs a box with the page foot… \makefootline
constructs a box with the page header. \makeheadline
size of column that must  contain text. . . . . . . . . . . . . . . . . . . . . . . . . . . \textfraction^{L}
box  containing a strut. . . . . . . . . . . . . . . . . . . . . . . . . . \strutbox
box  containing current tabs. . . . . . . . . . . . . . . . . . . . .  \tabs
math relation:  contains (∋). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ni
generate a table of  contents. . . . . . . . . . . . . . . . . . . . . . . . . . . \tableofcontents^{L}
add text to table  contents, figures, or tables. . . . . . . . . . . \addtocontents^{L}
put page  contents in a box of the proper height. . . . . . \pagebody
space between a box and its  contents in \fbox and \framebox. . . . . . . . . . \fboxsep^{L}
display the  contents of a register. . . . . . . . . . . . . . . . . . . . . . . \showthe^{*}
includes the section number in a  \contentsline command. . . . . . . . . . . . . . . \numberline^{L}
message.  number of lines of  context to be displayed in an error \errorcontextlines^{3}
the second token to a control sequence and  continues.  assigns . . . . . . . . . . . . . . . . . . . . \futurelet^{*}
math operator:  contour integral (∮). . . . . . . . . . . . . . . . . . . . . . . .  \oint
define a synonym for the current meaning of a  control sequence. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \let^{*}
an undefined  control sequence. . . . . . . . . . . . . . . . . . . . . . . . . . \undefined
assigns the second token to a  control sequence and continues. . . . . . . . . . . \futurelet^{*}
expand a  control sequence into character tokens. . . . . . \string^{*}
escape character in the output of  control sequence tokens. . . . . . . . . . . . . . . . . . \escapechar^{*}
redefinable scratch  control sequence used by \dospecials. . . . .  \do
scratch  control sequence used in preloading fonts.  \preloaded
control space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \␣^{*}
convert a number to a token string. . . . . . . . . \number^{*}
numerals.  convert a number to lower-case Roman \romannumeral^{*}
convert a numeric register to displayable form.  \the^{*}
place several  copies of a picture object. . . . . . . . . . . . . . . . . . \multiput^{L}
math operator: amalgamated sum, co-product (⨿). . . . . . . . . . . . . . . . . . . . . . . . . . . . . \amalg
math operator: large  co-product (∐). . . . . . . . . . . . . . . . . . . . . . . . . . . . . \coprod
use a  copy of a box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \copy^{*}
un-box a  copy of a vbox and add it to the vertical list.  \unvcopy^{*}
list.  un-box a  copy of an hbox and add it to the horizontal  \unhcopy^{*}
copyright symbol (\begin{matrix}c \\ ◯ \end{matrix}). . . . . . . . . . . . . . . . . . . . \copyright
italic  correction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \/^{*}
math function:  cosecant. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \csc
math function: arc  cosine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \arccos
math function:  cosine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \cos
math function: hyperbolic  cosine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \cosh
math function:  cotangent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \cot
math function: hyperbolic  cotangent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \coth
use a  count register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \count^{*}

define a name for a  count register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \countdef^{*}
allocate a new  count register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newcount
increment a  counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \addtocounter^{L}
increment and reference a  counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \refstepcounter^{L}
set value of a  counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \setcounter^{L}
increment a  counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \stepcounter^{L}
current value of the third-level item  counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \theenumiii^{L}
current value of the second-level item  counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \theenumii^{L}
current value of the first-level item  counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \theenumi^{L}
current value of the fourth-level item  counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \theenumiv^{L}
current value of the sixth-level item  counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \theenumvi^{L}
current value of the fifth-level item  counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \theenumv^{L}
current value of the figure  counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \thefigure^{L}
produce the value of a counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \value^{L}
display  counter as Arabic numerals. . . . . . . . . . . . . . . . . \arabic^{L}
display  counter as footnote symbol. . . . . . . . . . . . . . . . \fnsymbol^{L}
display  counter as lower-case letter. . . . . . . . . . . . . . . . .  \alph^{L}
display  counter as lower-case Roman numerals. . . . . .  \roman^{L}
display  counter as upper-case letter. . . . . . . . . . . . . . . .  \Alph^{L}
display  counter as upper-case Roman numerals. . . . .  \Roman^{L}
define a new  counter variable. . . . . . . . . . . . . . . . . . . . . . . . . \newcounter^{L}
associate a  counter with an item-type command. . . . . \usecounter^{L}
equivalent to  \cr, end of aligned text. . . . . . . . . . . . . . . . . . . . \endline
ensure a \cr in \halign. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \crcr^{*}
tokens to insert after every  \cr or nonredundant \crcr. . . . . . . . . . . . . . . . \everycr^{*}
tokens to insert after every \cr or nonredundant  \crcr. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \everycr^{*}
create a bibliography entry. . . . . . . . . . . . . . . . . \bibitem^{L}
create a box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \mbox^{L}
create a box in dashes, with positioning. … \dashbox^{L}
create a box of current line width. . . . . . . . . . .  \line
create a box with a single column of items. \shortstack^{L}
create a box with an underline. . . . . . . . . . . . . \underbar
create a box with horizontal mode material.  \hbox^{*}
create a box, with positioning. . . . . . . . . . . . . . \makebox^{L}
create a box with vertical mode material…  \vbox^{*}
the baseline at the top. create a box with vertical mode material with \vtop^{*}
create a centered vbox in a math list. . . . . . . \vcenter^{*}
create a footnote. . . . . . . . . . . . . . . . . . . . . . . . . . \footnote
create a formula with zero height. . . . . . . . . . .  \smash
create a multicolumn entry in an aligned table.  \span^{*}
create a rule or line. . . . . . . . . . . . . . . . . . . . . . . . .  \rule^{L}
create a small sample page. . . . . . . . . . . . . . . . . \minipage^{L}
formula.  create a zero-height hbox with the width of a \hphantom
create a zero-width box with text to the left.  \llap
create a zero-width box with text to the right.  \rlap
depth of a formula.  create a zero-width vbox the height and … \vphantom
create an aligned table. . . . . . . . . . . . . . . . . . . . .  \halign^{*}
create an index entry. . . . . . . . . . . . . . . . . . . . . . .  \index^{L}
create an underlined box with text in it… \underline^{L}
create and frame a box. . . . . . . . . . . . . . . . . . . . .  \fbox^{L}
create and frame a box, with positioning… \framebox^{L}
create and name a box. . . . . . . . . . . . . . . . . . . . .  \sbox^{L}
create and name a box, with positioning. … \savebox^{L}

punctuation.  create end-of-sentence space after following .  \@^{L}
create item label for a list environment… \makelabel^{L}
create some horizontal space. . . . . . . . . . . . . . . .  \hglue
create some vertical space. . . . . . . . . . . . . . . . . .  \vglue
penalty for  creating a widow line at top of page… \widowpenalty^{*}
display. penalty for creating a widow line before a \displaywidowpenalty^{*}
enable index  creation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \makeindex^{L}
begin and  cross reference an equation. . . . . . . . . . . . . . . . \equation^{L}
define a cross reference label. . . . . . . . . . . . . . . . . . . . . . . . \label^{L}
page number of a  cross reference label. . . . . . . . . . . . . . . . . . . . . . . . \pageref^{L}
refer to a cross reference label. . . . . . . . . . . . . . . . . . . . . . . . \ref^{L}
terminate a  \csname token list. . . . . . . . . . . . . . . . . . . . . . . . \endcsname^{*}
math operator: large  cup (⋃). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigcup
math operator: large square  cup (⨆). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigsqcup
math operator: union or  cup (∪). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \cup
math operator: square  cup (⊔). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sqcup
math delimiter: left  curly brace ({). . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lbrace
math delimiter: right  curly brace (}). . . . . . . . . . . . . . . . . . . . . . . . . . . . . \rbrace
current day of the month. . . . . . . . . . . . . . . . . . .  \day^{*}
number of the  current displayed equation. . . . . . . . . . . . . \theequation^{L}
current family number. . . . . . . . . . . . . . . . . . . . . .  \fam^{*}
insert a token after the  current group is completed. . . . . . . . . . . . . . \aftergroup^{*}
line. stop reading current input file at the end of the current . \endinput^{*}
current language used for hyphenation. . . . . \language^{3}
produces  current left page heading. . . . . . . . . . . . . . . . . . \leftmark^{L}
stop reading current input file at the end of the  current line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \endinput^{*}
width of the  current line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \linewidth^{L}
discard  current line in tabbing environment. . . . . . . .  \kill^{L}
create a box of  current line width. . . . . . . . . . . . . . . . . . . . . . . . . .  \line
append a discretionary item to the  current list. . . . . . . . . . . . . . . . . . . . . . . . . . \discretionary^{*}
fetch last box off the  current list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lastbox^{*}
fetch last kern off the  current list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lastkern^{*}
fetch last penalty off the  current list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lastpenalty^{*}
fetch last skip off the  current list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lastskip^{*}
place balanced text into a mark item on the current list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mark^{*}
remove a kern just added to the  current list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \unkern^{*}
remove a penalty just added to the  current list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \unpenalty^{*}
remove a skip just added to the  current list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \unskip^{*}
show the  current lists. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \showlists^{*}
define a synonym for the current meaning of a control sequence. . . . . . \let^{*}
current month of the year. . . . . . . . . . . . . . . . . .  \month^{*}
depth of the  current page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pagedepth^{*}
amount of filll space in  current page. . . . . . . . . . . . . . . . . . . . . \pagefilllstretch^{*}
amount of fill space in  current page. . . . . . . . . . . . . . . . . . . . . . \pagefillstretch^{*}
amount of fil space in  current page. . . . . . . . . . . . . . . . . . . . . . . . \pagefilstretch^{*}
amount of glue shrinkage in  current page. . . . . . . . . . . . . . . . . . . . . . . . . . . . \pageshrink^{*}
amount of glue stretch in  current page. . . . . . . . . . . . . . . . . . . . . . . . . . . \pagestretch^{*}
penalty at the  current page break. . . . . . . . . . . . . . . . . . . \outputpenalty^{*}
current page number. . . . . . . . . . . . . . . . . . . . . . . \pageno
current page number. . . . . . . . . . . . . . . . . . . . . . . \thepage^{L}
set  current page style. . . . . . . . . . . . . . . . . . . . \thispagestyle^{L}
value of \botmark just before  current page was boxed. . . . . . . . . . . . . . . . . . . . . \topmark^{*}
insert at  current position in page. . . . . . . . . . . . . . . . . . \midinsert

produces  current right page heading. . . . . . . . . . . . . . . . \rightmark^{L}
choose a math formula based on the  current style. . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathchoice^{*}
save  current tab stops in tabbing environment. . \pushtabs^{L}
box containing current tabs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tabs
name of  current TEX format package. . . . . . . . . . . . . . . . \fmtname
version of  current TEX format package. . . . . . . . . . . . . \fmtversion
current time of day. . . . . . . . . . . . . . . . . . . . . . . . .  \time^{*}
current value of the fifth-level item counter. \theenumv^{L}
current value of the figure counter. . . . . . . . \thefigure^{L}
current value of the first-level item counter.  \theenumi^{L}
counter.  current value of the fourth-level item . . . . . \theenumiv^{L}
counter.  current value of the second-level item . . . . . \theenumii^{L}
current value of the sixth-level item counter. \theenumvi^{L}
counter.  current value of the third-level item . . . . . \theenumiii^{L}
current year of our Lord. . . . . . . . . . . . . . . . . . . .  \year^{*}
math operator:  dagger (†). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \dagger
math operator: double  dagger (‡). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ddagger
dagger symbol (†). . . . . . . . . . . . . . . . . . . . . . . . . .  \dag
double  dagger symbol (‡). . . . . . . . . . . . . . . . . . . . . . . . . .  \ddag
math relation:  dash V (⊣). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \dashv
math relation:  dash (⊢). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vdash
create a box in  dashes, with positioning. . . . . . . . . . . . . . . . . . . . \dashbox^{L}
display today’s date. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \today^{L}
date on title page. . . . . . . . . . . . . . . . . . . . . . . . . .  \date^{L}
current time of day. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \time^{*}
turn on all  debugging commands. . . . . . . . . . . . . . . . . . . . \tracingall
declaration to center lines. . . . . . . . . . . . . . . . \centering^{L}
declare document style and options. . \documentstyle^{L}
declare the title. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \title^{L}
position accents.  default horizontal kern character to \defaultskewchar^{*}
loaded.  default hyphen when a font is … \defaulthyphenchar^{*}
operators. restore default limit placement on large math \displaylimits^{*}
default Plain output routine. . . . . . . . . . . . \plainoutput
define a binary math operator. . . . . . . . . . . . . . \mathbin^{*}
define a character as a delimiter. . . . . . . . . . . . \delcode^{*}
define a cross reference label. . . . . . . . . . . . . . . .  \label^{L}
define a delimiter for math mode. . . . . . . . . \delimiter^{*}
replacement text. define a global macro with expanded . . . . . . . \xdef^{*}
define a large math operator. . . . . . . . . . . . . . . . \mathop^{*}
define a macro. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \def^{*}
define a macro globally. . . . . . . . . . . . . . . . . . . . .  \gdef^{*}
text. define a macro with expanded replacement . \edef^{*}
define a math punctuation operator. . . . . . . \mathpunct^{*}
define a math relation operator. . . . . . . . . . . . . \mathrel^{*}
define a name for a character. . . . . . . . . . . . . . . \chardef^{*}
define a name for a count register. . . . . . . . . . \countdef^{*}
define a name for a dimension register. . . . . \dimendef^{*}
define a name for a math character. . . . . \mathchardef^{*}
define a name for a math skip register. … \muskipdef^{*}
define a name for a skip register. . . . . . . . . . . . \skipdef^{*}
define a name for a token list register. . . . . . . \toksdef^{*}
define a new box register. . . . . . . . . . . . . . . . \newsavebox^{L}
define a new command. . . . . . . . . . . . . . . . . . \newcommand^{L}

define a new counter variable. . . . . . . . . . . . \newcounter^{L}
define a new environment. . . . . . . . . . . \newenvironment^{L}
define a new font family. . . . . . . . . . . . . . . . . . . . \newfam
define a new help message. . . . . . . . . . . . . . . . . . \newhelp
define a new \if command. . . . . . . . . . . . . . . . .  \newif
hyphenation.  define a new language to be used for … \newlanguage^{3}
define a new length command. . . . . . . . . . . . \newlength^{L}
define a new theorem environment. . . . . . . \newtheorem^{L}
define a set of hyphenation patterns. . . . . . . \patterns^{*}
text and script sizes.  define a symbol that will work properly in \mathpalette
control sequence. define a synonym for the current meaning of a \let^{*}
define a tab stop in tabbing environment…  \=^{L}
define an inner math subformula. . . . . . . . . . \mathinner^{*}
define an ordinary math operator. . . . . . . . . . . \mathord^{*}
define an unusual paragraph shape. . . . . . . . . \parshape^{*}
define font magnification. . . . . . . . . . . . . . . . . . . \magstep
define horizontal tabs. . . . . . . . . . . . . . . . . . . . . . . \settabs
define kerning around math in text. … \mathsurround^{*}
internal Plain TEX operation to  define math text symbols. . . . . . . . . . . . . . . . \mathhexbox
define paragraph indentation. . . . . . . . . . . . . \parindent^{*}
define space between paragraphs. . . . . . . . . . . . \parskip^{*}
define the page output routine. . . . . . . . . . . . . . \output^{*}
footnotes.  define the rule separating a page and . \footnoterule
page is formatted.  define where text will be inserted when the . \insert^{*}
select smallest defined font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tiny^{L}
defines a carriage return as \\. . . . . . . . . . . . . . \obeycr^{L}
give a token list  defining the semantics of a token. . . . . . . . . . . \meaning^{*}
end a \settabs  definition. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \columns
the following macro  definition or register setting is global. . . . . . .  \global^{*}
math function:  degree of a polynomial (deg). . . . . . . . . . . . . . .  \deg
2.5-line left math  delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Biggl
2-line left math  delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \biggl
2.5-line middle math  delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Biggm
2-line middle math  delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \biggm
2.5-line right math  delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Biggr
2-line right math  delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \biggr
1.5-line left math  delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Bigl
1-line left math  delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bigl
1.5-line middle math  delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Bigm
1-line middle math  delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bigm
1.5-line right math  delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Bigr
1-line right math  delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bigr
generate a matrix with a left brace  delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \cases
define a character as a  delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \delcode^{*}
begin a new math list with a left delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \left^{*}
width of a null  delimiter. . . . . . . . . . . . . . . . . . . . . . . \nulldelimiterspace^{*}
end a math list with a right delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \right^{*}
math  delimiter: backslash (\). . . . . . . . . . . . . . . . . . . \backslash
word  delimiter character. . . . . . . . . . . . . . . . . . . . \boundarychar^{3}
math delimiter: double vertical bar (∥). . . . . . . . . . . \Vert
define a  delimiter for math mode. . . . . . . . . . . . . . . . . . \delimiter^{*}
select math spacing of a closing  delimiter for the next item. . . . . . . . . . . . . . . . \mathclose^{*}
select math spacing of an opening  delimiter for the next item. . . . . . . . . . . . . . . . . \mathopen^{*}

math  delimiter: left angle bracket (⟨). . . . . . . . . . . . . \langle
math  delimiter: left bracket ([). . . . . . . . . . . . . . . . . . . \lbrack
math delimiter: left ceiling bracket (⌈). . . . . . . . . . . . \lceil
math  delimiter: left curly brace ({). . . . . . . . . . . . . . . \lbrace
math delimiter: left floor bracket (⌊). . . . . . . . . . . . . \lfloor
math  delimiter: left group ({). . . . . . . . . . . . . . . . . . . \lgroup
math delimiter: right angle bracket (⟩). . . . . . . . . . . \rangle
math  delimiter: right bracket (]). . . . . . . . . . . . . . . . . . \rbrack
math delimiter: right ceiling bracket (⌉). . . . . . . . . . \rceil
math  delimiter: right curly brace (}). . . . . . . . . . . . . \rbrace
math  delimiter: right floor (⌋). . . . . . . . . . . . . . . . . . . . \rfloor
math  delimiter: right group (}). . . . . . . . . . . . . . . . . . \rgroup
1.5-line math  delimiter size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Big
1-line math  delimiter size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \big
2.5-line math  delimiter size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Bigg
2-line math  delimiter size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bigg
math  delimiter: up-and-down arrow (↕). . . . . . \updownarrow
math  delimiter: upward arrow (↑). . . . . . . . . . . . . . . . \uparrow
math delimiter: vertical bar (|). . . . . . . . . . . . . . . . . . . \vert
fraction with specified rule and  delimiters. . . . . . . . . . . . . . . . . . . . . . . . . \abovewithdelims^{*}
fraction without rule with given  delimiters. . . . . . . . . . . . . . . . . . . . . . . . . . \atopwithdelims^{*}
generate a matrix without  delimiters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \matrix
fraction with rule and given  delimiters. . . . . . . . . . . . . . . . . . . . . . . . . . \overwithdelims^{*}
generate a matrix with parentheses  delimiters. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pmatrix
included material. amount by which delimiters can fail to span . . . . . \delimitershortfall^{*}
ratio for variable  delimiters times 1000. . . . . . . . . . . . . . \delimiterfactor^{*}
math Greek letter:  delta (δ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \delta
math Greek letter: capital  delta (Δ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Delta
broken line.  demerits for a penultimate … \finalhyphendemerits^{*}
demerits for adjacent incompatible lines. \adjdemerits^{*}
lines.  demerits for consecutive broken \doublehyphendemerits^{*}
mark  denoting fifth-level items. . . . . . . . . . . . . . . . \labelitemv^{L}
mark  denoting first-level items. . . . . . . . . . . . . . . . \labelitemi^{L}
mark  denoting fourth-level items. . . . . . . . . . . . . \labelitemiv^{L}
mark  denoting second-level items. . . . . . . . . . . . . \labelitemii^{L}
mark  denoting sixth-level items. . . . . . . . . . . . . . \labelitemvi^{L}
mark  denoting third-level items. . . . . . . . . . . . . \labelitemiii^{L}
maximum boxed items shown at a given  depth. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \showboxbreadth^{*}
depth of a box. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \dp^{*}
create a zero-width vbox the height and  depth of a formula. . . . . . . . . . . . . . . . . . . . . . . . . \vphantom
maximum  depth of boxes in a split box. . . . . . . . . \splitmaxdepth^{*}
maximum  depth of boxes on explicit pages. . . . . . . . \boxmaxdepth^{*}
maximum  depth of boxes on main pages. . . . . . . . . . . . . . \maxdepth^{*}
math strut with height and  depth of parentheses. . . . . . . . . . . . . . . . . . . . . . \mathstrut
depth of the current page. . . . . . . . . . . . . . . . . \pagedepth^{*}
depth of the last box on the vertical list. . \prevdepth^{*}
maximum box  depth shown. . . . . . . . . . . . . . . . . . . . . . . . . . \showboxdepth^{*}
select old-style numerals with  descenders. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \oldstyle
desired page height. . . . . . . . . . . . . . . . . . . . . . . . \pagegoal^{*}
math function: det. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \det
select one of several entries  determined by an integer. . . . . . . . . . . . . . . . . . . \ifcase^{*}
diæresis or umlaut accent (ẍ). . . . . . . . . . . . . . .  \"

show  diagnostics on the terminal. . . . . . . . . . \tracingonline^{*}
three diagonal dots in math mode (...). . . . . . . . . . . \ddots
math operator:  diamond (⋄). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \diamond
math symbol:  diamond (◇). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \Diamond^{L}
math symbol: left half  diamond (⊲). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \lhd^{L}
math symbol: right half  diamond (⊳). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \rhd^{L}
math symbol:  diamond suit (♢). . . . . . . . . . . . . . . . . . . . . . \diamondsuit
math symbol: underlined left half  diamond (⊴). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \unlhd^{L}
math symbol: underlined right half  diamond (⊵). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \unrhd^{L}
insert words into hyphenation  dictionary. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \hyphenation^{*}
ligature  digraph symbol ae (æ). . . . . . . . . . . . . . . . . . . . . .  \ae
ligature  digraph symbol capital AE (Æ). . . . . . . . . . . . .  \AE
ligature  digraph symbol capital OE (Œ). . . . . . . . . . . .  \OE
ligature  digraph symbol oe (œ). . . . . . . . . . . . . . . . . . . . .  \oe
math function:  dim. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \dim
largest permissible  dimension. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \maxdimen
use a dimension register. . . . . . . . . . . . . . . . . . . . . . . . . . \dimen^{*}
define a name for a  dimension register. . . . . . . . . . . . . . . . . . . . . . . . . \dimendef^{*}
allocate a new  dimension register. . . . . . . . . . . . . . . . . . . . . . . . . \newdimen
test two  dimensions. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ifdim^{*}
math operator:  direct sum, circle plus (⊕). . . . . . . . . . . . . . . . . .  \oplus
read a file unless  disabled by \includeonly. . . . . . . . . . . . . . . . . . \include^{L}
discard current line in tabbing environment.  \kill^{L}
discretionary hyphen. . . . . . . . . . . . . . . . . . . . . . .  \-^{*}
penalty for line break after  discretionary hyphen. . . . . . . . . . . . . . . . . \hyphenpenalty^{*}
append a  discretionary item to the current list.  \discretionary^{*}
math operator:  discretionary multiply sign. . . . . . . . . . . . . . . . .  \*
penalty for creating a widow line before a  display. . . . . . . . . . . . . . . . . . . . . . . \displaywidowpenalty^{*}
penalty for breaking between lines of a  display. . . . . . . . . . . . . . . . . . . \interdisplaylinepenalty
penalty for page break just after a  display. . . . . . . . . . . . . . . . . . . . . . . . . \postdisplaypenalty^{*}
penalty for page break just before a  display. . . . . . . . . . . . . . . . . . . . . . . . . . \predisplaypenalty^{*}
length of text preceding a  display. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \predisplaysize^{*}
alignment.  display a stack of formulas without … \displaylines
display a sub-item. . . . . . . . . . . . . . . . . . . . . . . . . \itemitem
display a symbol from a font. . . . . . . . . . . . . . .  \symbol^{L}
display an item. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \item
indentation.  display an item without hanging . . . . . . . . \textindent
display counter as Arabic numerals. . . . . . . . . \arabic^{L}
display counter as footnote symbol. . . . . . . . . \fnsymbol^{L}
display counter as lower-case letter. . . . . . . . .  \alph^{L}
display counter as lower-case Roman numerals.  \roman^{L}
display counter as upper-case letter. . . . . . . . .  \Alph^{L}
numerals. display counter as upper-case Roman . . . . . . . \Roman^{L}
indentation of  display equations. . . . . . . . . . . . . . . . . . . . . . . . \mathindent^{L}
help message to  display if user asks for help. . . . . . . . . . . . . . . . . \errhelp^{*}
tokens to insert when  display math begins. . . . . . . . . . . . . . . . . . . \everydisplay^{*}
enter  display math mode. . . . . . . . . . . . . . . . . . . . . . . . .  $$^{*}
begin  display math mode. . . . . . . . . . . . . . . . . . . . . . . . .  \[^{L}
end  display math mode. . . . . . . . . . . . . . . . . . . . . . . . .  \]^{L}
set a formula flush left in  display style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lefteqn^{L}
display the contents of a register. . . . . . . . . . . . \showthe^{*}
display today’s date. . . . . . . . . . . . . . . . . . . . . . . .  \today^{L}
convert a numeric register to  displayable form. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \the^{*}

indentation of line for  displayed equation. . . . . . . . . . . . . . . . . . . \displayindent^{*}
width of a  displayed equation. . . . . . . . . . . . . . . . . . . . \displaywidth^{*}
number of the current  displayed equation. . . . . . . . . . . . . . . . . . . . . \theequation^{L}
suppress numbering of  displayed equations. . . . . . . . . . . . . . . . . . . . . . . . \nonumber^{L}
number of lines of context to be  displayed in an error message… \errorcontextlines^{3}
process input without  displaying errors. . . . . . . . . . . . . . . . . . . . . . . . . . \batchmode^{*}
extra space above  displays. . . . . . . . . . . . . . . . . . . . . . . . . . \abovedisplayskip^{*}
extra space just below  displays. . . . . . . . . . . . . . . . . . . . . . . . . . \belowdisplayskip^{*}
unit of measure for opening up displays. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \jot
extra space above  displays following short lines. \abovedisplayshortskip^{*}
extra space just below  displays following short lines. \belowdisplayshortskip^{*}
footnote separation  distance. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footnotesep^{L}
kern a given  distance. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \kern^{*}
lower a box a given  distance. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \lower^{*}
raise a box a given  distance. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \raise^{*}
raise a box a  distance. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \raisebox^{L}
distance between marginal note and text. \marginparsep^{L}
unit of  distance in picture environment. . . . . . . . . \unitlength^{L}
tabbing fields.  distance left by \' command between … \tabbingsep^{L}
fill a space with an evenly  distributed box or rule. . . . . . . . . . . . . . . . . . . . . \xleaders^{*}
math operator: div. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \div
divide a register by a value. . . . . . . . . . . . . . . . . \divide^{*}
start a major division of a long document. . . . . . . . . . . . . . . . \part^{L}
set the magnification for the  document. . . . . . . . . . . . . . . . . . . . . . . . . . . . \magnification
start a major division of a long document. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \part^{L}
declare  document style and options. . . . . . . . . . \documentstyle^{L}
dollar sign. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \$
redefinable scratch control sequence used by \dospecials. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \do
dot accent (ẋ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \.
math operator: large circle with  dot (⨀). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigodot
math operator: centered  dot (·). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \cdot
math accent: double  dot (ẍ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ddot
math accent:  dot (ẋ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \dot
lower  dot math symbol with special spacing. . . . . .  \ldotp
math operator: circle  dot (⊙). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \odot
dot under accent (\begin{matrix}x \\ . \end{matrix}). . . . . . . . . . . . . . . . . . . . . . . .  \d
centered  dot with special spacing in math mode. . . . .  \cdotp
math symbol:  dotless i (ı). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \imath
dotless i letter (ı). . . . . . . . . . . . . . . . . . . . . . . . . . .  \i
math symbol: dotless j (ȷ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \jmath
dotless j letter (ȷ). . . . . . . . . . . . . . . . . . . . . . . . . . .  \j
fill a space with  dots. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \dotfill
three centered  dots in math mode (⋯). . . . . . . . . . . . . . . . . . . .  \cdots
three low  dots in math mode (…). . . . . . . . . . . . . . . . . . . .  \ldots
three diagonal  dots in math mode (...). . . . . . . . . . . . . . . . . . . .  \ddots
math symbol: vertical  dots (⋮). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vdots
math relation:  dotted equal (≐). . . . . . . . . . . . . . . . . . . . . . . . . . .  \doteq
math symbol: downward  double arrow (⇓). . . . . . . . . . . . . . . . . . . . . . . . . \Downarrow
math symbol: left  double arrow (⇐). . . . . . . . . . . . . . . . . . . . . . . . \Leftarrow
math symbol: left-right  double arrow (⇔). . . . . . . . . . . . . . . . . . \Leftrightarrow
math symbol: long left  double arrow (⟸). . . . . . . . . . . . . . . . . . \Longleftarrow
math symbol: long left-right  double arrow (⟺). . . . . . . . . . . . \Longleftrightarrow
math symbol: long right  double arrow (⟹). . . . . . . . . . . . . . . . . \Longrightarrow

math symbol: right  double arrow (⇒). . . . . . . . . . . . . . . . . . . . . . . \Rightarrow
math symbol: upward  double arrow (⇑). . . . . . . . . . . . . . . . . . . . . . . . . . . \Uparrow
math symbol: up-and-down  double arrow (⇕). . . . . . . . . . . . . . . . . . . . . . . \Updownarrow
space between columns in  double column text. . . . . . . . . . . . . . . . . . . . . . . \columnsep^{L}
width of rule separating  double columns. . . . . . . . . . . . . . . . . . . . . . \columnseprule^{L}
math operator:  double dagger (‡). . . . . . . . . . . . . . . . . . . . . . . . . . . \ddagger
double dagger symbol (‡). . . . . . . . . . . . . . . . . . .  \ddag
math accent: double dot (ẍ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ddot
math symbol:  double relation bar (=). . . . . . . . . . . . . . . . . . . . . \Relbar
tabular environment.  double rule separation in array or … \doublerulesep^{L}
math symbol:  double vertical bar (‖). . . . . . . . . . . . . . . . . . . \Arrowvert
math delimiter: double vertical bar (∥). . . . . . . . . . . . . . . . . . . . . . \Vert
space between  double-column floats. . . . . . . . . . . . . . . . . . . \dblfloatsep^{L}
rule between  double-column floats and text. . . . . . . . . . . \dblfigrule^{L}
space between  double-column floats and text. . . . . \dbltextfloatsep^{L}
start a new page in  double-column format. . . . . . . . . . . . . . . . . . . . \twocolumn^{L}
size of float on  double-column page. . . . . . . . . \dblfloatpagefraction^{L}
math symbol:  down arrow (↓). . . . . . . . . . . . . . . . . . . . . . . . . . . \downarrow
math symbol: left harpoon  down (↽). . . . . . . . . . . . . . . . . . . . . . . . . \leftharpoondown
math symbol: right harpoon  down (⇁). . . . . . . . . . . . . . . . . . . . . . . . \rightharpoondown
math operator: large  down triangle (▽). . . . . . . . . . . . . . . . . \bigtriangledown
fill a space with a  downward brace. . . . . . . . . . . . . . . . . . . . . \downbracefill
math symbol:  downward double arrow (⇓). . . . . . . . . . . . . . \Downarrow
draw a line over a formula. . . . . . . . . . . . . . . . . \overline^{*}
draw a line under a formula. . . . . . . . . . . . . . \underline^{*}
text.  macro to  draw the rule separating footnotes from \footnoterule^{L}
environment.  cancel  effect of one \+ command in tabbing . . . . . . .  \<^{L}
flush insertions and  eject to a new page. . . . . . . . . . . . . . . . . . . \dosupereject
flush all insertions and  eject to a new page. . . . . . . . . . . . . . . . . . . . . . \supereject
an ellipsis, equivalent to \ldots in math mode. \dots
emphasis font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \em^{L}
test for an  empty box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ifvoid^{*}
an  empty hbox. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \null
an  empty list of tokens. . . . . . . . . . . . . . . . . . . . . . . .  \empty
math symbol:  empty set (∅). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \emptyset
en entry in a file of index terms. . . . . . . . . \indexentry^{L}
enable BibTEX bibliography interface. . \bibliography^{L}
enable glossary processing. . . . . . . . . . . . . \makeglossary^{L}
enable index creation. . . . . . . . . . . . . . . . . . . . . \makeindex^{L}
enable only specific \include commands. \includeonly^{L}
punctuation.  enable special spacing after . . . . . . \nonfrenchspacing
mark text last  encountered in a split box. . . . . . . . . . . . . \splitbotmark^{*}
mark text first  encountered in a split box. . . . . . . . . . \splitfirstmark^{*}
mark text last  encountered on page just boxed. . . . . . . . . . . . \botmark^{*}
mark text first  encountered on page just boxed. . . . . . . . . . . \firstmark^{*}
encourage a line break. . . . . . . . . . . . . . . . . . . . \linebreak^{L}
encourage a page break. . . . . . . . . . . . . . . . . . . \pagebreak^{L}
show unassignments when groups  end. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tracingrestores^{*}
punctuation. create end-of-sentence space after following . . . . . . . \@^{L}
ensure a \cr in \halign. . . . . . . . . . . . . . . . . . . .  \crcr^{*}
ensure footnote line separation. . . . . . . . . . . \footstrut
enter display math mode. . . . . . . . . . . . . . . . . . .  $$^{*}

enter math mode. . . . . . . . . . . . . . . . . . . . . . . . . . .  $^{*}
space between aligned tab  entries. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tabskip^{*}
select one of several  entries determined by an integer. . . . . . . . . . . . \ifcase^{*}
amount of extra space between  entries in an index. . . . . . . . . . . . . . . . . . . . . . \indexspace^{L}
create a bibliography  entry. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bibitem^{L}
write a glossary  entry. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \glossary^{L}
create an index  entry. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \index^{L}
the template in the alignment preamble for this  entry.  suppress . . . . . . . . . . . . . . . . . . . . . . . . .  \omit^{*}
en  entry in a file of index terms. . . . . . . . . . . . \indexentry^{L}
start an  entry in a list environment. . . . . . . . . . . . . . . . .  \item^{L}
create a multicolumn  entry in an aligned table. . . . . . . . . . . . . . . . . . .  \span^{*}
multicolumn  entry in array or tabular environment. . \multicolumn^{L}
permit an alignment  entry to stick out of its column. . . . . . . . . . . \hidewidth
adds an  entry to the specified list or table. . \addcontentsline^{L}
produce TEX accents in tabbing environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \a^{L}
column separation in array  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . \arraycolsep^{L}
begin array environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \array^{L}
space between rows of array or tabular  environment. . . . . . . . . . . . . . . . . . . . . . . . . . \arraystretch^{L}
beginning of an environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \begin^{L}
begin centering  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \center^{L}
circle in a picture  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \circle^{L}
multicolumn line in array or tabular environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \cline^{L}
double rule separation in array or tabular  environment. . . . . . . . . . . . . . . . . . . . . . . . . \doublerulesep^{L}
end array  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \endarray^{L}
end centering  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \endcenter^{L}
end flush left  environment. . . . . . . . . . . . . . . . . . . . . . . . . . \endflushleft^{L}
end flush right  environment. . . . . . . . . . . . . . . . . . . . . . . . . \endflushright^{L}
end of an environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \end^{L}
end picture  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \endpicture^{L}
end bibliography  environment. . . . . . . . . . . . . . . . . . . \endthebibliography^{L}
extra space before a column in array or tabular environment. add . . . . . . . . . . . . . . . . . . \extracolsep^{L}
begin flush left  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \flushleft^{L}
begin flush right  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \flushright^{L}
indent before the label in a list  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \itemindent^{L}
start an entry in a list environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \item^{L}
space between successive items in a list  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \itemsep^{L}
discard current line in tabbing  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \kill^{L}
cancel effect of one \+ command in tabbing  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \<^{L}
define a tab stop in tabbing  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \=^{L}
move to next tab position in tabbing  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \>^{L}
put text flush right in a column in tabbing  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \'^{L}
unindents left margin one tab stop in tabbing  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \-^{L}
label width in a list  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \labelwidth^{L}
left margin of a list  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leftmargin^{L}
line in a picture environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \line^{L}
set width of lines in picture  environment. . . . . . . . . . . . . . . . . . . . . . . . . \linethickness^{L}
indent second and subsequent paragraphs in a list  environment. . . . . . . . . . . . . . . . . . . . . . . . . \listparindent^{L}
create item label for a list  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \makelabel^{L}
multicolumn entry in array or tabular  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . \multicolumn^{L}
define a new  environment. . . . . . . . . . . . . . . . . . . . . . . . \newenvironment^{L}
define a new theorem  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newtheorem^{L}
oval in a picture environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \oval^{L}
space between paragraphs within an item in a list  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \parsep^{L}

begin picture  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \picture^{L}
restore tabs stops in tabbing  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \poptabs^{L}
save current tab stops in tabbing  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pushtabs^{L}
redefine an  environment. . . . . . . . . . . . . . . . . . . . . \renewenvironment^{L}
begin a tabbed line in an outer  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \+
indents left margin one tab stop in tabbing  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \+^{L}
begin theorem with special format in math  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \proclaim
begin a tabbed line in an inner  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tabalign
half the width separating columns in a tabular  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tabcolsep^{L}
unit of distance in picture  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \unitlength^{L}
vector in a picture  environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vector^{L}
extra vertical space when  environment starts a paragraph. . . . . . . . . . . \partopsep^{L}
horizontal line in array and tabular  environments. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \hline^{L}
vertical line in array and tabular  environments. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vline^{L}
math Greek letter:  epsilon (ϵ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \epsilon
math Greek letter: variant  epsilon (ε). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \varepsilon
math relation: approximately  equal (≈). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \approx
math relation: dotted  equal (≐). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \doteq
math relation: greater or  equal (≥). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ge
math relation: greater or  equal (≥). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \geq
math relation: less or  equal (≤). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \le
math relation: less or  equal (≤). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \leq
math relation: not  equal (≠). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ne
math relation: not  equal (≠). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \neq
math relation: not  equal (≠). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \not=
math relation: precedes or  equal (⪯). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \preceq
math relation: similar or  equal (≃). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \simeq
math relation: square subset or  equal (⊑). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \sqsubseteq
math relation: square superset or  equal (⊒). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \sqsupseteq
math relation: subset or  equal (⊆). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \subseteq
math relation: successor or equal (⪰). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \succeq
math relation: superset or  equal (⊇). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \supseteq
indentation of line for displayed  equation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \displayindent^{*}
width of a displayed  equation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \displaywidth^{*}
end an  equation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \endequation^{L}
begin and cross reference an  equation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \equation^{L}
number of the current displayed  equation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \theequation^{L}
stack one  equation above another. . . . . . . . . . . . . . . . . . . . \stackrel^{L}
equation number. . . . . . . . . . . . . . . . . . . . . . . . . . .  \eqno^{*}
left equation number. . . . . . . . . . . . . . . . . . . . . . . . . . . \leqno^{*}
align a stack of equations with  equation numbers. . . . . . . . . . . . . . . . . . . . . . . . \eqalignno
align a stack of equations with left  equation numbers. . . . . . . . . . . . . . . . . . . . . . . \leqalignno
align a stack of  equations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \eqalign
indentation of display  equations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathindent^{L}
suppress numbering of displayed  equations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \nonumber^{L}
align a stack of  equations with equation numbers. . . . . . . . . \eqalignno
align a stack of  equations with left equation numbers. … \leqalignno
math relation:  equivalence (≡). . . . . . . . . . . . . . . . . . . . . . . . . . . .  \equiv
superscript,  equivalent to ^. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sp
equivalent to \cr, end of aligned text. . . . . . . \endline
an ellipsis, equivalent to \ldots in math mode. . . . . . . . . \dots
equivalent to \par. . . . . . . . . . . . . . . . . . . . . . . . . . \endgraf
subscript,  equivalent to _. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sb

maximum overrun before overfull vbox  error. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vfuzz^{*}
number of lines of context to be displayed in an  error message. . . . . . . . . . . . . . . . . . . \errorcontextlines^{3}
write balanced  error message to the terminal. . . . . . . . . . . . \errmessage^{*}
process input without displaying  errors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \batchmode^{*}
limit for bad hbox  errors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \hbadness^{*}
process input without stopping for  errors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \nonstopmode^{*}
process TEX input without pausing for normal  errors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \scrollmode^{*}
limit for bad vbox  errors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vbadness^{*}
pause for normal  errors while processing TEX input. … \errorstopmode^{*}
sequence tokens.  escape character in the output of control  \escapechar^{*}
math Greek letter:  eta (η). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \eta
fill a space with an  evenly distributed box or rule. . . . . . . . . . . . . . \xleaders^{*}
height of text,  excluding head and foot. . . . . . . . . . . . . . . . . \textheight^{L}
execute a command from the terminal. . . . . . \typein^{L}
show commands before they are  executed. . . . . . . . . . . . . . . . . . . . . . . . . . \tracingcommands^{*}
math symbol:  exists quantifier (∃). . . . . . . . . . . . . . . . . . . . . . . .  \exists
math function:  exp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \exp
tokens.  expand a control sequence into character … \string^{*}
token.  expand the token following the next … \expandafter^{*}
non-space is found. read, expand, then ignore tokens until a …. \ignorespaces^{*}
show macros as they are  expanded. . . . . . . . . . . . . . . . . . . . . . . . . . . . \tracingmacros^{*}
define a macro with  expanded replacement text. . . . . . . . . . . . . . . . .  \edef^{*}
define a global macro with  expanded replacement text. . . . . . . . . . . . . . . . .  \xdef^{*}
name.  begin  expanding tokens to construct a command .  \csname^{*}
suppress  expansion of a character. . . . . . . . . . . . . . . . . . . \noexpand^{*}
penalty for line break after  explicit hyphen. . . . . . . . . . . . . . . . . . . . \exhyphenpenalty^{*}
maximum depth of boxes on  explicit pages. . . . . . . . . . . . . . . . . . . . . . . . . . \boxmaxdepth^{*}
select bold  extended font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bf
11 point bold  extended font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvbf^{L}
14 point bold  extended font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \frtnbf^{L}
9 point bold  extended font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ninbf^{L}
17 point bold  extended font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \svtnbf^{L}
12 point bold  extended font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twlbf^{L}
20 point bold  extended font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twtybf^{L}
10 point math  extension symbol font. . . . . . . . . . . . . . . . . . . . . .  \tenex
gives the  external file name for the given font. . . . . . . . \fontname^{*}
extra space above displays. . . . . . . \abovedisplayskip^{*}
following short lines.  extra space above displays . \abovedisplayshortskip^{*}
extra space added to top of page. . . . . . . . . \topmargin^{L}
extra space after subscript or superscript. \scriptspace^{*}
tabular environment. add extra space before a column in array or \extracolsep^{L}
amount of  extra space between entries in an index. . \indexspace^{L}
turn off  extra space between lines. . . . . . . . . \offinterlineskip
isn’t enough.  extra space between lines if \baselineskip    \lineskip^{*}
extra space in badly-stretched lines. \emergencystretch^{3}
extra space just below displays… \belowdisplayskip^{*}
displays following short lines.  extra space just below . . . . . \belowdisplayshortskip^{*}
add  extra vertical space. . . . . . . . . . . . . . . . . . . . . . . \addvspace^{L}
a paragraph.  extra vertical space when environment starts \partopsep^{L}
set a character’s space  factor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \sfcode^{*}
set a character’s spacing  factor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \spacefactor^{*}
amount by which delimiters can  fail to span included material. . \delimitershortfall^{*}
test always  false. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \iffalse^{*}

bold font  family. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bffam
italic font  family. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \itfam
define a new font  family. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newfam
slanted font  family. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \slfam
typewriter font  family. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ttfam
current  family number. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \fam^{*}
space between a box and its contents in  \fbox and \framebox. . . . . . . . . . . . . . . . . . . . . . \fboxsep^{L}
fetch last box off the current list. . . . . . . . . . . . \lastbox^{*}
fetch last kern off the current list. . . . . . . . . . \lastkern^{*}
fetch last penalty off the current list. … \lastpenalty^{*}
fetch last skip off the current list. . . . . . . . . . . \lastskip^{*}
place an accent over the next math  field. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathaccent^{*}
distance left by \' command between tabbing  fields. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tabbingsep^{L}
current value of the  fifth-level item counter. . . . . . . . . . . . . . . . . . . . . \theenumv^{L}
mark denoting  fifth-level items. . . . . . . . . . . . . . . . . . . . . . . . . . \labelitemv^{L}
width of left margin in  fifth-level list. . . . . . . . . . . . . . . . . . . . . . . . . . . \leftmarginv^{L}
end a floating  figure. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \endfigure^{L}
begin a floating figure. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \figure^{L}
current value of the  figure counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . \thefigure^{L}
generate a list of  figures. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \listoffigures^{L}
flush  figures and start a new page. . . . . . . . . . . . . . \clearpage^{L}
right-hand page. flush figures and tables and start a new . \cleardoublepage^{L}
add text to table contents,  figures, or tables. . . . . . . . . . . . . . . . . . . . . \addtocontents^{L}
amount of  fil space in current page. . . . . . . . . . . . \pagefilstretch^{*}
break a page unless there is a better  \filbreak below. . . . . . . . . . . . . . . . . . . . . . . . . . \filbreak
close an input  file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \closein^{*}
close an output  file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \closeout^{*}
test for end of  file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ifeof^{*}
read a  file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \input^{*}
name of the principal input  file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \jobname^{*}
pause after each line is read from a  file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pausing^{*}
read a line from a file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \read^{*}
allocate a new input  file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newread
allocate a new output  file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newwrite
write a token list to a file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \write^{*}
stop reading current input  file at the end of the current line. . . . . . . . . . . \endinput^{*}
open a  file for input. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \openin^{*}
open a  file for output. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \openout^{*}
terminate TEX and write a format file: INITEX only. . . . . . . . . . . . . . . . . . . . . . . . . . \dump^{*}
gives the external  file name for the given font. . . . . . . . . . . . . . . . \fontname^{*}
en entry in a  file of index terms. . . . . . . . . . . . . . . . . . . . . . . \indexentry^{L}
write to the log file only. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \wlog
read a  file unless disabled by \includeonly. . . . . . . \include^{L}
suppress writing all auxiliary  files. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \nofiles^{L}
fill a space with a downward brace… \downbracefill
fill a space with a left arrow. . . . . . . . . \leftarrowfill
fill a space with a repeated box or rule. . . . . \leaders^{*}
fill a space with a right arrow. . . . . . . \rightarrowfill
fill a space with a rule. . . . . . . . . . . . . . . . . . . . \hrulefill
rule.  fill a space with an evenly distributed box or \xleaders^{*}
fill a space with an upward brace. . . . . . . \upbracefill
fill a space with dots. . . . . . . . . . . . . . . . . . . . . . . \dotfill
amount of  fill space in current page. . . . . . . . . . \pagefillstretch^{*}

amount of  filll space in current page. . . . . . . . . \pagefilllstretch^{*}
terminate L^{A}TEX and flush the final page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \stop^{L}
finish processing input. . . . . . . . . . . . . . . . . . . . . .  \bye
current value of the  first-level item counter. . . . . . . . . . . . . . . . . . . . . \theenumi^{L}
mark denoting  first-level items. . . . . . . . . . . . . . . . . . . . . . . . . . \labelitemi^{L}
width of left margin in  first-level list. . . . . . . . . . . . . . . . . . . . . . . . . . . \leftmargini^{L}
math symbol: flat (♭). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \flat
vertical space around a  float in the middle of a page. . . . . . . . . . . . . . \intextsep^{L}
size of float on double-column page. \dblfloatpagefraction^{L}
end a  floating figure. . . . . . . . . . . . . . . . . . . . . . . . . . . . \endfigure^{L}
begin a  floating figure. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \figure^{L}
place rule between text and bottom  floats. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \botfigrule^{L}
fraction of column for bottom  floats. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bottomfraction^{L}
space between double-column  floats. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \dblfloatsep^{L}
fraction of two-column page for top  floats. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \dbltopfraction^{L}
portion of page that may be occupied by  floats. . . . . . . . . . . . . . . . . . . . . . . . . . . \floatpagefraction^{L}
space between  floats. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \floatsep^{L}
fraction of column for top  floats. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \topfraction^{L}
rule between double-column  floats and text. . . . . . . . . . . . . . . . . . . . . . . . . . \dblfigrule^{L}
space between double-column  floats and text. . . . . . . . . . . . . . . . . . . . \dbltextfloatsep^{L}
place rule between top  floats and text. . . . . . . . . . . . . . . . . . . . . . . . . . \topfigrule^{L}
space between  floats and the text. . . . . . . . . . . . . . . . . . . . \textfloatsep^{L}
math delimiter: left  floor bracket (⌊). . . . . . . . . . . . . . . . . . . . . . . . . . . . \lfloor
math delimiter: right  floor (⌋). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \rfloor
flush all insertions and eject to a new page. \supereject
flush figures and start a new page. . . . . . . . \clearpage^{L}
new right-hand page. flush figures and tables and start a \cleardoublepage^{L}
flush insertions and eject to a new page. \dosupereject
end  flush left environment. . . . . . . . . . . . . . . . . \endflushleft^{L}
begin  flush left environment. . . . . . . . . . . . . . . . . . . . \flushleft^{L}
set a formula  flush left in display style. . . . . . . . . . . . . . . . . . . . \lefteqn^{L}
flush lines left. . . . . . . . . . . . . . . . . . . . . . . . . . \raggedright^{L}
flush lines right. . . . . . . . . . . . . . . . . . . . . . . . . . \raggedleft^{L}
end  flush right environment. . . . . . . . . . . . . . \endflushright^{L}
begin  flush right environment. . . . . . . . . . . . . . . . . . \flushright^{L}
environment. put text flush right in a column in tabbing . . . . . . . . . . \'^{L}
flush right text on a line. . . . . . . . . . . . . . . . . . \rightline
flush text left on a line. . . . . . . . . . . . . . . . . . . . . \leftline
terminate L^{A}TEX and  flush the final page. . . . . . . . . . . . . . . . . . . . . . . . .  \stop^{L}
global.  the  following macro definition or register setting is  \global^{*}
as parameters.  the  following macro may have multiple paragraphs  \long^{*}
another macro.  the  following macro must not be called from …  \outer^{*}
create end-of-sentence space after  following punctuation. . . . . . . . . . . . . . . . . . . . . .  \@^{L}
extra space above displays  following short lines. . . . . . . . \abovedisplayshortskip^{*}
extra space just below displays  following short lines. . . . . . . . \belowdisplayshortskip^{*}
expand the token  following the next token. . . . . . . . . . . . . . . . \expandafter^{*}
select bold extended  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bf
math mode calligraphic letters  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \cal
8 point italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \egtit^{L}
8 point L^{A}TEX symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \egtly^{L}
8 point math italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \egtmi^{L}
8 point Roman  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \egtrm^{L}
8 point math symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \egtsy^{L}

11 point bold extended  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvbf^{L}
11 point italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvit^{L}
11 point L^{A}TEX symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvly^{L}
11 point math italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvmi^{L}
11 point Roman  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvrm^{L}
11 point sans serif  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvsf^{L}
11 point slanted  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvsl^{L}
11 point math symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvsy^{L}
11 point typewriter  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvtt^{L}
emphasis  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \em^{L}
5 point boldface Roman  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fivebf
5 point math italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fivei
5 point Roman  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fiverm
5 point math symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fivesy
5 point L^{A}TEX symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \fivly^{L}
5 point math italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \fivmi^{L}
5 point Roman  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \fivrm^{L}
5 point math symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \fivsy^{L}
load a  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \font^{*}
gives the external file name for the given  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fontname^{*}
14 point bold extended  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \frtnbf^{L}
14 point L^{A}TEX symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \frtnly^{L}
14 point math italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \frtnmi^{L}
14 point Roman  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \frtnrm^{L}
14 point math symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \frtnsy^{L}
select largest available font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \Huge^{L}
hyphenation character for this  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \hyphenchar^{*}
select italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \it
load a  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \load^{L}
math mode italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \mit
select a  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newfont^{L}
9 point bold extended  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ninbf^{L}
9 point italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ninit^{L}
9 point L^{A}TEX symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ninly^{L}
9 point math italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ninmi^{L}
9 point Roman  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ninrm^{L}
9 point math symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ninsy^{L}
9 point typewriter  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \nintt^{L}
select Roman  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \rm
select small caps  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sc^{L}
7 point bold Roman  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \sevenbf
7 point math italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \seveni
7 point Roman  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \sevenrm
7 point math symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \sevensy
7 point italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sevit^{L}
7 point L^{A}TEX symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sevly^{L}
7 point math italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sevmi^{L}
7 point Roman  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sevrm^{L}
7 point math symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sevsy^{L}
select sans serif  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sf^{L}
6 point L^{A}TEX symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sixly^{L}
6 point math italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sixmi^{L}
6 point Roman  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sixrm^{L}

6 point math symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sixsy^{L}
select slanted  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sl
17 point bold extended  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \svtnbf^{L}
17 point L^{A}TEX symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \svtnly^{L}
17 point math italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \svtnmi^{L}
17 point Roman  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \svtnrm^{L}
17 point math symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \svtnsy^{L}
display a symbol from a  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \symbol^{L}
10 point bold font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tenbf
10 point circle  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tencirc^{L}
10 point wide circle  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tencircw^{L}
10 point math extension symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tenex
10 point math italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \teni
10 point text italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tenit
10 point line  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tenln^{L}
10 point wide line  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tenlnw^{L}
10 point L^{A}TEX symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tenly^{L}
10 point math italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tenmi^{L}
10 point Roman  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tenrm
10 point sans serif  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tensf^{L}
10 point slanted  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tensl
10 point math symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tensy
10 point typewriter  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tentt
select smallest defined  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tiny^{L}
show characters not in the  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tracinglostchars^{*}
select typewriter  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tt
12 point bold extended  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twlbf^{L}
12 point italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twlit^{L}
12 point L^{A}TEX symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twlly^{L}
12 point math italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twlmi^{L}
12 point Roman  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twlrm^{L}
12 point sans serif  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twlsf^{L}
12 point slanted  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twlsl^{L}
12 point math symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twlsy^{L}
12 point typewriter  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twltt^{L}
20 point bold extended  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \twtybf^{L}
20 point L^{A}TEX symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \twtyly^{L}
20 point math italic  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \twtymi^{L}
20 point Roman  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \twtyrm^{L}
20 point math symbol  font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \twtysy^{L}
bold  font family. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bffam
italic  font family. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \itfam
define a new  font family. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \newfam
slanted  font family. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \slfam
typewriter  font family. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ttfam
select  font for non-math text. . . . . . . . . . . . . . . . . . . . . \textfont^{*}
select  font for small math scripts. . . . . . . . . . . . . . \scriptfont^{*}
select  font for very small math scripts. . \scriptscriptfont^{*}
select font four steps larger than normal size. . . . . . \huge^{L}
default hyphen when a  font is loaded. . . . . . . . . . . . . . . . . . . \defaulthyphenchar^{*}
suppress right justification of typewriter  font lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . \ttraggedright
define  font magnification. . . . . . . . . . . . . . . . . . . . . . . . . . \magstep
select  font one step larger than normal size. . . . . . .  \large^{L}

big  font size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \big^{L}
bigger  font size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Big^{L}
biggest  font size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \BIG^{L}
select footnote  font size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footnotesize^{L}
select normal  font size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \normalsize^{L}
select sub-subscript  font size. . . . . . . . . . . . . . . . . . . . . . . . . . \scriptscriptsize^{L}
select subscript or superscript  font size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \scriptsize^{L}
select small  font size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \small^{L}
bold  font style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bf^{L}
select  font three steps larger than normal size. . . . .  \LARGE^{L}
select  font two steps larger than normal size. . . . . .  \Large^{L}
a  font with no characters. . . . . . . . . . . . . . . . . . . . \nullfont^{*}
set a  font-related parameter. . . . . . . . . . . . . . . . . . . . \fontdimen^{*}
select bold math italic and symbol  fonts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \boldmath^{L}
scratch control sequence used in preloading  fonts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \preloaded
unselect bold math italic and symbol  fonts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \unboldmath^{L}
constructs a box with the page  foot. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \makefootline
height of text, excluding head and  foot. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \textheight^{L}
page  foot line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footline
height of page  footer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footheight^{L}
space between text and page  footer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footskip^{L}
create a  footnote. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footnote
penalty for breaking between lines of a  footnote. . . . . . . . . . . . . . . . \interfootnotelinepenalty
select  footnote font size. . . . . . . . . . . . . . . . . . . . . \footnotesize^{L}
place a footnote in a caption or other vertical list. \vfootnote
insertion class for  footnote inserts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footins
ensure  footnote line separation. . . . . . . . . . . . . . . . . . . \footstrut
insert a  footnote mark without text. . . . . . . . . . . \footnotemark^{L}
produces the  footnote number. . . . . . . . . . . . . . . . . . . . . . . \thefootnote^{L}
footnote separation distance. . . . . . . . . . . . \footnotesep^{L}
display counter as  footnote symbol. . . . . . . . . . . . . . . . . . . . . . . . . . . \fnsymbol^{L}
produce  footnote text without a mark. . . . . . . . . \footnotetext^{L}
add  footnote to title page. . . . . . . . . . . . . . . . . . . . . . . \thanks^{L}
space between main text and  footnotes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footins^{L}
define the rule separating a page and  footnotes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footnoterule
interline penalty for  footnotes. . . . . . . . . . . . . . . \interfootnotelinepenalty^{L}
macro to draw the rule separating  footnotes from text. . . . . . . . . . . . . . . . . . . \footnoterule^{L}
math symbol:  for-all quantifier (∀). . . . . . . . . . . . . . . . . . . . . . . . \forall
convert a numeric register to displayable form. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \the^{*}
start a new page in single-column  format. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \onecolumn^{L}
start a new page in double-column  format. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \twocolumn^{L}
terminate TEX and write a format file: INITEX only. . . . . . . . . . . . . . . . . . . \dump^{*}
begin theorem with special  format in math environment. . . . . . . . . . . . . . . \proclaim
name of current TEX  format package. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fmtname
version of current TEX  format package. . . . . . . . . . . . . . . . . . . . . . . . . . \fmtversion
where text will be inserted when the page is  formatted.  define . . . . . . . . . . . . . . . . . . . . . . . \insert^{*}
colon in a math  formula. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \colon
create a zero-height hbox with the width of a  formula. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \hphantom
horizontal brace over a math  formula. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \overbrace
left arrow over a math  formula. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \overleftarrow
draw a line over a  formula. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \overline^{*}
right arrow over a math  formula. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \overrightarrow
use the space taken by a  formula. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \phantom

specified root of a  formula. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \root
square root of a  formula. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sqrt
horizontal brace under a math  formula. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \underbrace
draw a line under a  formula. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \underline^{*}
a zero-width vbox the height and depth of a formula. create . . . . . . . . . . . . . . . . . . . . . . . . \vphantom
choose a math  formula based on the current style. . . . . . . \mathchoice^{*}
set a  formula flush left in display style. . . . . . . . . . . \lefteqn^{L}
create a formula with zero height. . . . . . . . . . . . . . . . . . . . \smash
medium space in math  formulas. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \medmuskip^{*}
thick space in math  formulas. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \thickmuskip^{*}
thin space in math  formulas. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \thinmuskip^{*}
display a stack of  formulas without alignment. . . . . . . . . . . \displaylines
expand, then ignore tokens until a non-space is  found.  read, . . . . . . . . . . . . . . . . . . . . . . \ignorespaces^{*}
select font  four steps larger than normal size. . . . . . . . . . .  \huge^{L}
current value of the  fourth-level item counter. . . . . . . . . . . . . . . . . \theenumiv^{L}
mark denoting  fourth-level items. . . . . . . . . . . . . . . . . . . . . . \labelitemiv^{L}
width of left margin in  fourth-level list. . . . . . . . . . . . . . . . . . . . . . . \leftmarginiv^{L}
generate a fraction. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \frac^{L}
fraction of column for bottom floats. \bottomfraction^{L}
fraction of column for top floats. . . . . . . . \topfraction^{L}
floats. fraction of two-column page for top \dbltopfraction^{L}
fraction with a rule. . . . . . . . . . . . . . . . . . . . . . . . .  \over^{*}
fraction with rule and given delimiters. \overwithdelims^{*}
fraction with rule thickness. . . . . . . . . . . . . . . . .  \above^{*}
delimiters.  fraction with specified rule and …. \abovewithdelims^{*}
fraction without a rule with braces. . . . . . . . .  \brace
fraction without a rule with brackets. . . . . . .  \brack
fraction without a rule with parentheses. … \choose
fraction without rule. . . . . . . . . . . . . . . . . . . . . . .  \atop^{*}
delimiters.  fraction without rule with given …. \atopwithdelims^{*}
protect  fragile commands and moving arguments… \protect^{L}
math symbol: imaginary,  Fraktur I (ℑ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Im
math symbol: real,  Fraktur R (ℜ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Re
create and  frame a box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \fbox^{L}
create and  frame a box, with positioning. . . . . . . . . . . . . . \framebox^{L}
put a  frame around some text. . . . . . . . . . . . . . . . . . . .  \frame^{L}
between a box and its contents in \fbox and  \framebox.  space . . . . . . . . . . . . . . . . . . . . . . \fboxsep^{L}
thickness of rule surrounding  framed box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fboxrule^{L}
math relation:  frown (⌢). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \frown
math  function: arc cosine. . . . . . . . . . . . . . . . . . . . . . . .  \arccos
math  function: arc sine. . . . . . . . . . . . . . . . . . . . . . . . . . . \arcsin
math  function: arc tangent. . . . . . . . . . . . . . . . . . . . . . . \arctan
math  function: arg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \arg
math  function: cosecant. . . . . . . . . . . . . . . . . . . . . . . . . .  \csc
math  function: cosine. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \cos
math  function: cotangent. . . . . . . . . . . . . . . . . . . . . . . . .  \cot
math  function: degree of a polynomial (deg). . . . . .  \deg
math  function: det. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \det
math  function: dim. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \dim
math  function: exp. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \exp
math  function: gcd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \gcd
math  function: hom. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \hom
math  function: hyperbolic cosine. . . . . . . . . . . . . . . . .  \cosh

math  function: hyperbolic cotangent. . . . . . . . . . . . .  \coth
math  function: hyperbolic sine. . . . . . . . . . . . . . . . . . .  \sinh
math  function: hyperbolic tangent. . . . . . . . . . . . . . . .  \tanh
math  function: inf. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \inf
math  function: ker. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ker
math  function: lg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \lg
math  function: lim. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \lim
math  function: limit infimum (liminf). . . . . . . . . . . \liminf
math  function: limit supremum (limsup). . . . . . . . . \limsup
math  function: ln. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ln
math  function: log. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \log
math  function: max. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \max
math  function: min. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \min
math  function: mod within parentheses ( (mod )).  \pmod
math  function: probability (Pr). . . . . . . . . . . . . . . . . .  \Pr
math  function: secant. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sec
math  function: sine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sin
math  function: sup. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sup
math  function: tangent. . . . . . . . . . . . . . . . . . . . . . . . . . .  \tan
math Greek letter:  gamma (γ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \gamma
math Greek letter: capital  gamma (Γ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Gamma
math function:  gcd. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \gcd
generate a fraction. . . . . . . . . . . . . . . . . . . . . . . . .  \frac^{L}
generate a list of figures. . . . . . . . . . . . . \listoffigures^{L}
generate a list of tables. . . . . . . . . . . . . . . \listoftables^{L}
columns. generate a matrix labeled on rows and \bordermatrix
generate a matrix with a left brace delimiter.  \cases
generate a matrix with parentheses delimiters. \pmatrix
generate a matrix without delimiters. . . . . . .  \matrix
generate a short amount of verbatim text…  \verb^{L}
generate a table of contents. . . . . . . \tableofcontents^{L}
generate an in-text citation of a reference…  \cite^{L}
German letter: sharp s (ß). . . . . . . . . . . . . . . . . .  \ss
math relation: gets (←). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \gets
token.  give a token list defining the semantics of a . \meaning^{*}
following macro definition or register setting is  global.  the . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \global^{*}
define a  global macro with expanded replacement text.  \xdef^{*}
set  global page style. . . . . . . . . . . . . . . . . . . . . . . . . . \pagestyle^{L}
override  \global specifications. . . . . . . . . . . . . . . . . . . \globaldefs^{*}
define a macro globally. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \gdef^{*}
write a  glossary entry. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \glossary^{L}
enable  glossary processing. . . . . . . . . . . . . . . . . . . . \makeglossary^{L}
amount of  glue shrinkage in current page. . . . . . . . . . . \pageshrink^{*}
amount of  glue stretch in current page. . . . . . . . . . . . \pagestretch^{*}
grave accent (è). . . . . . . . . . . . . . . . . . . . . . . . . . . .  \`
math accent:  grave (x̀). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \grave
math relation: much  greater (≫). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \gg
math relation:  greater or equal (≥). . . . . . . . . . . . . . . . . . . . . . . .  \ge
math relation:  greater or equal (≥). . . . . . . . . . . . . . . . . . . . . . . .  \geq
math  Greek letter: alpha (α). . . . . . . . . . . . . . . . . . . . .  \alpha
math  Greek letter: beta (β). . . . . . . . . . . . . . . . . . . . . .  \beta
math  Greek letter: capital delta (Δ). . . . . . . . . . . . . .  \Delta
math  Greek letter: capital gamma (Γ). . . . . . . . . . . .  \Gamma

math  Greek letter: capital lambda (Λ). . . . . . . . . . . . \Lambda
math  Greek letter: capital omega (Ω). . . . . . . . . . . .  \Omega
math  Greek letter: capital phi (Φ). . . . . . . . . . . . . . . .  \Phi
math  Greek letter: capital pi (Π). . . . . . . . . . . . . . . . .  \Pi
math  Greek letter: capital psi (Ψ). . . . . . . . . . . . . . . .  \Psi
math  Greek letter: capital sigma (Σ). . . . . . . . . . . . .  \Sigma
math  Greek letter: capital theta (Θ). . . . . . . . . . . . .  \Theta
math  Greek letter: capital upsilon (Υ). . . . . . . . . . . . \Upsilon
math  Greek letter: capital xi (Ξ). . . . . . . . . . . . . . . . .  \Xi
math  Greek letter: chi (χ). . . . . . . . . . . . . . . . . . . . . . . .  \chi
math  Greek letter: delta (δ). . . . . . . . . . . . . . . . . . . . . .  \delta
math  Greek letter: epsilon (ϵ). . . . . . . . . . . . . . . . . . . . \epsilon
math  Greek letter: eta (η). . . . . . . . . . . . . . . . . . . . . . . .  \eta
math  Greek letter: gamma (γ). . . . . . . . . . . . . . . . . . .  \gamma
math  Greek letter: kappa (κ). . . . . . . . . . . . . . . . . . . . .  \kappa
math  Greek letter: lambda (λ). . . . . . . . . . . . . . . . . . . \lambda
math  Greek letter: mu (μ). . . . . . . . . . . . . . . . . . . . . . .  \mu
math  Greek letter: nu (ν). . . . . . . . . . . . . . . . . . . . . . . .  \nu
math  Greek letter: omega (ω). . . . . . . . . . . . . . . . . . . .  \omega
math  Greek letter: phi (ϕ). . . . . . . . . . . . . . . . . . . . . . .  \phi
math  Greek letter: pi (π). . . . . . . . . . . . . . . . . . . . . . . . .  \pi
math  Greek letter: psi (ψ). . . . . . . . . . . . . . . . . . . . . . .  \psi
math  Greek letter: rho (ρ). . . . . . . . . . . . . . . . . . . . . . .  \rho
math  Greek letter: sigma (σ). . . . . . . . . . . . . . . . . . . . .  \sigma
math  Greek letter: tau (τ). . . . . . . . . . . . . . . . . . . . . . .  \tau
math  Greek letter: theta (θ). . . . . . . . . . . . . . . . . . . . . .  \theta
math  Greek letter: upsilon (υ). . . . . . . . . . . . . . . . . . . \upsilon
math  Greek letter: variant epsilon (ε). . . . . . . . . \varepsilon
math  Greek letter: variant phi (φ). . . . . . . . . . . . . . . . \varphi
math Greek letter: variant pi (ϖ). . . . . . . . . . . . . . . . \varpi
math  Greek letter: variant rho (ϱ). . . . . . . . . . . . . . . . \varrho
math  Greek letter: variant sigma (ς). . . . . . . . . . . . . \varsigma
math  Greek letter: variant theta (ϑ). . . . . . . . . . . . . \vartheta
math  Greek letter: xi (ξ). . . . . . . . . . . . . . . . . . . . . . . . .  \xi
math  Greek letter: zeta (ζ). . . . . . . . . . . . . . . . . . . . . . .  \zeta
begin a  group. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \begingroup^{*}
begin a  group. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bgroup
end a  group. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \egroup
end a  group. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \endgroup^{*}
show unassignments when  groups end. . . . . . . . . . . . . . . . . . . . . . . . \tracingrestores^{*}
ensure a \cr in  \halign. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \crcr^{*}
end a line in  \halign aligned text. . . . . . . . . . . . . . . . . . . . . . .  \cr^{*}
field separator in  \halign or \valign. . . . . . . . . . . . . . . . . . . . . . . .  &^{*}
insert unaligned material in  \halign or \valign. . . . . . . . . . . . . . . . . . . . . . . . \noalign^{*}
start an  \halign with \tabskip initialized to zero. . \ialign
hanging indent of a paragraph. . . . . . . . . . . . \@hangfrom^{L}
set  hanging indentation. . . . . . . . . . . . . . . . . . . . . \hangindent^{*}
display an item without  hanging indentation. . . . . . . . . . . . . . . . . . . . . \textindent
number of lines.  hanging indentation changes after specified \hangafter^{*}
math symbol: left  harpoon down (↽). . . . . . . . . . . . . . . . \leftharpoondown
math symbol: right  harpoon down (⇁). . . . . . . . . . . . . . . \rightharpoondown
math symbol: right left  harpoon (⇀ ↽). . . . . . . . . . . . . . . . . . . . \rightleftharpoons
math symbol: left  harpoon up (↼). . . . . . . . . . . . . . . . . . . . . \leftharpoonup

math symbol: right  harpoon up (⇀). . . . . . . . . . . . . . . . . . . . \rightharpoonup
math accent: hat (x̂). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \hat
math accent: wide  hat (x̂). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \widehat
pound,  hatch mark, sharp sign, octothorpe (#). . . . .  \#
math symbol: Planck’s constant or  h-bar (h̄). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \hbar
width of a paragraph or  \hbox. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \hsize^{*}
an empty  hbox. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \null
un-box an  hbox and add it to the horizontal list. . . . . . . \unhbox^{*}
un-box a copy of an  hbox and add it to the horizontal list. . . . . . . \unhcopy^{*}
tokens to insert when an  hbox begins. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \everyhbox^{*}
limit for bad  hbox errors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \hbadness^{*}
maximum overrun before overfull hbox messages occur. . . . . . . . . . . . . . . . . . . . . . . \hfuzz^{*}
store an  hbox or vbox in a box register. . . . . . . . . . . . . . \setbox^{*}
create a zero-height  hbox with the width of a formula. . . . . . . . . . \hphantom
height of page  header. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \headheight^{L}
constructs a box with the page  header. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \makeheadline
space between page  header and text. . . . . . . . . . . . . . . . . . . . . . . . . . . . \headsep^{L}
produces current left page  heading. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leftmark^{L}
produces current right page  heading. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \rightmark^{L}
set  heading for right pages. . . . . . . . . . . . . . . . . . . \markright^{L}
set  headings for left and right pages. . . . . . . . . . . \markboth^{L}
math symbol:  heart suit (♡). . . . . . . . . . . . . . . . . . . . . . . . . . . . \heartsuit
justify page bottoms to the same  height. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \flushbottom^{L}
justify page bottoms to the same  height. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \normalbottom
put page contents in a box of the proper  height. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pagebody
desired page  height. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pagegoal^{*}
justify page bottoms to their natural  height. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \raggedbottom^{L}
create a formula with zero height. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \smash
create a zero-width vbox the  height and depth of a formula. . . . . . . . . . . . . \vphantom
math strut with  height and depth of parentheses. . . . . . . . . . \mathstrut
height of a box. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ht^{*}
height of page footer. . . . . . . . . . . . . . . . . . . . \footheight^{L}
height of page header. . . . . . . . . . . . . . . . . . . . \headheight^{L}
natural  height of page so far. . . . . . . . . . . . . . . . . . . . . . \pagetotal^{*}
height of text, excluding head and foot. . \textheight^{L}
height of text on a page or \vbox. . . . . . . . . . .  \vsize^{*}
help message to display if user asks for  help. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \errhelp^{*}
set @ non-alphabetic to  hide internal commands. . . . . . . . . . . . . . . . \makeatother^{L}
math function: hom. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \hom
math symbol:  hook left arrow (↩). . . . . . . . . . . . . . . . . \hookleftarrow
math symbol: left  hook ( ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \lhook
math operator: logical not,  hook (¬). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \lnot
math symbol: right  hook ( ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \rhook
math symbol:  hook right arrow (↪). . . . . . . . . . . . . . \hookrightarrow
test for a  horizontal box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ifhbox^{*}
lower left piece of a  horizontal brace. . . . . . . . . . . . . . . . . . . . . . . . . . . . \braceld
upper left piece of a  horizontal brace. . . . . . . . . . . . . . . . . . . . . . . . . . . . \bracelu
lower right piece of a  horizontal brace. . . . . . . . . . . . . . . . . . . . . . . . . . . . \bracerd
upper right piece of a  horizontal brace. . . . . . . . . . . . . . . . . . . . . . . . . . . . \braceru
horizontal brace over a math formula. . . . . \overbrace
horizontal brace under a math formula. . \underbrace
accents. default horizontal kern character to position \defaultskewchar^{*}
horizontal kern to position accents. . . . . . . . . \skewchar^{*}

environments.  horizontal line in array and tabular . . . . . . . .  \hline^{L}
un-box an hbox and add it to the  horizontal list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \unhbox^{*}
un-box a copy of an hbox and add it to the  horizontal list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \unhcopy^{*}
test for  horizontal mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . \ifhmode^{*}
switch to  horizontal mode from vertical mode. . . . . \leavevmode
create a box with horizontal mode material. . . . . . . . . . . . . . . . . . . \hbox^{*}
horizontal offset of a page. . . . . . . . . . . . . . . . . . \hoffset^{*}
produce a  horizontal rule. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \hrule^{*}
breakable small  horizontal skip. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \enskip
large breakable  horizontal skip. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \qquad
breakable medium  horizontal skip. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \quad
unbreakable small  horizontal space. . . . . . . . . . . . . . . . . . . . . . . . . . . . \enspace
infinitely stretchable  horizontal space. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \hfil^{*}
more infinitely stretchable  horizontal space. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \hfill^{*}
cancel infinitely stretchable  horizontal space. . . . . . . . . . . . . . . . . . . . . . . . . . . . \hfilneg^{*}
create some  horizontal space. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \hglue
add  horizontal space. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \hskip^{*}
skip  horizontal space. . . . . . . . . . . . . . . . . . . . . . . . . . . . \hspace^{L}
infinitely stretchable and shrinkable horizontal space. . . . . . . . . . . . . . . . . . . . . . . . . . . . \hss^{*}
unbreakable tiny negative math mode  horizontal space. . . . . . . . . . . . . . . . . . . . . . . \negthinspace
unbreakable small  horizontal space. . . . . . . . . . . . . . . . . . . . . . . . . . \thinspace
insert horizontal space in math mode. . . . . . . . . . . . . \mskip^{*}
define  horizontal tabs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \settabs
long  Hungarian umlaut accent (ő). . . . . . . . . . . . . . .  \H
math function:  hyperbolic cosine. . . . . . . . . . . . . . . . . . . . . . . . . . .  \cosh
math function:  hyperbolic cotangent. . . . . . . . . . . . . . . . . . . . . . .  \coth
math function:  hyperbolic sine. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sinh
math function:  hyperbolic tangent. . . . . . . . . . . . . . . . . . . . . . . . .  \tanh
discretionary  hyphen. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \-^{*}
penalty for line break after explicit  hyphen. . . . . . . . . . . . . . . . . . . . . . . . . . . . \exhyphenpenalty^{*}
penalty for line break after discretionary  hyphen. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \hyphenpenalty^{*}
default  hyphen when a font is loaded… \defaulthyphenchar^{*}
penalty if page break after  hyphenated line. . . . . . . . . . . . . . . . . . . . . . \brokenpenalty^{*}
letters. positive if hyphenating words beginning with capital . . \uchyph^{*}
current language used for  hyphenation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \language^{3}
define a new language to be used for  hyphenation. . . . . . . . . . . . . . . . . . . . . . . . . . . \newlanguage^{3}
specify language to be used for  hyphenation. . . . . . . . . . . . . . . . . . . . . . . . . . . \setlanguage^{3}
badness tolerance after  hyphenation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tolerance^{*}
word.  minimum number of characters after  hyphenation at the end of a . . \righthyphenposition^{3}
word.  minimum number of characters before  hyphenation at the start of a . . \lefthyphenposition^{3}
hyphenation character for this font. . . . . . \hyphenchar^{*}
insert words into  hyphenation dictionary. . . . . . . . . . . . . . . . . \hyphenation^{*}
badness tolerance before  hyphenation is attempted. . . . . . . . . . . . . \pretolerance^{*}
define a set of  hyphenation patterns. . . . . . . . . . . . . . . . . . . . . . \patterns^{*}
show  hyphenations of given words. . . . . . . . . . . . \showhyphens
alternative to an  \if. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \else^{*}
end of an  \if clause. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \fi^{*}
define a new  \if command. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \newif
separate cases in an  \ifcase. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \or^{*}
read, expand, then  ignore tokens until a non-space is found. \ignorespaces^{*}
math symbol: imaginary, Fraktur I (ℑ). . . . . . . . . . . . . . . . . . . . \Im
perform a \read or \write  immediately. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \immediate^{*}
post-processing.  include balanced text in DVI file for . . . . . . . \special^{*}

enable only specific  \include commands. . . . . . . . . . . . . . . . . . . \includeonly^{L}
amount by which delimiters can fail to span  included material. . . . . . . . . . . . . . \delimitershortfall^{*}
read a file unless disabled by  \includeonly. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \include^{L}
\contentsline command.  includes the section number in a . . . . . . . . \numberline^{L}
demerits for adjacent  incompatible lines. . . . . . . . . . . . . . . . . . . . . . \adjdemerits^{*}
\parindent.  increase left and right margins by . . . . . . . . . \narrower
increase line separation in math mode. . . . . . \openup
increment a counter. . . . . . . . . . . . . . . . . . . \addtocounter^{L}
increment a counter. . . . . . . . . . . . . . . . . . . . \stepcounter^{L}
increment and reference a counter. . \refstepcounter^{L}
increment the value of a length command. \addtolength^{L}
environment.  indent before the label in a list . . . . . . . . . . \itemindent^{L}
hanging  indent of a paragraph. . . . . . . . . . . . . . . . . . . . \@hangfrom^{L}
paragraph by \parindent. indent second and subsequent lines in a …. \hang
paragraphs in a list environment.  indent second and subsequent . . . . . . . \listparindent^{L}
set hanging  indentation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \hangindent^{*}
start a paragraph without  indentation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \noindent^{*}
define paragraph  indentation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \parindent^{*}
display an item without hanging  indentation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \textindent
of lines. hanging indentation changes after specified number \hangafter^{*}
indentation of display equations. . . . . . . . . \mathindent^{L}
equation.  indentation of line for displayed . . . . . \displayindent^{*}
start a paragraph  indented \parindent. . . . . . . . . . . . . . . . . . . . . . . \indent^{*}
environment. indents left margin one tab stop in tabbing . \+^{L}
amount of extra space between entries in an  index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \indexspace^{L}
second-level item in an  index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \subitem^{L}
third-level item in an  index. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \subsubitem^{L}
enable  index creation. . . . . . . . . . . . . . . . . . . . . . . . . . . . \makeindex^{L}
create an index entry. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \index^{L}
en entry in a file of  index terms. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \indexentry^{L}
math function:  inf. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \inf
math function: limit  infimum (liminf). . . . . . . . . . . . . . . . . . . . . . . . . .  \liminf
space.  infinitely stretchable and shrinkable horizontal  \hss^{*}
space.  infinitely stretchable and shrinkable vertical  \vss^{*}
infinitely stretchable horizontal space. . . . . . .  \hfil^{*}
more infinitely stretchable horizontal space. . . . . . . \hfill^{*}
cancel  infinitely stretchable horizontal space. . . . . . . \hfilneg^{*}
infinitely stretchable space. . . . . . . . . . . . . . . . . . \stretch^{L}
infinitely stretchable vertical space. . . . . . . . .  \vfil^{*}
more infinitely stretchable vertical space. . . . . . . . . \vfill^{*}
cancel  infinitely stretchable vertical space. . . . . . . . . \vfilneg^{*}
math symbol: infinity (∞). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \infty
inhibit a page break over a region. . . . . . . . . . \samepage^{L}
terminate TEX and write a format file: INITEX only. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \dump^{*}
start an \halign with \tabskip  initialized to zero. . . . . . . . . . . . . . . . . . . . . . . . . . . \ialign
begin a tabbed line in an  inner environment. . . . . . . . . . . . . . . . . . . . . . . . . \tabalign
define an  inner math subformula. . . . . . . . . . . . . . . . . . . \mathinner^{*}
finish processing input. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bye
pause for normal errors while processing TEX  input. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \errorstopmode^{*}
open a file for  input. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \openin^{*}
close an  input file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \closein^{*}
name of the principal  input file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \jobname^{*}
allocate a new  input file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newread

stop reading current  input file at the end of the current line. . . . . \endinput^{*}
character placed at the right end of an  input line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \endlinechar^{*}
verbatim line breaks in  input text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \obeylines
verbatim spaces in  input text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \obeyspaces
process  input without displaying errors. . . . . . . . . . . \batchmode^{*}
process TEX input without pausing for normal errors. \scrollmode^{*}
process  input without stopping for errors. . . . . . . \nonstopmode^{*}
end of an  insert. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \endinsert
insert a footnote mark without text… \footnotemark^{L}
completed.  insert a token after the current group is . \aftergroup^{*}
assignment command.  insert a token after the next . . . . . . . \afterassignment^{*}
insert a whole page. . . . . . . . . . . . . . . . . . . . . . \pageinsert
tokens to  insert after every \cr or nonredundant \crcr. \everycr^{*}
insert at current position in page. . . . . . . . . \midinsert
insert horizontal space in math mode. . . . . . .  \mskip^{*}
allocate a new  insert register. . . . . . . . . . . . . . . . . . . . . . . . . . . . \newinsert
insert text at the top of the page. . . . . . . . . \topinsert
\valign.  insert unaligned material in \halign or …. \noalign^{*}
insert vertical material into a paragraph. … \vadjust^{*}
tokens to  insert when a paragraph begins. . . . . . . . . . . . \everypar^{*}
tokens to  insert when a vbox begins. . . . . . . . . . . . . . . . \everyvbox^{*}
tokens to  insert when an hbox begins. . . . . . . . . . . . . . . \everyhbox^{*}
tokens to  insert when display math begins. . . . . . \everydisplay^{*}
tokens to  insert when math in text begins. . . . . . . . . . \everymath^{*}
tokens to  insert when the job begins. . . . . . . . . . . . . . . . . \everyjob^{*}
insert words into hyphenation dictionary. \hyphenation^{*}
define where text will be  inserted when the page is formatted. . . . . . . . \insert^{*}
insertion class for footnote inserts. . . . . . . . . . \footins
insertion class for inserts at the top of a page.  \topins
flush  insertions and eject to a new page. . . . . \dosupereject
flush all  insertions and eject to a new page. . . . . . . \supereject
sum of penalties for split  insertions on the page. . . . . . . . . . . . . \insertpenalties^{*}
penalty for  insertions that are split. . . . . . . . . . . . \floatingpenalty^{*}
insertion class for footnote  inserts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footins
insertion class for  inserts at the top of a page. . . . . . . . . . . . . . . . . \topins
select one of several entries determined by an  integer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ifcase^{*}
test for odd  integer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ifodd^{*}
test two  integers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ifnum^{*}
math operator:  integral (∫). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \int
math operator: contour  integral (∮). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \oint
math operator: small  integral (∫). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \smallint
enable BibTEX bibliography  interface. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bibliography^{L}
footnotes.  interline penalty for …. \interfootnotelinepenalty^{L}
box. suppress interline space before next vertical \nointerlineskip
set @ alphabetic to access  internal commands. . . . . . . . . . . . . . . . . . . . \makeatletter^{L}
set @ non-alphabetic to hide  internal commands. . . . . . . . . . . . . . . . . . . . . \makeatother^{L}
test for an  internal mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ifinner^{*}
long arrows.  internal Plain TEX command to piece together \joinrel
math text symbols.  internal Plain TEX operation to define . . \mathhexbox
internal Plain TEX space used for centering. \centering
math operator:  intersection or cap (∩). . . . . . . . . . . . . . . . . . . . . .  \cap
generate an  in-text citation of a reference. . . . . . . . . . . . . . .  \cite^{L}
use  in-text style for math. . . . . . . . . . . . . . . . . . . . . \textstyle^{*}

math symbol:  iota (ι). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \iota
select bold math  italic and symbol fonts. . . . . . . . . . . . . . . . . . . . . \boldmath^{L}
unselect bold math  italic and symbol fonts. . . . . . . . . . . . . . . . . . \unboldmath^{L}
italic correction. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \/^{*}
8 point  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \egtit^{L}
8 point math  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \egtmi^{L}
11 point  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvit^{L}
11 point math  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvmi^{L}
5 point math  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \fivei
5 point math  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \fivmi^{L}
14 point math  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \frtnmi^{L}
select  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \it
math mode  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \mit
9 point  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ninit^{L}
9 point math  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ninmi^{L}
7 point math  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \seveni
7 point  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sevit^{L}
7 point math  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sevmi^{L}
6 point math  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sixmi^{L}
17 point math  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \svtnmi^{L}
10 point math  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \teni
10 point text  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tenit
10 point math  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tenmi^{L}
12 point  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twlit^{L}
12 point math  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twlmi^{L}
20 point math  italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twtymi^{L}
italic font family. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \itfam
display an item. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \item
space between a label and text of a list  item. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \labelsep^{L}
math spacing of a closing delimiter for the next  item.  select . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathclose^{*}
math spacing of an opening delimiter for the next  item.  select . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathopen^{*}
space between first list  item and preceding paragraph. . . . . . . . . . . . . . \topsep^{L}
current value of the third-level  item counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . \theenumiii^{L}
current value of the second-level  item counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \theenumii^{L}
current value of the first-level  item counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \theenumi^{L}
current value of the fourth-level  item counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \theenumiv^{L}
current value of the sixth-level  item counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \theenumvi^{L}
current value of the fifth-level  item counter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \theenumv^{L}
space between paragraphs within an  item in a list environment. . . . . . . . . . . . . . . . . . \parsep^{L}
second-level  item in an index. . . . . . . . . . . . . . . . . . . . . . . . . . . . \subitem^{L}
third-level  item in an index. . . . . . . . . . . . . . . . . . . . . . . . . \subsubitem^{L}
create  item label for a list environment. . . . . . . . . . \makelabel^{L}
place balanced text into a mark item on the current list. . . . . . . . . . . . . . . . . . . . . \mark^{*}
append a discretionary  item to the current list. . . . . . . . . . . . . . . \discretionary^{*}
reference a BibTEX  item without citation. . . . . . . . . . . . . . . . . . . . . . . \nocite^{L}
display an  item without hanging indentation. . . . . . . \textindent
penalty between list  items. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \@itempenalty^{L}
mark denoting third-level  items. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \labelitemiii^{L}
mark denoting second-level  items. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \labelitemii^{L}
mark denoting first-level  items. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \labelitemi^{L}
mark denoting fourth-level  items. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \labelitemiv^{L}
mark denoting sixth-level  items. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \labelitemvi^{L}
mark denoting fifth-level  items. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \labelitemv^{L}

create a box with a single column of  items. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \shortstack^{L}
space between successive  items in a list environment. . . . . . . . . . . . . . . . . \itemsep^{L}
maximum boxed  items shown at a given depth. . . . . . . \showboxbreadth^{*}
associate a counter with an  item-type command. . . . . . . . . . . . . . . . . . . . . \usecounter^{L}
math symbol: dotless  j (ȷ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \jmath
dotless  j letter (ȷ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \j
math symbol:  join (⋈). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Join^{L}
math  join operator: large V (⋁). . . . . . . . . . . . . . . . . . \bigvee
math binary operator: join or V (∨). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vee
suppress bottom  justification of page text. . . . . . . . . . . . . . \raggedbottom
suppress right  justification of paragraph lines. . . . . . . . . \raggedright
suppress right  justification of typewriter font lines. . \ttraggedright
justify page bottoms to the same height. \flushbottom^{L}
justify page bottoms to the same height. \normalbottom
height.  justify page bottoms to their natural . \raggedbottom^{L}
math Greek letter:  kappa (κ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \kappa
math function:  ker. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ker
kern a given distance. . . . . . . . . . . . . . . . . . . . . . .  \kern^{*}
default horizontal kern character to position accents. \defaultskewchar^{*}
kern in math mode. . . . . . . . . . . . . . . . . . . . . . . . .  \mkern^{*}
remove a  kern just added to the current list. . . . . . . . . . \unkern^{*}
fetch last  kern off the current list. . . . . . . . . . . . . . . . . . . . \lastkern^{*}
horizontal  kern to position accents. . . . . . . . . . . . . . . . . . . . \skewchar^{*}
define  kerning around math in text. . . . . . . . . . \mathsurround^{*}
math symbol: small script  L (ℓ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ell
Polish letter: slashed  L (ł). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \l
Polish letter: upper-case slashed  L (Ł). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \L
define a cross reference  label. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \label^{L}
page number of a cross reference  label. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pageref^{L}
refer to a cross reference label. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ref^{L}
space between a  label and text of a list item. . . . . . . . . . . . . . . . \labelsep^{L}
create item  label for a list environment. . . . . . . . . . . . . . . \makelabel^{L}
indent before the  label in a list environment. . . . . . . . . . . . . . . \itemindent^{L}
label width in a list environment. . . . . . . . \labelwidth^{L}
generate a matrix  labeled on rows and columns. . . . . . . . . . \bordermatrix
start a  labeled paragraph. . . . . . . . . . . . . . . . . . . . . . . . \paragraph^{L}
start a  labeled sub-level paragraph. . . . . . . . . . . \subparagraph^{L}
math Greek letter:  lambda (λ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lambda
math Greek letter: capital  lambda (Λ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \Lambda
define a new  language to be used for hyphenation… \newlanguage^{3}
specify  language to be used for hyphenation… \setlanguage^{3}
current  language used for hyphenation. . . . . . . . . . . . . \language^{3}
vertical skip a  large amount. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigskip^{L}
large breakable horizontal skip. . . . . . . . . . . . . .  \qquad
math operator:  large cap (⋂). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigcap
math operator:  large circle (◯). . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigcirc
math operator:  large circle with dot (⨀). . . . . . . . . . . . . . . . . . . \bigodot
math operator:  large circle with plus (⨁). . . . . . . . . . . . . . . . . . \bigoplus
math operator:  large circle with times (⨂). . . . . . . . . . . . . . . \bigotimes
math operator:  large co-product (∐). . . . . . . . . . . . . . . . . . . . . . . \coprod
math operator: large cup (⋃). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigcup
math operator:  large down triangle (▽). . . . . . . . . . . \bigtriangledown
define a  large math operator. . . . . . . . . . . . . . . . . . . . . . . . \mathop^{*}

restore default limit placement on  large math operators. . . . . . . . . . . . . . . . . \displaylimits^{*}
math operator: large product (∏). . . . . . . . . . . . . . . . . . . . . . . . . . \prod
math operator:  large square cup (⨆). . . . . . . . . . . . . . . . . . . . . . . \bigsqcup
math operator: large sum (∑). . . . . . . . . . . . . . . . . . . . . . . . . . . . . \sum
math operator:  large U plus (⨄). . . . . . . . . . . . . . . . . . . . . . . . . . . \biguplus
math operator:  large up triangle (△). . . . . . . . . . . . . . . . \bigtriangleup
math join operator: large V (⋁). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigvee
large vertical break. . . . . . . . . . . . . . . . . . . . . . . . . \bigskip
large vertical space or a good page break… \bigbreak
math meet operator:  large wedge (⋀). . . . . . . . . . . . . . . . . . . . . . . . . . . \bigwedge
select font four steps  larger than normal size. . . . . . . . . . . . . . . . . . . . .  \huge^{L}
select font one step  larger than normal size. . . . . . . . . . . . . . . . . . . . .  \large^{L}
select font three steps  larger than normal size. . . . . . . . . . . . . . . . . . . . .  \LARGE^{L}
select font two steps  larger than normal size. . . . . . . . . . . . . . . . . . . . .  \Large^{L}
select  largest available font. . . . . . . . . . . . . . . . . . . . . . .  \Huge^{L}
largest permissible dimension. . . . . . . . . . . . . . \maxdimen
fetch  last box off the current list. . . . . . . . . . . . . . . . . \lastbox^{*}
depth of the  last box on the vertical list. . . . . . . . . . . . . . . \prevdepth^{*}
mark text  last encountered in a split box. . . . . . . . \splitbotmark^{*}
mark text  last encountered on page just boxed. . . . . . . . \botmark^{*}
fetch  last kern off the current list. . . . . . . . . . . . . . . . \lastkern^{*}
space at the end of the  last line of a paragraph. . . . . . . . . . . . . . . . . \parfillskip^{*}
number of lines in the  last paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \prevgraf^{*}
fetch  last penalty off the current list. . . . . . . . . \lastpenalty^{*}
number of output routine calls since  last \shipout. . . . . . . . . . . . . . . . . . . . . . . . . . . \deadcycles^{*}
fetch  last skip off the current list. . . . . . . . . . . . . . . . \lastskip^{*}
remove  last skip on the list. . . . . . . . . . . . . . . . . \removelastskip
terminate L^{A}TEX and flush the final page. . . . . . . . . . . . . . \stop^{L}
L^{A}TEX logo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \LaTeX^{L}
8 point  L^{A}TEX symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \egtly^{L}
11 point  L^{A}TEX symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \elvly^{L}
5 point  L^{A}TEX symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \fivly^{L}
14 point  L^{A}TEX symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . \frtnly^{L}
9 point  L^{A}TEX symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \ninly^{L}
7 point  L^{A}TEX symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \sevly^{L}
6 point  L^{A}TEX symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \sixly^{L}
17 point  L^{A}TEX symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . \svtnly^{L}
10 point  L^{A}TEX symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \tenly^{L}
12 point  L^{A}TEX symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \twlly^{L}
20 point  L^{A}TEX symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . \twtyly^{L}
an ellipsis, equivalent to \ldots in math mode. . . . . . . . . . . . . . . . . . . . . . \dots
math symbol:  leads to (⇝). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leadsto^{L}
create a zero-width box with text to the left. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \llap
add a box to the vertical list shifted  left. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \moveleft^{*}
flush lines  left. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \raggedright^{L}
math symbol: long  left and right arrow (⟷). . . . . \longleftrightarrow
increase  left and right margins by \parindent. . . . . . \narrower
set headings for  left and right pages. . . . . . . . . . . . . . . . . . . . . . . . \markboth^{L}
math delimiter: left angle bracket (⟨). . . . . . . . . . . . . . . . . . . . . . . \langle
fill a space with a  left arrow. . . . . . . . . . . . . . . . . . . . . . . . . . . . \leftarrowfill
math symbol: hook  left arrow (↩). . . . . . . . . . . . . . . . . . . . . . . \hookleftarrow
math symbol:  left arrow (←). . . . . . . . . . . . . . . . . . . . . . . . . . . . \leftarrow
math symbol: long  left arrow (⟵). . . . . . . . . . . . . . . . . . . . . . \longleftarrow

left arrow over a math formula. . . . . . . \overleftarrow
generate a matrix with a left brace delimiter. . . . . . . . . . . . . . . . . . . . . . . . . \cases
math delimiter:  left bracket ([). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lbrack
distance  left by \' command between tabbing fields. \tabbingsep^{L}
math delimiter: left ceiling bracket (⌈). . . . . . . . . . . . . . . . . . . . . . \lceil
math delimiter:  left curly brace ({). . . . . . . . . . . . . . . . . . . . . . . . . \lbrace
begin a new math list with a left delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \left^{*}
math symbol:  left double arrow (⇐). . . . . . . . . . . . . . . . . . . . \Leftarrow
math symbol: long  left double arrow (⟸). . . . . . . . . . . . . . \Longleftarrow
end flush  left environment. . . . . . . . . . . . . . . . . . . . . . \endflushleft^{L}
begin flush  left environment. . . . . . . . . . . . . . . . . . . . . . . . . . \flushleft^{L}
left equation number. . . . . . . . . . . . . . . . . . . . . . .  \leqno^{*}
align a stack of equations with  left equation numbers. . . . . . . . . . . . . . . . . . . \leqalignno
math delimiter:  left floor bracket (⌊). . . . . . . . . . . . . . . . . . . . . . . . \lfloor
math delimiter:  left group ({). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lgroup
math symbol:  left half diamond (⊲). . . . . . . . . . . . . . . . . . . . . . .  \lhd^{L}
math symbol: underlined  left half diamond (⊴). . . . . . . . . . . . . . . . . . . . . . .  \unlhd^{L}
left hand margin on even pages. . . . . \evensidemargin^{L}
left hand margin on odd pages. . . . . . . \oddsidemargin^{L}
math symbol:  left harpoon down (↽). . . . . . . . . . . . \leftharpoondown
math symbol: right  left harpoon (⇀ ↽). . . . . . . . . . . . . . . . \rightleftharpoons
math symbol:  left harpoon up (↼). . . . . . . . . . . . . . . . . \leftharpoonup
math symbol: left hook ( ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lhook
set a formula flush  left in display style. . . . . . . . . . . . . . . . . . . . . . . . . \lefteqn^{L}
width of  left margin in fifth-level list. . . . . . . . . . . . \leftmarginv^{L}
width of  left margin in first-level list. . . . . . . . . . . . \leftmargini^{L}
width of  left margin in fourth-level list. . . . . . . . . \leftmarginiv^{L}
width of  left margin in second-level list. . . . . . . . . \leftmarginii^{L}
width of  left margin in sixth-level list. . . . . . . . . . \leftmarginvi^{L}
width of  left margin in third-level list. . . . . . . . . \leftmarginiii^{L}
left margin of a list environment. . . . . . . . . \leftmargin^{L}
environment.  unindents  left margin one tab stop in tabbing . . . . . . . . .  \-^{L}
environment.  indents  left margin one tab stop in tabbing . . . . . . . . .  \+^{L}
2.5-line  left math delimiter. . . . . . . . . . . . . . . . . . . . . . . . .  \Biggl
2-line  left math delimiter. . . . . . . . . . . . . . . . . . . . . . . . .  \biggl
1.5-line  left math delimiter. . . . . . . . . . . . . . . . . . . . . . . . .  \Bigl
1-line  left math delimiter. . . . . . . . . . . . . . . . . . . . . . . . .  \bigl
math symbol:  left moustache ({). . . . . . . . . . . . . . . . . . . . . . \lmoustache
space to the  left of a paragraph. . . . . . . . . . . . . . . . . . . . . . . . . \leftskip^{*}
flush text  left on a line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leftline
produces current  left page heading. . . . . . . . . . . . . . . . . . . . . . . . . . \leftmark^{L}
lower  left piece of a horizontal brace. . . . . . . . . . . . . . \braceld
upper  left piece of a horizontal brace. . . . . . . . . . . . . . \bracelu
math symbol: left quote (‘). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lq
math operator:  left triangle (◁). . . . . . . . . . . . . . . . . . . . . . . \triangleleft
math symbol:  left-right arrow (↔). . . . . . . . . . . . . . . . \leftrightarrow
math symbol:  left-right double arrow (⇔). . . . . . . . . \Leftrightarrow
math symbol: long  left-right double arrow (⟺). . \Longleftrightarrow
increment the value of a  length command. . . . . . . . . . . . . . . . . . . . . . . \addtolength^{L}
define a new  length command. . . . . . . . . . . . . . . . . . . . . . . . . . \newlength^{L}
set a  length command. . . . . . . . . . . . . . . . . . . . . . . . . . \setlength^{L}
set a  length command to width of text. . . . . . . . \settowidth^{L}

length of text preceding a display… \predisplaysize^{*}
math relation: much  less (≪). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ll
math relation:  less or equal (≤). . . . . . . . . . . . . . . . . . . . . . . . . . .  \le
math relation:  less or equal (≤). . . . . . . . . . . . . . . . . . . . . . . . . . .  \leq
display counter as lower-case  letter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \alph^{L}
display counter as upper-case  letter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Alph^{L}
Scandinavian  letter: a with circle (å). . . . . . . . . . . . . . . . . . . . .  \aa
math Greek  letter: alpha (α). . . . . . . . . . . . . . . . . . . . . . . . . . . .  \alpha
math Greek  letter: beta (β). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \beta
Scandinavian  letter: capital A with circle (Å). . . . . . . . . . . . .  \AA
math Greek  letter: capital delta (Δ). . . . . . . . . . . . . . . . . . . .  \Delta
math Greek  letter: capital gamma (Γ). . . . . . . . . . . . . . . . . .  \Gamma
math Greek  letter: capital lambda (Λ). . . . . . . . . . . . . . . . . . \Lambda
Norwegian  letter: capital O with slash (Ø). . . . . . . . . . . . .  \O
math Greek  letter: capital omega (Ω). . . . . . . . . . . . . . . . . . .  \Omega
math Greek  letter: capital phi (Φ). . . . . . . . . . . . . . . . . . . . . .  \Phi
math Greek  letter: capital pi (Π). . . . . . . . . . . . . . . . . . . . . . .  \Pi
math Greek  letter: capital psi (Ψ). . . . . . . . . . . . . . . . . . . . . .  \Psi
math Greek  letter: capital sigma (Σ). . . . . . . . . . . . . . . . . . . .  \Sigma
math Greek  letter: capital theta (Θ). . . . . . . . . . . . . . . . . . . .  \Theta
math Greek  letter: capital upsilon (Υ). . . . . . . . . . . . . . . . . . \Upsilon
math Greek  letter: capital xi (Ξ). . . . . . . . . . . . . . . . . . . . . . . .  \Xi
math Greek  letter: chi (χ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \chi
math Greek  letter: delta (δ). . . . . . . . . . . . . . . . . . . . . . . . . . . .  \delta
math Greek  letter: epsilon (ϵ). . . . . . . . . . . . . . . . . . . . . . . . . . . \epsilon
math Greek  letter: eta (η). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \eta
math Greek  letter: gamma (γ). . . . . . . . . . . . . . . . . . . . . . . . . .  \gamma
dotless i  letter (ı). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \i
dotless j  letter (ȷ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \j
math Greek  letter: kappa (κ). . . . . . . . . . . . . . . . . . . . . . . . . . .  \kappa
math Greek  letter: lambda (λ). . . . . . . . . . . . . . . . . . . . . . . . . . \lambda
math Greek  letter: mu (μ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \mu
math Greek  letter: nu (ν). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \nu
Norwegian  letter: o with slash (ø). . . . . . . . . . . . . . . . . . . . .  \o
math Greek  letter: omega (ω). . . . . . . . . . . . . . . . . . . . . . . . . . .  \omega
math Greek  letter: phi (ϕ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \phi
math Greek  letter: pi (π). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \pi
math Greek  letter: psi (ψ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \psi
math Greek  letter: rho (ρ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \rho
German  letter: sharp s (ß). . . . . . . . . . . . . . . . . . . . . . . . . .  \ss
math Greek  letter: sigma (σ). . . . . . . . . . . . . . . . . . . . . . . . . . .  \sigma
Polish  letter: slashed L (ł). . . . . . . . . . . . . . . . . . . . . . . . .  \l
math Greek  letter: tau (τ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tau
math Greek  letter: theta (θ). . . . . . . . . . . . . . . . . . . . . . . . . . . .  \theta
Polish  letter: upper-case slashed L (Ł). . . . . . . . . . . .  \L
math Greek  letter: upsilon (υ). . . . . . . . . . . . . . . . . . . . . . . . . . \upsilon
math Greek  letter: variant epsilon (ε). . . . . . . . . . . . . . . . \varepsilon
math Greek  letter: variant phi (φ). . . . . . . . . . . . . . . . . . . . . . \varphi
math Greek  letter: variant pi (ϖ). . . . . . . . . . . . . . . . . . . . . . .  \varpi
math Greek  letter: variant rho (ϱ). . . . . . . . . . . . . . . . . . . . . .  \varrho
math Greek  letter: variant sigma (ς). . . . . . . . . . . . . . . . . . . \varsigma
math Greek  letter: variant theta (ϑ). . . . . . . . . . . . . . . . . . . . \vartheta
math Greek  letter: xi (ξ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \xi

math Greek  letter: zeta (ζ). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \zeta
if hyphenating words beginning with capital  letters.  positive . . . . . . . . . . . . . . . . . . . . . . . . . \uchyph^{*}
math mode calligraphic  letters font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \cal
math function:  lg. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \lg
ligature digraph symbol ae (æ). . . . . . . . . . . . .  \ae
ligature digraph symbol capital AE (Æ)….  \AE
ligature digraph symbol capital OE (Œ)….  \OE
ligature digraph symbol oe (œ). . . . . . . . . . . . .  \oe
math function: lim. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lim
limit for bad hbox errors. . . . . . . . . . . . . . . . . . . \hbadness^{*}
limit for bad vbox errors. . . . . . . . . . . . . . . . . . . \vbadness^{*}
math function:  limit infimum (liminf). . . . . . . . . . . . . . . . . . . . . \liminf
operators. restore default limit placement on large math . . . . . . . \displaylimits^{*}
math function:  limit supremum (limsup). . . . . . . . . . . . . . . . . . . \limsup
place math  limits above and below math operators. . . . . \limits^{*}
penalty if page break after hyphenated  line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \brokenpenalty^{*}
center text in a  line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \centerline
current input file at the end of the current line. stop reading . . . . . . . . . . . . . . . . . . . . . . \endinput^{*}
character placed at the right end of an input  line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \endlinechar^{*}
demerits for a penultimate broken  line. . . . . . . . . . . . . . . . . . . . . . . . . . . \finalhyphendemerits^{*}
page foot  line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footline
page head  line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \headline
flush text left on a  line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leftline
width of the current  line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \linewidth^{L}
flush right text on a  line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \rightline
create a rule or  line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \rule^{L}
require strict  line and page breaks. . . . . . . . . . . . . . . . . . . . . . . .  \fussy^{L}
accept looser  line and page breaks. . . . . . . . . . . . . . . . . . . . . . . . \sloppy^{L}
penalty for creating a widow  line at top of page. . . . . . . . . . . . . . . . . . . . \widowpenalty^{*}
penalty for creating a widow  line before a display. . . . . . . . . . \displaywidowpenalty^{*}
allow a  line break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \allowbreak
force a line break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \break
encourage a  line break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \linebreak^{L}
suppress a  line break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \nolinebreak^{L}
permit a line break after a slash. . . . . . . . . . . . . . . . . . . . . \slash
penalty for  line break after binary operation. . . . . . \binoppenalty^{*}
penalty for line break after discretionary hyphen. \hyphenpenalty^{*}
penalty for  line break after explicit hyphen. … \exhyphenpenalty^{*}
penalty for  line break after math relation. . . . . . . . . . . \relpenalty^{*}
force a line break in a paragraph. . . . . . . . . . . . . . . . . . . \\^{L}
force a  line break in a paragraph. . . . . . . . . . . . . . . . . . . \newline^{L}
verbatim  line breaks in input text. . . . . . . . . . . . . . . . . . \obeylines
10 point line font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tenln^{L}
10 point wide  line font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tenlnw^{L}
indentation of  line for displayed equation. . . . . . . . . . . \displayindent^{*}
read a line from a file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \read^{*}
amount added to badness of every  line in a paragraph. . . . . . . . . . . . . . . . . . . . . \linepenalty^{*}
line in a picture environment. . . . . . . . . . . . . . .  \line^{L}
character that starts a new output  line in a write statement. . . . . . . . . . . . . . . \newlinechar^{*}
begin a tabbed  line in an inner environment. . . . . . . . . . . . . . . \tabalign
begin a tabbed  line in an outer environment. . . . . . . . . . . . . . . .  \+
horizontal  line in array and tabular environments. . . . .  \hline^{L}
vertical  line in array and tabular environments. . . . .  \vline^{L}

multicolumn  line in array or tabular environment. . . . . . . .  \cline^{L}
end a  line in \halign aligned text. . . . . . . . . . . . . . . .  \cr^{*}
discard current  line in tabbing environment. . . . . . . . . . . . . . . .  \kill^{L}
pause after each  line is read from a file. . . . . . . . . . . . . . . . . . . . . . \pausing^{*}
space at the end of the last  line of a paragraph. . . . . . . . . . . . . . . . . . . . . \parfillskip^{*}
penalty if page break after first  line of paragraph. . . . . . . . . . . . . . . . . . . . . . . \clubpenalty^{*}
prohibit a  line or page break. . . . . . . . . . . . . . . . . . . . . . . . . . \nobreak
specify penalty for a  line or page break. . . . . . . . . . . . . . . . . . . . . . . . . . \penalty^{*}
draw a  line over a formula. . . . . . . . . . . . . . . . . . . . . . . . . \overline^{*}
ensure footnote  line separation. . . . . . . . . . . . . . . . . . . . . . . . . . . . \footstrut
increase  line separation in math mode. . . . . . . . . . . . . . . \openup
vertical strut to preserve line spacing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \strut
set  line spacing to normal values. . . . . . \normalbaselines
draw a  line under a formula. . . . . . . . . . . . . . . . . . . . . . \underline^{*}
create a box of current line width. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \line
show  line-break calculations. . . . . . . . . . \tracingparagraphs^{*}
extra space above displays following short  lines. . . . . . . . . . . . . . . . . . . . . . . \abovedisplayshortskip^{*}
demerits for adjacent incompatible  lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \adjdemerits^{*}
normal space between  lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \baselineskip^{*}
extra space just below displays following short  lines. . . . . . . . . . . . . . . . . . . . . . . \belowdisplayshortskip^{*}
declaration to center  lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \centering^{L}
demerits for consecutive broken  lines. . . . . . . . . . . . . . . . . . . . . . . . . \doublehyphendemerits^{*}
extra space in badly-stretched  lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \emergencystretch^{3}
indentation changes after specified number of lines. hanging . . . . . . . . . . . . . . . . . . . . . . . . \hangafter^{*}
additional penalty for page break between  lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \interlinepenalty^{*}
turn off extra space between  lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \offinterlineskip
suppress right justification of paragraph  lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \raggedright
suppress right justification of typewriter font  lines. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ttraggedright
thick lines for  lines and circles. . . . . . . . . . . . . . . . . . . . . . . . . \thicklines^{L}
thin lines for  lines and circles. . . . . . . . . . . . . . . . . . . . . . . . . . \thinlines^{L}
thick  lines for lines and circles. . . . . . . . . . . . . . . . . \thicklines^{L}
thin  lines for lines and circles. . . . . . . . . . . . . . . . . . \thinlines^{L}
extra space between  lines if \baselineskip isn’t enough. . . . . . . . \lineskip^{*}
force a change to the number of  lines in a paragraph. . . . . . . . . . . . . . . . . . . . . . \looseness^{*}
indent second and subsequent lines in a paragraph by \parindent. . . . . . . . \hang
minimum space between  lines in a vertical list. . . . . . . . . . . . . . . . \lineskiplimit^{*}
set width of  lines in picture environment. . . . . . . . . \linethickness^{L}
number of  lines in the last paragraph. . . . . . . . . . . . . . . . . \prevgraf^{*}
flush  lines left. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \raggedright^{L}
penalty for breaking between  lines of a display. . . . . . . . . \interdisplaylinepenalty
penalty for breaking between  lines of a footnote. . . . . . \interfootnotelinepenalty
an error message. number of lines of context to be displayed in \errorcontextlines^{3}
flush  lines right. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \raggedleft^{L}
value of normal  \lineskip. . . . . . . . . . . . . . . . . . . . . . . . . \normallineskip

           value of normal  \lineskiplimit. . . . . . . . . . . . . \normallineskiplimit

penalty at the beginning of a  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \@beginparpenalty^{L}
append a discretionary item to the current  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \discretionary^{*}
terminate a \csname token  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \endcsname^{*}
penalty at the end of a  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \@endparpenalty^{L}
fetch last box off the current  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lastbox^{*}
fetch last kern off the current  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lastkern^{*}
fetch last penalty off the current  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lastpenalty^{*}
fetch last skip off the current  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lastskip^{*}

width of left margin in third-level  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leftmarginiii^{L}
width of left margin in second-level  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leftmarginii^{L}
width of left margin in first-level  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leftmargini^{L}
width of left margin in fourth-level  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leftmarginiv^{L}
width of left margin in sixth-level  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leftmarginvi^{L}
width of left margin in fifth-level  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leftmarginv^{L}
minimum space between lines in a vertical  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lineskiplimit^{*}
balanced text into a mark item on the current list. place . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mark^{*}
depth of the last box on the vertical  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \prevdepth^{*}
remove last skip on the  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \removelastskip
right margin in a  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \rightmargin^{L}
un-box an hbox and add it to the horizontal  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \unhbox^{*}
a copy of an hbox and add it to the horizontal list. un-box . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \unhcopy^{*}
remove a kern just added to the current  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \unkern^{*}
remove a penalty just added to the current  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \unpenalty^{*}
remove a skip just added to the current  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \unskip^{*}
un-box a vbox and add it to the vertical  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \unvbox^{*}
a copy of a vbox and add it to the vertical list. un-box . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \unvcopy^{*}
create a centered vbox in a math  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vcenter^{*}
place a footnote in a caption or other vertical  list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vfootnote
give a token  list defining the semantics of a token. . . . . . . \meaning^{*}
indent before the label in a  list environment. . . . . . . . . . . . . . . . . . . . . . . . . \itemindent^{L}
start an entry in a list environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . \item^{L}
space between successive items in a  list environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . \itemsep^{L}
label width in a  list environment. . . . . . . . . . . . . . . . . . . . . . . . . \labelwidth^{L}
left margin of a  list environment. . . . . . . . . . . . . . . . . . . . . . . . . \leftmargin^{L}
indent second and subsequent paragraphs in a  list environment. . . . . . . . . . . . . . . . . . . . . \listparindent^{L}
create item label for a  list environment. . . . . . . . . . . . . . . . . . . . . . . . . . \makelabel^{L}
space between paragraphs within an item in a  list environment. . . . . . . . . . . . . . . . . . . . . . . . . . . . \parsep^{L}
space between a label and text of a  list item. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \labelsep^{L}
space between first  list item and preceding paragraph. . . . . . . . . . \topsep^{L}
penalty between  list items. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \@itempenalty^{L}
generate a  list of figures. . . . . . . . . . . . . . . . . . . . . . . . \listoffigures^{L}
generate a  list of tables. . . . . . . . . . . . . . . . . . . . . . . . . . \listoftables^{L}
produce a vertical  list of the body of a page. . . . . . . . . . . . . \pagecontents
an empty list of tokens. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \empty
adds an entry to the specified  list or table. . . . . . . . . . . . . . . . . . . . . . . \addcontentsline^{L}
use a token list register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \toks^{*}
define a name for a token  list register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \toksdef^{*}
add a box to the vertical  list shifted left. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \moveleft^{*}
add a box to the vertical  list shifted right. . . . . . . . . . . . . . . . . . . . . . . . . . \moveright^{*}
write a token  list to a file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \write^{*}
begin a new math  list with a left delimiter. . . . . . . . . . . . . . . . . . . .  \left^{*}
end a math  list with a right delimiter. . . . . . . . . . . . . . . . . . .  \right^{*}
show the current  lists. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \showlists^{*}
math function: ln. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ln
load a font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \font^{*}
load a font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \load^{L}
default hyphen when a font is  loaded. . . . . . . . . . . . . . . . . . . . . . . . . . \defaulthyphenchar^{*}
math function:  log. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \log
write to the  log file only. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \wlog
math operator:  logical and (∧). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \land
math operator:  logical not, hook (¬). . . . . . . . . . . . . . . . . . . . . . .  \lnot

math symbol:  logical or (∨). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \lor
L^{A}TEX  logo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \LaTeX^{L}
“TEX”  logo. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \TeX
begin a  loop. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \loop
end of a  \loop body. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \repeat
accept  looser line and page breaks. . . . . . . . . . . . . . . . . \sloppy^{L}
current year of our  Lord. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \year^{*}
three  low dots in math mode (…). . . . . . . . . . . . . . . .  \ldots
lower a box a given distance. . . . . . . . . . . . . . . .  \lower^{*}
lower dot math symbol with special spacing.  \ldotp
lower left piece of a horizontal brace. . . . . . . . \braceld
lower right piece of a horizontal brace. . . . . . \bracerd
translate balanced text to  lower-case. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lowercase^{*}
lower-case code for a character. . . . . . . . . . . . .  \lccode^{*}
display counter as  lower-case letter. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \alph^{L}
display counter as  lower-case Roman numerals. . . . . . . . . . . . . . . .  \roman^{L}
convert a number to  lower-case Roman numerals. . . . . . . . . . . \romannumeral^{*}
define a  macro. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \def^{*}
following macro must not be called from another  macro.  the . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \outer^{*}
the following  macro definition or register setting is global.  \global^{*}
define a  macro globally. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \gdef^{*}
parameters.  the following  macro may have multiple paragraphs as ….  \long^{*}
the following  macro must not be called from another macro.  \outer^{*}
footnotes from text.  macro to draw the rule separating …. \footnoterule^{L}
define a  macro with expanded replacement text. . . . .  \edef^{*}
define a global  macro with expanded replacement text. . . . .  \xdef^{*}
macron or bar accent (ō). . . . . . . . . . . . . . . . . . .  \=
show  macros as they are expanded. . . . . . . . \tracingmacros^{*}
define font  magnification. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \magstep
use 11 point  magnification. . . . . . . . . . . . . . . . . . . . . . . . . . \magstephalf
set the  magnification for the document. . . . . . \magnification
magnification ratio times 1000. . . . . . . . . . . . . .  \mag^{*}
maximum depth of boxes on  main pages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \maxdepth^{*}
space between  main text and footnotes. . . . . . . . . . . . . . . . . . . . \footins^{L}
math symbol:  maps to char ( ). . . . . . . . . . . . . . . . . . . . . . . . . \mapstochar
math symbol: long  maps to (⟼). . . . . . . . . . . . . . . . . . . . . . . . . . . \longmapsto
math symbol:  maps to (↦). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mapsto
right  margin in a list. . . . . . . . . . . . . . . . . . . . . . . . \rightmargin^{L}
width of left  margin in fifth-level list. . . . . . . . . . . . . . . . \leftmarginv^{L}
width of left  margin in first-level list. . . . . . . . . . . . . . . . \leftmargini^{L}
width of left  margin in fourth-level list. . . . . . . . . . . . . \leftmarginiv^{L}
width of left  margin in second-level list. . . . . . . . . . . . . \leftmarginii^{L}
width of left  margin in sixth-level list. . . . . . . . . . . . . . \leftmarginvi^{L}
width of left  margin in third-level list. . . . . . . . . . . . . \leftmarginiii^{L}
left  margin of a list environment. . . . . . . . . . . . . \leftmargin^{L}
left hand  margin on even pages. . . . . . . . . . . . . . . \evensidemargin^{L}
left hand  margin on odd pages. . . . . . . . . . . . . . . . . \oddsidemargin^{L}
unindents left  margin one tab stop in tabbing environment.  \-^{L}
indents left  margin one tab stop in tabbing environment.  \+^{L}
produce a  marginal note. . . . . . . . . . . . . . . . . . . . . . . . . . . . \marginpar^{L}
distance between  marginal note and text. . . . . . . . . . . . . . . . \marginparsep^{L}
vertical space between  marginal notes. . . . . . . . . . . . . . . . . . . . . . . \marginparpush^{L}
width of  marginal notes. . . . . . . . . . . . . . . . . . . . . \marginparwidth^{L}

page.  put  marginal notes on normal side of . . \normalmarginpar^{L}
page.  put  marginal notes on opposite side of \reversemarginpar^{L}
increase left and right  margins by \parindent. . . . . . . . . . . . . . . . . . . . \narrower
produce footnote text without a  mark. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footnotetext^{L}
mark denoting fifth-level items. . . . . . . . . . \labelitemv^{L}
mark denoting first-level items. . . . . . . . . . \labelitemi^{L}
mark denoting fourth-level items. . . . . . . \labelitemiv^{L}
mark denoting second-level items. . . . . . . \labelitemii^{L}
mark denoting sixth-level items. . . . . . . . \labelitemvi^{L}
mark denoting third-level items. . . . . . . \labelitemiii^{L}
place balanced text into a  mark item on the current list. . . . . . . . . . . . . . .  \mark^{*}
pound, hatch  mark, sharp sign, octothorpe (#). . . . . . . . . . .  \#
box.  mark text first encountered in a split \splitfirstmark^{*}
boxed.  mark text first encountered on page just . . \firstmark^{*}
mark text last encountered in a split box. \splitbotmark^{*}
boxed.  mark text last encountered on page just …. \botmark^{*}
subdivision.  mark the beginning of a major . . . . . . . . \beginsection
insert a footnote  mark without text. . . . . . . . . . . . . . . . . . . . \footnotemark^{L}
by which delimiters can fail to span included  material.  amount . . . . . . . . . . \delimitershortfall^{*}
create a box with horizontal mode  material. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \hbox^{*}
create a box with vertical mode  material. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vbox^{*}
use in-text style for  math. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \textstyle^{*}
math accent: acute (x́). . . . . . . . . . . . . . . . . . . . .  \acute
math accent: bar (x̄). . . . . . . . . . . . . . . . . . . . . . .  \bar
math accent: bar under (x̄). . . . . . . . . . . . . . . . .  \b
math accent: breve (x̆). . . . . . . . . . . . . . . . . . . . .  \breve
math accent: check (x̌). . . . . . . . . . . . . . . . . . . . .  \check
math accent: dot (ẋ). . . . . . . . . . . . . . . . . . . . . . .  \dot
math accent: double dot (ẍ). . . . . . . . . . . . . . . .  \ddot
math accent: grave (x̀). . . . . . . . . . . . . . . . . . . . .  \grave
math accent: hat (x̂). . . . . . . . . . . . . . . . . . . . . . .  \hat
math accent: tilde (x̃). . . . . . . . . . . . . . . . . . . . . .  \tilde
math accent: vector (x⃗). . . . . . . . . . . . . . . . . . . .  \vec
math accent: wide hat (x̂). . . . . . . . . . . . . . . . . . \widehat
math accent: wide tilde (x̃). . . . . . . . . . . . . . . \widetilde
tokens to insert when display  math begins. . . . . . . . . . . . . . . . . . . . . . . . . . \everydisplay^{*}
math binary operator: join or V (∨). . . . . . . .  \vee
math binary operator: meet or wedge (∧). .  \wedge
math binary operator: wreath product (≀). .  \wr
define a name for a  math character. . . . . . . . . . . . . . . . . . . . . . . . . \mathchardef^{*}
specify a  math character code. . . . . . . . . . . . . . . . . . . . . . . \mathchar^{*}
set a  math character’s processing category type.  \mathcode^{*}
2.5-line left  math delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Biggl
2-line left  math delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \biggl
2.5-line middle  math delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Biggm
2-line middle  math delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \biggm
2.5-line right  math delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Biggr
2-line right  math delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \biggr
1.5-line left  math delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Bigl
1-line left  math delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bigl
1.5-line middle  math delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Bigm
1-line middle  math delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bigm
1.5-line right  math delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Bigr

1-line right  math delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bigr
math delimiter: backslash (\). . . . . . . . . . . . . \backslash
math delimiter: double vertical bar (∥). . . . .  \Vert
math delimiter: left angle bracket (⟨). . . . . . . \langle
math delimiter: left bracket ([). . . . . . . . . . . . . \lbrack
math delimiter: left ceiling bracket (⌈). . . . . .  \lceil
math delimiter: left curly brace ({). . . . . . . . . \lbrace
math delimiter: left floor bracket (⌊). . . . . . . . \lfloor
math delimiter: left group ({). . . . . . . . . . . . .  \lgroup
math delimiter: right angle bracket (⟩). . . . .  \rangle
math delimiter: right bracket (]). . . . . . . . . . . . \rbrack
math delimiter: right ceiling bracket (⌉)….  \rceil
math delimiter: right curly brace (}). . . . . . .  \rbrace
math delimiter: right floor (⌋). . . . . . . . . . . . . . \rfloor
math delimiter: right group (}). . . . . . . . . . . . \rgroup
1.5-line  math delimiter size. . . . . . . . . . . . . . . . . . . . . . . . .  \Big
1-line  math delimiter size. . . . . . . . . . . . . . . . . . . . . . . . .  \big
2.5-line  math delimiter size. . . . . . . . . . . . . . . . . . . . . . . . .  \Bigg
2-line  math delimiter size. . . . . . . . . . . . . . . . . . . . . . . . .  \bigg
math delimiter: up-and-down arrow (↕). \updownarrow
math delimiter: upward arrow (↑). . . . . . . . . . \uparrow
math delimiter: vertical bar (|). . . . . . . . . . . . .  \vert
begin theorem with special format in  math environment. . . . . . . . . . . . . . . . . . . . . . . . . \proclaim
10 point math extension symbol font. . . . . . . . . . . . . . . . \tenex
place an accent over the next  math field. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathaccent^{*}
put a  math field under a radical. . . . . . . . . . . . . . . . . . \radical^{*}
colon in a math formula. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \colon
horizontal brace over a  math formula. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \overbrace
left arrow over a  math formula. . . . . . . . . . . . . . . . . . . . . . . . \overleftarrow
right arrow over a  math formula. . . . . . . . . . . . . . . . . . . . . . . \overrightarrow
horizontal brace under a  math formula. . . . . . . . . . . . . . . . . . . . . . . . . . . \underbrace
choose a math formula based on the current style. \mathchoice^{*}
medium space in  math formulas. . . . . . . . . . . . . . . . . . . . . . . . . . . . \medmuskip^{*}
thick space in  math formulas. . . . . . . . . . . . . . . . . . . . . . . . . \thickmuskip^{*}
thin space in  math formulas. . . . . . . . . . . . . . . . . . . . . . . . . . \thinmuskip^{*}
math function: arc cosine. . . . . . . . . . . . . . . . . . . \arccos
math function: arc sine. . . . . . . . . . . . . . . . . . . . . \arcsin
math function: arc tangent. . . . . . . . . . . . . . . . . \arctan
math function: arg. . . . . . . . . . . . . . . . . . . . . . . . .  \arg
math function: cosecant. . . . . . . . . . . . . . . . . . . .  \csc
math function: cosine. . . . . . . . . . . . . . . . . . . . . .  \cos
math function: cotangent. . . . . . . . . . . . . . . . . . .  \cot
math function: degree of a polynomial (deg).  \deg
math function: det. . . . . . . . . . . . . . . . . . . . . . . . .  \det
math function: dim. . . . . . . . . . . . . . . . . . . . . . . .  \dim
math function: exp. . . . . . . . . . . . . . . . . . . . . . . . .  \exp
math function: gcd. . . . . . . . . . . . . . . . . . . . . . . . .  \gcd
math function: hom. . . . . . . . . . . . . . . . . . . . . . . .  \hom
math function: hyperbolic cosine. . . . . . . . . . .  \cosh
math function: hyperbolic cotangent. . . . . . .  \coth
math function: hyperbolic sine. . . . . . . . . . . . .  \sinh
math function: hyperbolic tangent. . . . . . . . . .  \tanh

math function: inf. . . . . . . . . . . . . . . . . . . . . . . . . .  \inf
math function: ker. . . . . . . . . . . . . . . . . . . . . . . . .  \ker
math function: lg. . . . . . . . . . . . . . . . . . . . . . . . . .  \lg
math function: lim. . . . . . . . . . . . . . . . . . . . . . . . .  \lim
math function: limit infimum (liminf). . . . . \liminf
math function: limit supremum (limsup)… \limsup
math function: ln. . . . . . . . . . . . . . . . . . . . . . . . . .  \ln
math function: log. . . . . . . . . . . . . . . . . . . . . . . . .  \log
math function: max. . . . . . . . . . . . . . . . . . . . . . . .  \max
math function: min. . . . . . . . . . . . . . . . . . . . . . . .  \min
( (mod )). math function: mod within parentheses …. \pmod
math function: probability (Pr). . . . . . . . . . . .  \Pr
math function: secant. . . . . . . . . . . . . . . . . . . . . .  \sec
math function: sine. . . . . . . . . . . . . . . . . . . . . . . .  \sin
math function: sup. . . . . . . . . . . . . . . . . . . . . . . . .  \sup
math function: tangent. . . . . . . . . . . . . . . . . . . . .  \tan
math Greek letter: alpha (α). . . . . . . . . . . . . . .  \alpha
math Greek letter: beta (β). . . . . . . . . . . . . . . .  \beta
math Greek letter: capital delta (Δ). . . . . . . .  \Delta
math Greek letter: capital gamma (Γ). . . . . .  \Gamma
math Greek letter: capital lambda (Λ). . . . . . \Lambda
math Greek letter: capital omega (Ω). . . . . . .  \Omega
math Greek letter: capital phi (Φ). . . . . . . . . .  \Phi
math Greek letter: capital pi (Π). . . . . . . . . . .  \Pi
math Greek letter: capital psi (Ψ). . . . . . . . . .  \Psi
math Greek letter: capital sigma (Σ). . . . . . .  \Sigma
math Greek letter: capital theta (Θ). . . . . . . .  \Theta
math Greek letter: capital upsilon (Υ). . . . . . \Upsilon
math Greek letter: capital xi (Ξ). . . . . . . . . . .  \Xi
math Greek letter: chi (χ). . . . . . . . . . . . . . . . . .  \chi
math Greek letter: delta (δ). . . . . . . . . . . . . . . .  \delta
math Greek letter: epsilon (ϵ). . . . . . . . . . . . . . \epsilon
math Greek letter: eta (η). . . . . . . . . . . . . . . . . .  \eta
math Greek letter: gamma (γ). . . . . . . . . . . . .  \gamma
math Greek letter: kappa (κ). . . . . . . . . . . . . . .  \kappa
math Greek letter: lambda (λ). . . . . . . . . . . . . \lambda
math Greek letter: mu (μ). . . . . . . . . . . . . . . . .  \mu
math Greek letter: nu (ν). . . . . . . . . . . . . . . . . .  \nu
math Greek letter: omega (ω). . . . . . . . . . . . . .  \omega
math Greek letter: phi (ϕ). . . . . . . . . . . . . . . . .  \phi
math Greek letter: pi (π). . . . . . . . . . . . . . . . . . .  \pi
math Greek letter: psi (ψ). . . . . . . . . . . . . . . . . .  \psi
math Greek letter: rho (ρ). . . . . . . . . . . . . . . . .  \rho
math Greek letter: sigma (σ). . . . . . . . . . . . . . .  \sigma
math Greek letter: tau (τ). . . . . . . . . . . . . . . . .  \tau
math Greek letter: theta (θ). . . . . . . . . . . . . . . .  \theta
math Greek letter: upsilon (υ). . . . . . . . . . . . . . \upsilon
math Greek letter: variant epsilon (ε)… \varepsilon
math Greek letter: variant phi (φ). . . . . . . . . . \varphi
math Greek letter: variant pi (ϖ). . . . . . . . . .  \varpi
math Greek letter: variant rho (ϱ). . . . . . . . . . \varrho
math Greek letter: variant sigma (ς). . . . . . . \varsigma
math Greek letter: variant theta (ϑ). . . . . . . \vartheta

math Greek letter: xi (ξ). . . . . . . . . . . . . . . . . . .  \xi
math Greek letter: zeta (ζ). . . . . . . . . . . . . . . . .  \zeta
define kerning around  math in text. . . . . . . . . . . . . . . . . . . . . . . . . . \mathsurround^{*}
tokens to insert when  math in text begins. . . . . . . . . . . . . . . . . . . . . . . \everymath^{*}
select bold  math italic and symbol fonts. . . . . . . . . . . . . . . \boldmath^{L}
unselect bold  math italic and symbol fonts. . . . . . . . . . . . \unboldmath^{L}
8 point  math italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \egtmi^{L}
11 point  math italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvmi^{L}
5 point  math italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \fivei
5 point  math italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \fivmi^{L}
14 point  math italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \frtnmi^{L}
9 point  math italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ninmi^{L}
7 point  math italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \seveni
7 point  math italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sevmi^{L}
6 point  math italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sixmi^{L}
17 point  math italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \svtnmi^{L}
10 point  math italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \teni
10 point  math italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tenmi^{L}
12 point  math italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twlmi^{L}
20 point  math italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twtymi^{L}
math join operator: large V (⋁). . . . . . . . . . . . \bigvee
place math limits above and below math operators. \limits^{*}
create a centered vbox in a  math list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vcenter^{*}
begin a new  math list with a left delimiter. . . . . . . . . . . . . .  \left^{*}
end a  math list with a right delimiter. . . . . . . . . . . . .  \right^{*}
math meet operator: large wedge (⋀). . . . . . \bigwedge
medium space in  math mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \>
centered dot with special spacing in  math mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \cdotp
define a delimiter for  math mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \delimiter^{*}
enter  math mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  $^{*}
enter display  math mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  $$^{*}
an ellipsis, equivalent to \ldots in  math mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \dots
test for  math mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ifmmode^{*}
begin  math mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \(^{L}
begin display  math mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \[^{L}
end  math mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \)^{L}
end display  math mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \]^{L}
medium space in  math mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \:^{L}
kern in  math mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \mkern^{*}
insert horizontal space in  math mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \mskip^{*}
increase line separation in  math mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \openup
math mode calligraphic letters font. . . . . . . . .  \cal
three centered dots in math mode (⋯). . . . . . . . . . . . . . . . . . . . . . . . . . . . \cdots
unbreakable tiny negative  math mode horizontal space. . . . . . . . . . \negthinspace
math mode italic font. . . . . . . . . . . . . . . . . . . . . .  \mit
three low dots in math mode (…). . . . . . . . . . . . . . . . . . . . . . . . . . . . \ldots
math mode negative thin space. . . . . . . . . . . . .  \!
three diagonal dots in math mode (...). . . . . . . . . . . . . . . . . . . . . . . . . . . . \ddots
define a binary  math operator. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathbin^{*}
define a large  math operator. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathop^{*}
define an ordinary  math operator. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathord^{*}
(⨿). math operator: amalgamated sum, co-product \amalg
math operator: asterisk (∗). . . . . . . . . . . . . . . . .  \ast

math operator: binary modulo (mod). . . . . .  \bmod
math operator: bullet (•). . . . . . . . . . . . . . . . . . . \bullet
math operator: centered dot (·). . . . . . . . . . . . .  \cdot
math operator: circle (∘). . . . . . . . . . . . . . . . . . .  \circ
math operator: circle dot (⊙). . . . . . . . . . . . . .  \odot
math operator: circle minus (⊖). . . . . . . . . . . . \ominus
math operator: circle slash (⊘). . . . . . . . . . . . . \oslash
math operator: contour integral (∮). . . . . . . .  \oint
math operator: dagger (†). . . . . . . . . . . . . . . . . . \dagger
math operator: diamond (⋄). . . . . . . . . . . . . . . . \diamond
math operator: direct sum, circle plus (⊕). .  \oplus
math operator: discretionary multiply sign.  \*
math operator: div. . . . . . . . . . . . . . . . . . . . . . . . .  \div
math operator: double dagger (‡). . . . . . . . . . . \ddagger
math operator: integral (∫). . . . . . . . . . . . . . . .  \int
math operator: intersection or cap (∩). . . . . .  \cap
math operator: large cap (⋂). . . . . . . . . . . . . . . \bigcap
math operator: large circle (◯). . . . . . . . . . . . \bigcirc
math operator: large circle with dot (⨀)… \bigodot
math operator: large circle with plus (⨁). . \bigoplus
math operator: large circle with times (⨂). \bigotimes
math operator: large co-product (∐). . . . . . . \coprod
math operator: large cup (⋃). . . . . . . . . . . . . .  \bigcup
(▽).  math operator: large down triangle \bigtriangledown
math operator: large product (∏). . . . . . . . . .  \prod
math operator: large square cup (⨆). . . . . . . \bigsqcup
math operator: large sum (∑). . . . . . . . . . . . .  \sum
math operator: large U plus (⨄). . . . . . . . . . . \biguplus
math operator: large up triangle (△). \bigtriangleup
math operator: left triangle (◁). . . . . . . \triangleleft
math operator: logical and (∧). . . . . . . . . . . . .  \land
math operator: logical not, hook (¬). . . . . . .  \lnot
math operator: minus plus (∓). . . . . . . . . . . . .  \mp
math operator: negate (¬). . . . . . . . . . . . . . . . . .  \neg
math operator: not in (∉). . . . . . . . . . . . . . . . . .  \notin
math operator: not ( ). . . . . . . . . . . . . . . . . . . .  \not
math operator: plus or minus (±). . . . . . . . . .  \pm
math operator: right triangle (▷). . . . . \triangleright
math operator: set minus (\). . . . . . . . . . . . . . \setminus
math operator: small integral (∫). . . . . . . . . . \smallint
math operator: square cap (⊓). . . . . . . . . . . . .  \sqcap
math operator: square cup (⊔). . . . . . . . . . . . .  \sqcup
math operator: star (⋆). . . . . . . . . . . . . . . . . . . . .  \star
(⊗). math operator: tensor product, circle times \otimes
math operator: times (×). . . . . . . . . . . . . . . . . .  \times
math operator: U plus (⊎). . . . . . . . . . . . . . . . .  \uplus
math operator: union or cup (∪). . . . . . . . . . .  \cup
restore default limit placement on large  math operators. . . . . . . . . . . . . . . . . . . . . . \displaylimits^{*}
place math limits above and below  math operators. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \limits^{*}
place superscripts and subscripts after  math operators. . . . . . . . . . . . . . . . . . . . . . . . . . . . \nolimits^{*}
define a  math punctuation operator. . . . . . . . . . . . . . . \mathpunct^{*}
penalty for line break after  math relation. . . . . . . . . . . . . . . . . . . . . . . . . . . \relpenalty^{*}
math relation: approximately equal (≈)….  \approx

math relation: asymptote (≍). . . . . . . . . . . . . .  \asymp
math relation: bowtie (▷◁). . . . . . . . . . . . . . . . . . \bowtie
math relation: congruent (≅). . . . . . . . . . . . . . .  \cong
math relation: contains (∋). . . . . . . . . . . . . . . . .  \ni
math relation: dash V (⊣). . . . . . . . . . . . . . . . . .  \dashv
math relation: dash (⊢). . . . . . . . . . . . . . . . . . . .  \vdash
math relation: dotted equal (≐). . . . . . . . . . . .  \doteq
math relation: equivalence (≡). . . . . . . . . . . . .  \equiv
math relation: frown (⌢). . . . . . . . . . . . . . . . . .  \frown
math relation: gets (←). . . . . . . . . . . . . . . . . . . .  \gets
math relation: greater or equal (≥). . . . . . . . .  \ge
math relation: greater or equal (≥). . . . . . . . .  \geq
math relation: if and only if ( ⟺ ). . . . . . . .  \iff
math relation: in (∈). . . . . . . . . . . . . . . . . . . . . . .  \in
math relation: less or equal (≤). . . . . . . . . . . .  \le
math relation: less or equal (≤). . . . . . . . . . . .  \leq
math relation: mid (|). . . . . . . . . . . . . . . . . . . . . .  \mid
math relation: models (⊧). . . . . . . . . . . . . . . . .  \models
math relation: much greater (≫). . . . . . . . . . .  \gg
math relation: much less (≪). . . . . . . . . . . . . . .  \ll
math relation: not equal (≠). . . . . . . . . . . . . . .  \ne
math relation: not equal (≠). . . . . . . . . . . . . . .  \neq
math relation: not equal (≠). . . . . . . . . . . . . . .  \not=
define a  math relation operator. . . . . . . . . . . . . . . . . . . . . \mathrel^{*}
math relation: owns (∋). . . . . . . . . . . . . . . . . . . .  \owns
math relation: parallel (∥). . . . . . . . . . . . . . . . . \parallel
math relation: perpendicular (⊥). . . . . . . . . . .  \perp
math relation: precedes or equal (⪯). . . . . . .  \preceq
math relation: precedes (≺). . . . . . . . . . . . . . . .  \prec
math relation: proportional to (∝). . . . . . . . .  \propto
math relation: similar or equal (≃). . . . . . . . .  \simeq
math relation: similar (∼). . . . . . . . . . . . . . . . . .  \sim
math relation: smile (⌣). . . . . . . . . . . . . . . . . . .  \smile
math relation: square subset or equal (⊑). \sqsubseteq
(⊒).  math relation: square superset or equal . \sqsupseteq
math relation: subset or equal (⊆). . . . . . . . . \subseteq
math relation: subset (⊂). . . . . . . . . . . . . . . . . .  \subset
math relation: successor or equal (⪰). . . . . . . \succeq
math relation: successor (≻). . . . . . . . . . . . . . . .  \succ
math relation: superset or equal (⊇). . . . . . . \supseteq
math relation: superset (⊃). . . . . . . . . . . . . . . .  \supset
math relation: to (→). . . . . . . . . . . . . . . . . . . . . .  \to
put symbols over  math relations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \buildrel
use normal  math script style. . . . . . . . . . . . . . . . . . . . . . \displaystyle^{*}
use very small  math script style. . . . . . . . . . . . . . . . \scriptscriptstyle^{*}
use small  math script style. . . . . . . . . . . . . . . . . . . . . . . \scriptstyle^{*}
select font for small  math scripts. . . . . . . . . . . . . . . . . . . . . . . . . . . . \scriptfont^{*}
select font for very small  math scripts. . . . . . . . . . . . . . . . . . . . . \scriptscriptfont^{*}
math skip register. . . . . . . . . . . . . . . . . . . . . . . . . . \muskip^{*}
define a name for a  math skip register. . . . . . . . . . . . . . . . . . . . . . . . \muskipdef^{*}
allocate a new  math skip register. . . . . . . . . . . . . . . . . . . . . . . . \newmuskip
next item.  select  math spacing of a closing delimiter for the  \mathclose^{*}
next item.  select  math spacing of an opening delimiter for the \mathopen^{*}

parentheses.  math strut with height and depth of . . . . . \mathstrut
define an inner  math subformula. . . . . . . . . . . . . . . . . . . . . . . . . \mathinner^{*}
math symbol: aleph (ℵ). . . . . . . . . . . . . . . . . . . .  \aleph
math symbol: bottom (⊥). . . . . . . . . . . . . . . . . .  \bot
math symbol: box (□). . . . . . . . . . . . . . . . . . . . . .  \Box^{L}
math symbol: club suit (♣). . . . . . . . . . . . . . . . \clubsuit
math symbol: diamond (◇). . . . . . . . . . . . . . . . \Diamond^{L}
math symbol: diamond suit (♢). . . . . . . . \diamondsuit
math symbol: dotless i (ı). . . . . . . . . . . . . . . . . .  \imath
math symbol: dotless j (ȷ). . . . . . . . . . . . . . . . . .  \jmath
math symbol: double relation bar (=). . . . . . \Relbar
math symbol: double vertical bar (‖). . . . . \Arrowvert
math symbol: down arrow (↓). . . . . . . . . . . . \downarrow
math symbol: downward double arrow (⇓). \Downarrow
math symbol: empty set (∅). . . . . . . . . . . . . . . \emptyset
math symbol: exists quantifier (∃). . . . . . . . . . \exists
math symbol: flat (♭). . . . . . . . . . . . . . . . . . . . . . .  \flat
8 point  math symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \egtsy^{L}
11 point  math symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \elvsy^{L}
5 point  math symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \fivesy
5 point  math symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \fivsy^{L}
14 point  math symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \frtnsy^{L}
9 point  math symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \ninsy^{L}
7 point  math symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . \sevensy
7 point  math symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \sevsy^{L}
6 point  math symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \sixsy^{L}
17 point  math symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \svtnsy^{L}
10 point  math symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \tensy
12 point  math symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \twlsy^{L}
20 point  math symbol font. . . . . . . . . . . . . . . . . . . . . . . . . .  \twtysy^{L}
math symbol: for-all quantifier (∀). . . . . . . . .  \forall
math symbol: heart suit (♡). . . . . . . . . . . . . . \heartsuit
math symbol: hook left arrow (↩)… \hookleftarrow
math symbol: hook right arrow (↪). \hookrightarrow
math symbol: imaginary, Fraktur I (ℑ). . . . .  \Im
math symbol: infinity (∞). . . . . . . . . . . . . . . . . .  \infty
math symbol: iota (ι). . . . . . . . . . . . . . . . . . . . . .  \iota
math symbol: join (⋈). . . . . . . . . . . . . . . . . . . . .  \Join^{L}
math symbol: leads to (⇝). . . . . . . . . . . . . . . . . \leadsto^{L}
math symbol: left arrow (←). . . . . . . . . . . . . \leftarrow
math symbol: left double arrow (⇐). . . . . . \Leftarrow
math symbol: left half diamond (⊲). . . . . . . .  \lhd^{L}
(↽).  math symbol: left harpoon down . . \leftharpoondown
math symbol: left harpoon up (↼). . \leftharpoonup
math symbol: left hook ( ). . . . . . . . . . . . . . . . .  \lhook
math symbol: left moustache ({). . . . . . . \lmoustache
math symbol: left quote (‘). . . . . . . . . . . . . . . . .  \lq
math symbol: left-right arrow (↔). . \leftrightarrow
(⇔).  math symbol: left-right double arrow \Leftrightarrow
math symbol: logical or (∨). . . . . . . . . . . . . . . .  \lor
arrow (⟷).  math symbol: long left and right \longleftrightarrow
math symbol: long left arrow (⟵). . \longleftarrow

(⟸).  math symbol: long left double arrow . \Longleftarrow
double arrow (⟺).  math symbol: long left-right … \Longleftrightarrow
math symbol: long maps to (⟼). . . . . . . \longmapsto
math symbol: long right arrow (⟶). \longrightarrow
(⟹).  math symbol: long right double arrow \Longrightarrow
math symbol: maps to char ( ). . . . . . . . . . \mapstochar
math symbol: maps to (↦). . . . . . . . . . . . . . . . . \mapsto
math symbol: mho (℧). . . . . . . . . . . . . . . . . . . . .  \mho^{L}
math symbol: nabla (∇). . . . . . . . . . . . . . . . . . . .  \nabla
math symbol: natural (♮). . . . . . . . . . . . . . . . . . . \natural
math symbol: northeast arrow (↗). . . . . . . . . \nearrow
math symbol: northwest arrow (↖). . . . . . . . \nwarrow
math symbol: paragraph (¶). . . . . . . . . . . . . . .  \P
math symbol: partial (∂). . . . . . . . . . . . . . . . . . . \partial
math symbol: Planck’s constant or h-bar (h̄).  \hbar
math symbol: prime (x′). . . . . . . . . . . . . . . . . . .  \prime
math symbol: real, Fraktur R (ℜ). . . . . . . . . .  \Re
math symbol: relation bar (−). . . . . . . . . . . . .  \relbar
math symbol: right arrow (→). . . . . . . . . . \rightarrow
math symbol: right double arrow (⇒)… \Rightarrow
math symbol: right half diamond (⊳). . . . . . .  \rhd^{L}
(⇁).  math symbol: right harpoon down \rightharpoondown
math symbol: right harpoon up (⇀). \rightharpoonup
math symbol: right hook ( ). . . . . . . . . . . . . . . .  \rhook
(⇀ ↽). math symbol: right left harpoon \rightleftharpoons
math symbol: right moustache (}). . . . . . \rmoustache
math symbol: section (§). . . . . . . . . . . . . . . . . . .  \S
math symbol: sharp (♯). . . . . . . . . . . . . . . . . . . . .  \sharp
math symbol: small script L (ℓ). . . . . . . . . . . .  \ell
math symbol: southeast arrow (↘). . . . . . . . . \searrow
math symbol: southwest arrow (↙). . . . . . . . \swarrow
math symbol: spade suit (♠). . . . . . . . . . . . . \spadesuit
math symbol: square subset (⊏). . . . . . . . . . . \sqsubset^{L}
math symbol: square superset (⊐). . . . . . . . . \sqsupset^{L}
math symbol: surd (√). . . . . . . . . . . . . . . . . . . . .  \surd
math symbol: top (⊤). . . . . . . . . . . . . . . . . . . . . .  \top
math symbol: triangle (△). . . . . . . . . . . . . . . . . \triangle
(⊴).  math symbol: underlined left half diamond .  \unlhd^{L}
(⊵).  math symbol: underlined right half diamond  \unrhd^{L}
(⇕).  math symbol: up-and-down double arrow \Updownarrow
math symbol: upward double arrow (⇑)…. \Uparrow
math symbol: vertical bar (|). . . . . . . . . . . . \arrowvert
math symbol: vertical bar (∥). . . . . . . . . . . . . .  \|
math symbol: vertical dots (⋮). . . . . . . . . . . . . .  \vdots
math symbol: Weierstrass p (℘). . . . . . . . . . . .  \wp
lower dot math symbol with special spacing. . . . . . . . . . \ldotp
internal Plain TEX operation to define  math text symbols. . . . . . . . . . . . . . . . . . . . . . \mathhexbox
generate a  matrix labeled on rows and columns. . \bordermatrix
generate a matrix with a left brace delimiter. . . . . . . . . . . \cases
generate a  matrix with parentheses delimiters. . . . . . . . . \pmatrix
generate a  matrix without delimiters. . . . . . . . . . . . . . . . . .  \matrix
math function:  max. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \max

maximum box depth shown. . . . . . . . . . . \showboxdepth^{*}
given depth.  maximum boxed items shown at a . . \showboxbreadth^{*}
maximum depth of boxes in a split box. \splitmaxdepth^{*}
pages.  maximum depth of boxes on explicit … \boxmaxdepth^{*}
maximum depth of boxes on main pages… \maxdepth^{*}
messages occur. maximum overrun before overfull hbox . . . . . \hfuzz^{*}
maximum overrun before overfull vbox error.  \vfuzz^{*}
restores a carriage return to its usual  meaning. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \restorecr^{L}
vertical skip a  medium amount. . . . . . . . . . . . . . . . . . . . . . . . . . . . \medskip^{L}
breakable medium horizontal skip. . . . . . . . . . . . . . . . . . . . . \quad
medium space in math formulas. . . . . . . . . . \medmuskip^{*}
medium space in math mode. . . . . . . . . . . . . . .  \>
medium space in math mode. . . . . . . . . . . . . . .  \:^{L}
medium vertical space. . . . . . . . . . . . . . . . . . . . . . \medskip
medium vertical space or a good page break. \medbreak
\medskip space. . . . . . . . . . . . . . . . . . . . . . \medskipamount
math  meet operator: large wedge (⋀). . . . . . . . . . . . \bigwedge
math binary operator: meet or wedge (∧). . . . . . . . . . . . . . . . . . . . . . . . . . \wedge
show statistics about  memory usage. . . . . . . . . . . . . . . . . . . . . . . . \tracingstats^{*}
of lines of context to be displayed in an error message. number . . . . . . . . . . . . \errorcontextlines^{3}
define a new help  message. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newhelp
write a  message on the terminal. . . . . . . . . . . . . . . . . . . . \typeout^{L}
help  message to display if user asks for help. . . . . \errhelp^{*}
write balanced error  message to the terminal. . . . . . . . . . . . . . . . . \errmessage^{*}
maximum overrun before overfull hbox  messages occur. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \hfuzz^{*}
math symbol:  mho (℧). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \mho^{L}
math relation:  mid (|). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \mid
2.5-line  middle math delimiter. . . . . . . . . . . . . . . . . . . . . .  \Biggm
2-line  middle math delimiter. . . . . . . . . . . . . . . . . . . . . .  \biggm
1.5-line  middle math delimiter. . . . . . . . . . . . . . . . . . . . . .  \Bigm
1-line  middle math delimiter. . . . . . . . . . . . . . . . . . . . . .  \bigm
vertical space around a float in the  middle of a page. . . . . . . . . . . . . . . . . . . . . . . . . . \intextsep^{L}
math function: min. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \min
after hyphenation at the end of a word.  minimum number of characters \righthyphenposition^{3}
before hyphenation at the start of a word. minimum number of characters \lefthyphenposition^{3}
vertical list.  minimum space between lines in a … \lineskiplimit^{*}
math operator: circle  minus (⊖). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ominus
math operator:  minus plus (∓). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \mp
math operator: plus or  minus (±). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \pm
math operator: set  minus (\). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \setminus
math function:  mod within parentheses ( (mod )). . . . . . . . .  \pmod
medium space in math  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \>
thick space (math  mode). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \;
thin space (math  mode). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \,
centered dot with special spacing in math  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \cdotp
define a delimiter for math  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \delimiter^{*}
enter math  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  $^{*}
enter display math  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  $$^{*}
an ellipsis, equivalent to \ldots in math  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \dots
test for horizontal  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ifhmode^{*}
test for an internal  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ifinner^{*}
test for math  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ifmmode^{*}
test for vertical  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ifvmode^{*}

begin display math  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \[^{L}
begin math  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \(^{L}
end display math  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \]^{L}
end math  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \)^{L}
medium space in math  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \:^{L}
switch to horizontal mode from vertical  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leavevmode
kern in math  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \mkern^{*}
insert horizontal space in math  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \mskip^{*}
increase line separation in math  mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \openup
math  mode calligraphic letters font. . . . . . . . . . . . . . .  \cal
three centered dots in math  mode (⋯). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \cdots
switch to horizontal  mode from vertical mode. . . . . . . . . . . . . . . . \leavevmode
unbreakable tiny negative math  mode horizontal space. . . . . . . . . . . . . . . . \negthinspace
math  mode italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \mit
three low dots in math  mode (…). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ldots
create a box with horizontal  mode material. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \hbox^{*}
create a box with vertical  mode material. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vbox^{*}
create a box with vertical  mode material with the baseline at the top.  \vtop^{*}
math  mode negative thin space. . . . . . . . . . . . . . . . . . .  \!
three diagonal dots in math  mode (...). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ddots
math relation:  models (⊧). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \models
math operator: binary  modulo (mod). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bmod
current day of the  month. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \day^{*}
current  month of the year. . . . . . . . . . . . . . . . . . . . . . . . . .  \month^{*}
math symbol: left  moustache ({). . . . . . . . . . . . . . . . . . . . . . . . . . \lmoustache
math symbol: right  moustache (}). . . . . . . . . . . . . . . . . . . . . . . . . . \rmoustache
environment. move to next tab position in tabbing . . . . . . . \>^{L}
protect fragile commands and  moving arguments. . . . . . . . . . . . . . . . . . . . . . . . . . \protect^{L}
math Greek letter:  mu (μ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \mu
math relation:  much greater (≫). . . . . . . . . . . . . . . . . . . . . . . . . .  \gg
math relation:  much less (≪). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ll
create a  multicolumn entry in an aligned table. . . . . .  \span^{*}
environment.  multicolumn entry in array or tabular . . \multicolumn^{L}
environment.  multicolumn line in array or tabular . . . . . . . .  \cline^{L}

         multiple of normal \baselineskip. \baselinestretch^{L}

the following macro may have  multiple paragraphs as parameters. . . . . . . . . .  \long^{*}
multiply a register by a value. . . . . . . . . . . . . . \multiply^{*}
math operator: discretionary  multiply sign. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \*
math symbol:  nabla (∇). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \nabla
begin expanding tokens to construct a command  name. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \csname^{*}
use a  named box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \usebox^{L}
justify page bottoms to their  natural height. . . . . . . . . . . . . . . . . . . . . . . . \raggedbottom^{L}
natural height of page so far. . . . . . . . . . . . . . \pagetotal^{*}
math symbol:  natural (♮). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \natural
math operator: negate (¬). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \neg
unbreakable tiny  negative math mode horizontal space. . \negthinspace
math mode negative thin space. . . . . . . . . . . . . . . . . . . . . . . . . \!
insert a token after the  next assignment command. . . . . . . . . \afterassignment^{*}
put an accent over the  next character. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \accent^{*}
test the  next character. . . . . . . . . . . . . . . . . . . . . . . . . \@ifnextchar^{L}
select math spacing of a closing delimiter for the  next item. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathclose^{*}
math spacing of an opening delimiter for the  next item.  select . . . . . . . . . . . . . . . . . . . . . . . \mathopen^{*}

place an accent over the  next math field. . . . . . . . . . . . . . . . . . . . . . . . . . \mathaccent^{*}
move to next tab position in tabbing environment… \>^{L}
expand the token following the  next token. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \expandafter^{*}
suppress interline space before  next vertical box. . . . . . . . . . . . . . . . . . \nointerlineskip
commands. set @ non-alphabetic to hide internal . . . . . . . . . \makeatother^{L}
select font for  non-math text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \textfont^{*}
tokens to insert after every \cr or  nonredundant \crcr. . . . . . . . . . . . . . . . . . . . . . . \everycr^{*}
read, expand, then ignore tokens until a  non-space is found. . . . . . . . . . . . . . . . . . . . \ignorespaces^{*}

multiple of  normal \baselineskip. . . . . . . . . . . . \baselinestretch^{L}
normal \baselineskip. . . . . . . . \normalbaselineskip

process TEX input without pausing for  normal errors. . . . . . . . . . . . . . . . . . . . . . . . . . . \scrollmode^{*}
input. pause for normal errors while processing TEX . . \errorstopmode^{*}
select  normal font size. . . . . . . . . . . . . . . . . . . . . . . . . \normalsize^{L}
value of  normal \lineskip. . . . . . . . . . . . . . . . . . \normallineskip

value of  normal \lineskiplimit. . . . . . \normallineskiplimit

use  normal math script style. . . . . . . . . . . . . . \displaystyle^{*}
put marginal notes on  normal side of page. . . . . . . . . . . . . . . . \normalmarginpar^{L}
select font four steps larger than  normal size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \huge^{L}
select font one step larger than  normal size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \large^{L}
select font three steps larger than  normal size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \LARGE^{L}
select font two steps larger than  normal size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Large^{L}
normal space between lines. . . . . . . . . . . . \baselineskip^{*}
set line spacing to  normal values. . . . . . . . . . . . . . . . . . . . . \normalbaselines
math symbol:  northeast arrow (↗). . . . . . . . . . . . . . . . . . . . . . . \nearrow
math symbol:  northwest arrow (↖). . . . . . . . . . . . . . . . . . . . . . . \nwarrow
Norwegian letter: capital O with slash (Ø). .  \O
Norwegian letter: o with slash (ø). . . . . . . . . .  \o
the following macro must  not be called from another macro. . . . . . . . . . .  \outer^{*}
math relation:  not equal (≠). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ne
math relation:  not equal (≠). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \neq
math relation:  not equal (≠). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \not=
math operator: logical  not, hook (¬). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \lnot
math operator:  not in (∉). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \notin
show characters  not in the font. . . . . . . . . . . . . . . . . . . \tracinglostchars^{*}
math operator: not ( ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \not
produce a marginal  note. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \marginpar^{L}
distance between marginal  note and text. . . . . . . . . . . . . . . . . . . . . . . . . \marginparsep^{L}
vertical space between marginal  notes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \marginparpush^{L}
width of marginal  notes. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \marginparwidth^{L}
put marginal  notes on normal side of page. . . . . . \normalmarginpar^{L}
put marginal  notes on opposite side of page. … \reversemarginpar^{L}
do  nothing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \relax^{*}
math Greek letter:  nu (ν). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \nu
width of a  null delimiter. . . . . . . . . . . . . . . . . . \nulldelimiterspace^{*}
most recently allocated register  number. . . . . . . . . . . . . . . . . . . . . . . . . . \allocationnumber
equation  number. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \eqno^{*}
current family  number. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \fam^{*}
typeset page  number. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \folio
left equation  number. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \leqno^{*}
current page  number. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \pageno
produces the footnote  number. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \thefootnote^{L}
current page  number. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \thepage^{L}
produce a  numbered caption. . . . . . . . . . . . . . . . . . . . . . . . . . \caption^{L}

suppress  numbering of displayed equations. . . . . . . . . . \nonumber^{L}
specify page  numbering style. . . . . . . . . . . . . . . . . . . . . \pagenumbering^{L}
align a stack of equations with equation  numbers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \eqalignno
align a stack of equations with left equation  numbers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leqalignno
suppress page  numbers. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \nopagenumbers
display counter as Arabic  numerals. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \arabic^{L}
display counter as lower-case Roman  numerals. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \roman^{L}
display counter as upper-case Roman  numerals. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Roman^{L}
convert a number to lower-case Roman  numerals. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \romannumeral^{*}
select old-style  numerals with descenders. . . . . . . . . . . . . . . . . . \oldstyle
specify a character by its  numeric code. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \char^{*}
convert a  numeric register to displayable form. . . . . . . .  \the^{*}
Norwegian letter:  o with slash (ø). . . . . . . . . . . . . . . . . . . . . . . . . . . .  \o
Norwegian letter: capital  O with slash (Ø). . . . . . . . . . . . . . . . . . . . . . . . . . .  \O
place several copies of a picture  object. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \multiput^{L}
place a picture object. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \put^{L}
portion of page that may be  occupied by floats. . . . . . . . . . . . . . . \floatpagefraction^{L}
maximum overrun before overfull hbox messages  occur. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \hfuzz^{*}
pound, hatch mark, sharp sign,  octothorpe (#). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \#
test for  odd integer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ifodd^{*}
left hand margin on  odd pages. . . . . . . . . . . . . . . . . . . . . . . . . . . \oddsidemargin^{L}
ligature digraph symbol  oe (œ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \oe
ligature digraph symbol capital  OE (Œ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \OE
horizontal  offset of a page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \hoffset^{*}
vertical  offset of a page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \voffset^{*}
select  old-style numerals with descenders. . . . . . . . . \oldstyle
math Greek letter:  omega (ω). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \omega
math Greek letter: capital  omega (Ω). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Omega
advance \pageno by  one. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \advancepageno
terminate TEX and write a format file: INITEX  only. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \dump^{*}
write to the log file  only. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \wlog
select math spacing of an  opening delimiter for the next item. . . . . . . . \mathopen^{*}
unit of measure for opening up displays. . . . . . . . . . . . . . . . . . . . . . . . \jot
penalty for line break after binary  operation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \binoppenalty^{*}
internal Plain TEX  operation to define math text symbols… \mathhexbox
define a binary math  operator. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathbin^{*}
define a large math  operator. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathop^{*}
define an ordinary math  operator. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathord^{*}
define a math punctuation  operator. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathpunct^{*}
define a math relation  operator. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathrel^{*}
math  operator: amalgamated sum, co-product (⨿).  \amalg
math  operator: asterisk (∗). . . . . . . . . . . . . . . . . . . . . . .  \ast
math  operator: binary modulo (mod). . . . . . . . . . . .  \bmod
math  operator: bullet (•). . . . . . . . . . . . . . . . . . . . . . . . . \bullet
math  operator: centered dot (·). . . . . . . . . . . . . . . . . .  \cdot
math  operator: circle (∘). . . . . . . . . . . . . . . . . . . . . . . . .  \circ
math  operator: circle dot (⊙). . . . . . . . . . . . . . . . . . . .  \odot
math  operator: circle minus (⊖). . . . . . . . . . . . . . . . . . \ominus
math  operator: circle slash (⊘). . . . . . . . . . . . . . . . . . . \oslash
math operator: contour integral (∮). . . . . . . . . . . . . . \oint
math  operator: dagger (†). . . . . . . . . . . . . . . . . . . . . . . . \dagger
math  operator: diamond (⋄). . . . . . . . . . . . . . . . . . . . . . \diamond
math  operator: direct sum, circle plus (⊕). . . . . . . .  \oplus

math  operator: discretionary multiply sign. . . . . . .  \*
math  operator: div. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \div
math  operator: double dagger (‡). . . . . . . . . . . . . . . . \ddagger
math  operator: integral (∫). . . . . . . . . . . . . . . . . . . . . .  \int
math  operator: intersection or cap (∩). . . . . . . . . . .  \cap
math binary  operator: join or V (∨). . . . . . . . . . . . . . . . . . . . .  \vee
math  operator: large cap (⋂). . . . . . . . . . . . . . . . . . . . . \bigcap
math  operator: large circle (◯). . . . . . . . . . . . . . . . . . \bigcirc
math  operator: large circle with dot (⨀). . . . . . . . . \bigodot
math  operator: large circle with plus (⨁). . . . . . . \bigoplus
math  operator: large circle with times (⨂). . . . . \bigotimes
math  operator: large co-product (∐). . . . . . . . . . . . . \coprod
math  operator: large cup (⋃). . . . . . . . . . . . . . . . . . . .  \bigcup
math  operator: large down triangle (▽).  \bigtriangledown
math  operator: large product (∏). . . . . . . . . . . . . . . .  \prod
math  operator: large square cup (⨆). . . . . . . . . . . . . \bigsqcup
math operator: large sum (∑). . . . . . . . . . . . . . . . . . . \sum
math  operator: large U plus (⨄). . . . . . . . . . . . . . . . . \biguplus
math  operator: large up triangle (△). . . . . . \bigtriangleup
math join  operator: large V (⋁). . . . . . . . . . . . . . . . . . . . . . \bigvee
math meet  operator: large wedge (⋀). . . . . . . . . . . . . . . . . \bigwedge
math  operator: left triangle (◁). . . . . . . . . . . . . \triangleleft
math  operator: logical and (∧). . . . . . . . . . . . . . . . . . .  \land
math  operator: logical not, hook (¬). . . . . . . . . . . . .  \lnot
math binary  operator: meet or wedge (∧). . . . . . . . . . . . . . . .  \wedge
math  operator: minus plus (∓). . . . . . . . . . . . . . . . . . .  \mp
math  operator: negate (¬). . . . . . . . . . . . . . . . . . . . . . .  \neg
math  operator: not in (∉). . . . . . . . . . . . . . . . . . . . . . . .  \notin
math  operator: not ( ). . . . . . . . . . . . . . . . . . . . . . . . . .  \not
math  operator: plus or minus (±). . . . . . . . . . . . . . . .  \pm
math  operator: right triangle (▷). . . . . . . . . . . \triangleright
math  operator: set minus (\). . . . . . . . . . . . . . . . . . . . \setminus
math  operator: small integral (∫). . . . . . . . . . . . . . . . \smallint
math  operator: square cap (⊓). . . . . . . . . . . . . . . . . . .  \sqcap
math  operator: square cup (⊔). . . . . . . . . . . . . . . . . . .  \sqcup
math  operator: star (⋆). . . . . . . . . . . . . . . . . . . . . . . . . .  \star
math  operator: tensor product, circle times (⊗). .  \otimes
math  operator: times (×). . . . . . . . . . . . . . . . . . . . . . . .  \times
math  operator: U plus (⊎). . . . . . . . . . . . . . . . . . . . . . .  \uplus
math  operator: union or cup (∪). . . . . . . . . . . . . . . . .  \cup
math binary  operator: wreath product (≀). . . . . . . . . . . . . . .  \wr
restore default limit placement on large math  operators. . . . . . . . . . . . . . . . . . . . . . . . . . . . \displaylimits^{*}
place math limits above and below math  operators. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \limits^{*}
place superscripts and subscripts after math  operators. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \nolimits^{*}
put marginal notes on  opposite side of page. . . . . . . . . . . . . \reversemarginpar^{L}
declare document style and  options. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \documentstyle^{L}
define an  ordinary math operator. . . . . . . . . . . . . . . . . . . . \mathord^{*}
begin a tabbed line in an outer environment. . . . . . . . . . . . . . . . . . . . . . . . . . \+
open a file for  output. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \openout^{*}
close an  output file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \closeout^{*}
allocate a new  output file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newwrite
character that starts a new  output line in a write statement. . . . . . . . \newlinechar^{*}
escape character in the  output of control sequence tokens. . . . . . . \escapechar^{*}

trace  output processing. . . . . . . . . . . . . . . . . . . . . . . \showoutput^{L}
define the page output routine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \output^{*}
default Plain  output routine. . . . . . . . . . . . . . . . . . . . . . . . . \plainoutput
upper bound on  output routine calls. . . . . . . . . . . . . . . . . . \maxdeadcycles^{*}
number of output routine calls since last \shipout. \deadcycles^{*}
oval in a picture environment. . . . . . . . . . . . . . .  \oval^{L}
draw a line  over a formula. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \overline^{*}
horizontal brace  over a math formula. . . . . . . . . . . . . . . . . . . . . . \overbrace
left arrow  over a math formula. . . . . . . . . . . . . . . . . \overleftarrow
right arrow  over a math formula. . . . . . . . . . . . . . . . \overrightarrow
inhibit a page break  over a region. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \samepage^{L}
put symbols  over math relations. . . . . . . . . . . . . . . . . . . . . . . . \buildrel
put an accent  over the next character. . . . . . . . . . . . . . . . . . . . . \accent^{*}
place an accent  over the next math field. . . . . . . . . . . . . . . . . \mathaccent^{*}
width of rules appended to  overfull boxes. . . . . . . . . . . . . . . . . . . . . . . . . \overfullrule^{*}
maximum overrun before  overfull hbox messages occur. . . . . . . . . . . . . . .  \hfuzz^{*}
maximum overrun before  overfull vbox error. . . . . . . . . . . . . . . . . . . . . . . . . .  \vfuzz^{*}
override \global specifications. . . . . . . . . . \globaldefs^{*}
maximum  overrun before overfull hbox messages occur.  \hfuzz^{*}
maximum  overrun before overfull vbox error. . . . . . . . . .  \vfuzz^{*}
math relation:  owns (∋). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \owns
math symbol: Weierstrass  p (℘). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \wp
name of current TEX format  package. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fmtname
version of current TEX format  package. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fmtversion
additional authors on title  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \and^{L}
figures and tables and start a new right-hand  page.  flush . . . . . . . . . . . . . . . . . . . . \cleardoublepage^{L}
flush figures and start a new  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \clearpage^{L}
date on title page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \date^{L}
size of float on double-column  page. . . . . . . . . . . . . . . . . . . . . . . . \dblfloatpagefraction^{L}
flush insertions and eject to a new  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \dosupereject
horizontal offset of a  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \hoffset^{*}
sum of penalties for split insertions on the  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \insertpenalties^{*}
vertical space around a float in the middle of a  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \intextsep^{L}
insert at current position in  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \midinsert
create a small sample  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \minipage^{L}
start a new  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newpage^{L}
produce a vertical list of the body of a  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pagecontents
depth of the current  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pagedepth^{*}
amount of filll space in current  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pagefilllstretch^{*}
amount of fill space in current  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pagefillstretch^{*}
amount of fil space in current  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pagefilstretch^{*}
insert a whole  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pageinsert
amount of glue shrinkage in current  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pageshrink^{*}
amount of glue stretch in current  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pagestretch^{*}
put marginal notes on normal side of  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \normalmarginpar^{L}
put marginal notes on opposite side of  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \reversemarginpar^{L}
terminate L^{A}TEX and flush the final page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \stop^{L}
flush all insertions and eject to a new  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \supereject
width of printing on  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \textwidth^{L}
add footnote to title  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \thanks^{L}
insertion class for inserts at the top of a  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \topins
insert text at the top of the  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \topinsert
extra space added to top of  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \topmargin^{L}

space at the top of a  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \topskip^{*}
vertical offset of a  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \voffset^{*}
penalty for creating a widow line at top of  page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \widowpenalty^{*}
define the rule separating a  page and footnotes. . . . . . . . . . . . . . . . . . . . \footnoterule
title  page author. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \author^{L}
justify  page bottoms to the same height. . . . . . . \flushbottom^{L}
justify  page bottoms to the same height. . . . . . \normalbottom
justify  page bottoms to their natural height. . \raggedbottom^{L}
large vertical space or a good  page break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigbreak
force a page break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \eject
good  page break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \goodbreak
medium vertical space or a good  page break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \medbreak
prohibit a line or  page break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \nobreak
suppress a  page break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \nopagebreak^{L}
penalty at the current  page break. . . . . . . . . . . . . . . . . . . . . . . . . . . \outputpenalty^{*}
encourage a  page break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pagebreak^{L}
specify penalty for a line or  page break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \penalty^{*}
small vertical space and a good  page break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \smallbreak
penalty if page break after first line of paragraph. \clubpenalty^{*}
penalty if  page break after hyphenated line. . . . . \brokenpenalty^{*}
additional penalty for  page break between lines. . . . . . . . . \interlinepenalty^{*}
penalty for  page break just after a display.  \postdisplaypenalty^{*}
penalty for  page break just before a display.  \predisplaypenalty^{*}
inhibit a  page break over a region. . . . . . . . . . . . . . . . . . . \samepage^{L}
require strict line and page breaks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fussy^{L}
accept looser line and  page breaks. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \sloppy^{L}
put  page contents in a box of the proper height. \pagebody
constructs a box with the  page foot. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \makefootline
page foot line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footline
height of  page footer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footheight^{L}
space between text and  page footer. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footskip^{L}
fraction of two-column  page for top floats. . . . . . . . . . . . . . . . . . \dbltopfraction^{L}
page head line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \headline
height of  page header. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \headheight^{L}
constructs a box with the  page header. . . . . . . . . . . . . . . . . . . . . . . . . . . \makeheadline
space between  page header and text. . . . . . . . . . . . . . . . . . . . . . . \headsep^{L}
produces current left  page heading. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leftmark^{L}
produces current right  page heading. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \rightmark^{L}
desired  page height. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pagegoal^{*}
start a new  page in double-column format. . . . . . . . . . . . \twocolumn^{L}
start a new  page in single-column format. . . . . . . . . . . . . \onecolumn^{L}
define where text will be inserted when the  page is formatted. . . . . . . . . . . . . . . . . . . . . . . . . . . \insert^{*}
mark text last encountered on  page just boxed. . . . . . . . . . . . . . . . . . . . . . . . . . . . \botmark^{*}
mark text first encountered on  page just boxed. . . . . . . . . . . . . . . . . . . . . . . . . . \firstmark^{*}
typeset page number. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \folio
current  page number. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pageno
current  page number. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \thepage^{L}
page number of a cross reference label. . . . . . \pageref^{L}
specify  page numbering style. . . . . . . . . . . . . . . . \pagenumbering^{L}
suppress  page numbers. . . . . . . . . . . . . . . . . . . . . . . . \nopagenumbers
height of text on a page or \vbox. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vsize^{*}
define the  page output routine. . . . . . . . . . . . . . . . . . . . . . . . \output^{*}
natural height of  page so far. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pagetotal^{*}

set global  page style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pagestyle^{L}
set current  page style. . . . . . . . . . . . . . . . . . . . . . . . . . . \thispagestyle^{L}
suppress bottom justification of  page text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \raggedbottom
floats.  portion of  page that may be occupied by . . \floatpagefraction^{L}
below.  break a  page unless there is a better \filbreak … \filbreak
value of \botmark just before current  page was boxed. . . . . . . . . . . . . . . . . . . . . . . . . . . . \topmark^{*}
show  page-break calculations. . . . . . . . . . . . . . . \tracingpages^{*}
advance  \pageno by one. . . . . . . . . . . . . . . . . . . . . . \advancepageno
maximum depth of boxes on explicit  pages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \boxmaxdepth^{*}
left hand margin on even  pages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \evensidemargin^{L}
set headings for left and right  pages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \markboth^{L}
set heading for right  pages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \markright^{L}
maximum depth of boxes on main  pages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \maxdepth^{*}
left hand margin on odd  pages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \oddsidemargin^{L}
penalty if page break after first line of  paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \clubpenalty^{*}
hanging indent of a  paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \@hangfrom^{L}
force a line break in a paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \\^{L}
space to the left of a  paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leftskip^{*}
amount added to badness of every line in a  paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \linepenalty^{*}
force a change to the number of lines in a  paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \looseness^{*}
force a line break in a  paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newline^{L}
start a labeled  paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \paragraph^{L}
space at the end of the last line of a  paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \parfillskip^{*}
extra vertical space when environment starts a  paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \partopsep^{L}
number of lines in the last  paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \prevgraf^{*}
space to the right of a  paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \rightskip^{*}
end a paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \par^{*}
start a labeled sub-level  paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . \subparagraph^{L}
space between first list item and preceding  paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \topsep^{L}
insert vertical material into a  paragraph. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vadjust^{*}
tokens to insert when a  paragraph begins. . . . . . . . . . . . . . . . . . . . . . . . . . \everypar^{*}
indent second and subsequent lines in a paragraph by \parindent. . . . . . . . . . . . . . . . . . \hang
put a  paragraph in a box. . . . . . . . . . . . . . . . . . . . . . . . . \parbox^{L}
define  paragraph indentation. . . . . . . . . . . . . . . . . . . . \parindent^{*}
start a  paragraph indented \parindent. . . . . . . . . . . . \indent^{*}
suppress right justification of  paragraph lines. . . . . . . . . . . . . . . . . . . . . . . . \raggedright
width of a  paragraph or \hbox. . . . . . . . . . . . . . . . . . . . . . . .  \hsize^{*}
math symbol:  paragraph (¶). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \P
define an unusual  paragraph shape. . . . . . . . . . . . . . . . . . . . . . . . . . . \parshape^{*}
start a  paragraph without indentation. . . . . . . . . . . . . \noindent^{*}
define space between  paragraphs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \parskip^{*}
the following macro may have multiple paragraphs as parameters. . . . . . . . . . . . . . . . . . \long^{*}
indent second and subsequent  paragraphs in a list environment. . . . . \listparindent^{L}
environment. space between paragraphs within an item in a list . . . . . . . . . \parsep^{L}
math relation:  parallel (∥). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \parallel
set a font-related  parameter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fontdimen^{*}
following macro may have multiple paragraphs as parameters. the . . . . . . . . . . . . . . . . . . . . . . . . \long^{*}
fraction without a rule with  parentheses. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \choose
math strut with height and depth of  parentheses. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathstrut
generate a matrix with  parentheses delimiters. . . . . . . . . . . . . . . . . . . . . . \pmatrix
math function: mod within  parentheses ( (mod )). . . . . . . . . . . . . . . . . . . . .  \pmod
second and subsequent lines in a paragraph by  \parindent.  indent . . . . . . . . . . . . . . . . . . . .  \hang
start a paragraph indented  \parindent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \indent^{*}

increase left and right margins by  \parindent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \narrower
math symbol:  partial (∂). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \partial
define a set of hyphenation  patterns. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \patterns^{*}
pause after each line is read from a file. . . . . \pausing^{*}
processing TEX input.  pause for normal errors while . . . . . . . . \errorstopmode^{*}
process TEX input without  pausing for normal errors. . . . . . . . . . . . . . . . \scrollmode^{*}
page. sum of penalties for split insertions on the \insertpenalties^{*}
penalty at the beginning of a list.  \@beginparpenalty^{L}
penalty at the current page break. … \outputpenalty^{*}
penalty at the end of a list. . . . . . . . . \@endparpenalty^{L}
penalty between list items. . . . . . . . . . . . \@itempenalty^{L}
specify  penalty for a line or page break. . . . . . . . . . . . . \penalty^{*}
between lines of a display.  penalty for breaking . . . . . \interdisplaylinepenalty
between lines of a footnote.  penalty for breaking …. \interfootnotelinepenalty
of page.  penalty for creating a widow line at top \widowpenalty^{*}
before a display.  penalty for creating a widow line \displaywidowpenalty^{*}
interline  penalty for footnotes… \interfootnotelinepenalty^{L}
penalty for insertions that are split. \floatingpenalty^{*}
operation.  penalty for line break after binary …. \binoppenalty^{*}
discretionary hyphen.  penalty for line break after . . . . . . . . . . \hyphenpenalty^{*}
hyphen. penalty for line break after explicit \exhyphenpenalty^{*}
penalty for line break after math relation. \relpenalty^{*}
lines. additional penalty for page break between . . \interlinepenalty^{*}
a display.  penalty for page break just after \postdisplaypenalty^{*}
a display.  penalty for page break just before \predisplaypenalty^{*}
paragraph.  penalty if page break after first line of . \clubpenalty^{*}
line.  penalty if page break after hyphenated \brokenpenalty^{*}
remove a  penalty just added to the current list. . . . . \unpenalty^{*}
fetch last  penalty off the current list. . . . . . . . . . . . . \lastpenalty^{*}
demerits for a  penultimate broken line. . . . . . . \finalhyphendemerits^{*}
percent sign (%). . . . . . . . . . . . . . . . . . . . . . . . . . . .  \%
perform a \read or \write immediately. . \immediate^{*}
perform arithmetic on a register. . . . . . . . . . . . \advance^{*}
largest  permissible dimension. . . . . . . . . . . . . . . . . . . . . \maxdimen
permit a line break after a slash. . . . . . . . . . . .  \slash
column.  permit an alignment entry to stick out of its \hidewidth
math relation:  perpendicular (⊥). . . . . . . . . . . . . . . . . . . . . . . . . .  \perp
math Greek letter:  phi (ϕ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \phi
math Greek letter: capital  phi (Φ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Phi
math Greek letter: variant  phi (φ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \varphi
math Greek letter:  pi (π). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \pi
math Greek letter: capital  pi (Π). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Pi
math Greek letter: variant  pi (ϖ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \varpi
circle in a  picture environment. . . . . . . . . . . . . . . . . . . . . . . . \circle^{L}
end  picture environment. . . . . . . . . . . . . . . . . . . . . \endpicture^{L}
line in a picture environment. . . . . . . . . . . . . . . . . . . . . . . . \line^{L}
set width of lines in  picture environment. . . . . . . . . . . . . . . . . \linethickness^{L}
oval in a picture environment. . . . . . . . . . . . . . . . . . . . . . . . \oval^{L}
begin  picture environment. . . . . . . . . . . . . . . . . . . . . . . . \picture^{L}
unit of distance in  picture environment. . . . . . . . . . . . . . . . . . . . . \unitlength^{L}
vector in a  picture environment. . . . . . . . . . . . . . . . . . . . . . . . \vector^{L}
place several copies of a  picture object. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \multiput^{L}
place a  picture object. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \put^{L}

lower left  piece of a horizontal brace. . . . . . . . . . . . . . . . . . \braceld
upper left  piece of a horizontal brace. . . . . . . . . . . . . . . . . . \bracelu
lower right  piece of a horizontal brace. . . . . . . . . . . . . . . . . . \bracerd
upper right  piece of a horizontal brace. . . . . . . . . . . . . . . . . . \braceru
piece of a vertical brace (|). . . . . . . . . . . . . . \bracevert
internal Plain TEX command to  piece together long arrows. . . . . . . . . . . . . . . . . . \joinrel
character placed at the right end of an input line. \endlinechar^{*}
restore default limit  placement on large math operators… \displaylimits^{*}
default  Plain output routine. . . . . . . . . . . . . . . . . . . \plainoutput
arrows.  internal  Plain TEX command to piece together long . \joinrel
symbols.  internal  Plain TEX operation to define math text  \mathhexbox
internal  Plain TEX space used for centering. . . . . . . \centering
special  Plain TEX space used in alignment. . . . . . . . . \hideskip
math symbol: Planck’s constant or h-bar (h̄). . . . . . . . . . . . . . \hbar
math operator: large circle with  plus (⨁). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigoplus
math operator: large U  plus (⨄). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \biguplus
math operator: minus  plus (∓). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \mp
math operator: direct sum, circle  plus (⊕). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \oplus
math operator:  plus or minus (±). . . . . . . . . . . . . . . . . . . . . . . . . .  \pm
math operator: U  plus (⊎). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \uplus
Polish letter: slashed L (ł). . . . . . . . . . . . . . . . . .  \l
Polish letter: upper-case slashed L (Ł). . . . . .  \L
math function: degree of a polynomial (deg). . . . . . . . . . . . . . . . . . . . . . . . . . . \deg
occupied by floats.  portion of page that may be …. \floatpagefraction^{L}
create a box in dashes, with  positioning. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \dashbox^{L}
create and frame a box, with  positioning. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \framebox^{L}
create a box, with  positioning. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \makebox^{L}
create and name a box, with  positioning. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \savebox^{L}
capital letters.  positive if hyphenating words beginning with  \uchyph^{*}
(#).  pound, hatch mark, sharp sign, octothorpe .  \#
British  pound symbol (£). . . . . . . . . . . . . . . . . . . . . . . . . .  \pounds^{L}
suppress the template in the alignment  preamble for this entry. . . . . . . . . . . . . . . . . . . . .  \omit^{*}
math relation:  precedes or equal (⪯). . . . . . . . . . . . . . . . . . . . . . . \preceq
math relation: precedes (≺). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \prec
length of text  preceding a display. . . . . . . . . . . . . . . . . \predisplaysize^{*}
space between first list item and  preceding paragraph. . . . . . . . . . . . . . . . . . . . . . . . \topsep^{L}
scratch control sequence used in  preloading fonts. . . . . . . . . . . . . . . . . . . . . . . . . . \preloaded
vertical strut to  preserve line spacing. . . . . . . . . . . . . . . . . . . . . . .  \strut
math symbol:  prime (x′). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \prime
TEX  primitive \par. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \@@par^{L}
name of the  principal input file. . . . . . . . . . . . . . . . . . . . . . . . . . \jobname^{*}
width of  printing on page. . . . . . . . . . . . . . . . . . . . . . . . . . \textwidth^{L}
math function: probability (Pr). . . . . . . . . . . . . . . . . . . . . . . . . . . . \Pr
process input without displaying errors… \batchmode^{*}
process input without stopping for errors. \nonstopmode^{*}
normal errors.  process TEX input without pausing for . . \scrollmode^{*}
enable glossary  processing. . . . . . . . . . . . . . . . . . . . . . . . . . . . \makeglossary^{L}
trace output  processing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \showoutput^{L}
set a character’s  processing category type. . . . . . . . . . . . . . . . . . . . \catcode^{*}
set a math character’s  processing category type. . . . . . . . . . . . . . . . . . . \mathcode^{*}
finish processing input. . . . . . . . . . . . . . . . . . . . . . . . . . . . \bye
pause for normal errors while  processing TEX input. . . . . . . . . . . . . . . . \errorstopmode^{*}

produce a horizontal rule. . . . . . . . . . . . . . . . . . .  \hrule^{*}
produce a marginal note. . . . . . . . . . . . . . . . . . \marginpar^{L}
produce a numbered caption. . . . . . . . . . . . . . . . \caption^{L}
page. produce a vertical list of the body of a \pagecontents
produce a vertical rule. . . . . . . . . . . . . . . . . . . . . .  \vrule^{*}
produce footnote text without a mark.  \footnotetext^{L}
produce TEX accents in tabbing environment.  \a^{L}
produce the title. . . . . . . . . . . . . . . . . . . . . . . . . \maketitle^{L}
produce the value of a counter. . . . . . . . . . . . . .  \value^{L}
produces current left page heading. . . . . . . . . \leftmark^{L}
produces current right page heading. . . . . . \rightmark^{L}
produces the footnote number. . . . . . . . . . \thefootnote^{L}
math operator: tensor  product, circle times (⊗). . . . . . . . . . . . . . . . . . . \otimes
math operator: large  product (∏). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \prod
math binary operator: wreath  product (≀). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \wr
prohibit a line or page break. . . . . . . . . . . . . . . \nobreak
put page contents in a box of the  proper height. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pagebody
define a symbol that will work  properly in text and script sizes. . . . . . . . \mathpalette
math relation:  proportional to (∝). . . . . . . . . . . . . . . . . . . . . . . . . \propto
arguments.  protect fragile commands and moving . . . . . . \protect^{L}
math Greek letter:  psi (ψ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \psi
math Greek letter: capital  psi (Ψ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Psi
suppress special spacing after  punctuation. . . . . . . . . . . . . . . . . . . . . . . . . \frenchspacing
create end-of-sentence space after following punctuation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \@^{L}
enable special spacing after  punctuation. . . . . . . . . . . . . . . . . . . . . . \nonfrenchspacing
define a math  punctuation operator. . . . . . . . . . . . . . . . . . . . . \mathpunct^{*}
math symbol: exists  quantifier (∃). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \exists
math symbol: for-all  quantifier (∀). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \forall
math symbol: left  quote (‘). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \lq
right  quote (’). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \rq
math symbol: real, Fraktur  R (ℜ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Re
put a math field under a  radical. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \radical^{*}
raise a box a distance. . . . . . . . . . . . . . . . . . . . . . \raisebox^{L}
raise a box a given distance. . . . . . . . . . . . . . . .  \raise^{*}
1000.  ratio for variable delimiters times . \delimiterfactor^{*}
magnification ratio times 1000. . . . . . . . . . . . . . . . . . . . . . . . . . . . \mag^{*}
read a file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \input^{*}
read a file unless disabled by \includeonly.  \include^{L}
read a line from a file. . . . . . . . . . . . . . . . . . . . . . .  \read^{*}
non-space is found.  read, expand, then ignore tokens until a \ignorespaces^{*}
pause after each line is  read from a file. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pausing^{*}
processing.  read in the .AUX files and disable alpha @     \document^{L}
perform a  \read or \write immediately. . . . . . . . . . . . . \immediate^{*}
current line.  stop  reading current input file at the end of the  \endinput^{*}
math symbol:  real, Fraktur R (ℜ). . . . . . . . . . . . . . . . . . . . . . . . .  \Re
most  recently allocated register number. \allocationnumber
\dospecials. redefinable scratch control sequence used by \do
redefine a command. . . . . . . . . . . . . . . . . . . \renewcommand^{L}
redefine an environment. . . . . . . . . . \renewenvironment^{L}
refer to a cross reference label. . . . . . . . . . . . . .  \ref^{L}
generate an in-text citation of a reference. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \cite^{L}
reference a BibTEX item without citation… \nocite^{L}
increment and  reference a counter. . . . . . . . . . . . . . . . . \refstepcounter^{L}

begin and cross  reference an equation. . . . . . . . . . . . . . . . . . . . . . \equation^{L}
define a cross reference label. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \label^{L}
page number of a cross  reference label. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pageref^{L}
refer to a cross reference label. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ref^{L}
inhibit a page break over a  region. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \samepage^{L}
perform arithmetic on a  register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \advance^{*}
use a count register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \count^{*}
define a name for a count  register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \countdef^{*}
use a dimension register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \dimen^{*}
define a name for a dimension  register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \dimendef^{*}
math skip  register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \muskip^{*}
define a name for a math skip  register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \muskipdef^{*}
allocate a new box  register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newbox
allocate a new count  register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newcount
allocate a new dimension  register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newdimen
define a new box  register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newsavebox^{L}
store an hbox or vbox in a box  register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \setbox^{*}
display the contents of a  register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \showthe^{*}
use a skip register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \skip^{*}
define a name for a skip  register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \skipdef^{*}
allocate a new insert  register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newinsert
allocate a new math skip  register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newmuskip
allocate a new skip  register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newskip
allocate a new token  register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newtoks
use a token list register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \toks^{*}
define a name for a token list  register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \toksdef^{*}
divide a  register by a value. . . . . . . . . . . . . . . . . . . . . . . . . . \divide^{*}
multiply a  register by a value. . . . . . . . . . . . . . . . . . . . . . . . . \multiply^{*}
most recently allocated  register number. . . . . . . . . . . . . . . . . . \allocationnumber
the following macro definition or  register setting is global. . . . . . . . . . . . . . . . . . . . \global^{*}
convert a numeric register to displayable form. . . . . . . . . . . . . . . . . \the^{*}
penalty for line break after math  relation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \relpenalty^{*}
math  relation: approximately equal (≈). . . . . . . . . . \approx
math relation: asymptote (≍). . . . . . . . . . . . . . . . . . . . \asymp
math symbol:  relation bar (−). . . . . . . . . . . . . . . . . . . . . . . . . . . . \relbar
math symbol: double  relation bar (=). . . . . . . . . . . . . . . . . . . . . . . . . . . . \Relbar
math  relation: bowtie (▷◁). . . . . . . . . . . . . . . . . . . . . . . . \bowtie
math  relation: congruent (≅). . . . . . . . . . . . . . . . . . . . .  \cong
math  relation: contains (∋). . . . . . . . . . . . . . . . . . . . . . .  \ni
math  relation: dash V (⊣). . . . . . . . . . . . . . . . . . . . . . . .  \dashv
math  relation: dash (⊢). . . . . . . . . . . . . . . . . . . . . . . . . .  \vdash
math  relation: dotted equal (≐). . . . . . . . . . . . . . . . . .  \doteq
math  relation: equivalence (≡). . . . . . . . . . . . . . . . . . .  \equiv
math  relation: frown (⌢). . . . . . . . . . . . . . . . . . . . . . . .  \frown
math  relation: gets (←). . . . . . . . . . . . . . . . . . . . . . . . . .  \gets
math  relation: greater or equal (≥). . . . . . . . . . . . . . .  \ge
math  relation: greater or equal (≥). . . . . . . . . . . . . . .  \geq
math  relation: if and only if ( ⟺ ). . . . . . . . . . . . . .  \iff
math  relation: in (∈). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \in
math  relation: less or equal (≤). . . . . . . . . . . . . . . . . .  \le
math  relation: less or equal (≤). . . . . . . . . . . . . . . . . .  \leq
math  relation: mid (|). . . . . . . . . . . . . . . . . . . . . . . . . . . .  \mid
math  relation: models (⊧). . . . . . . . . . . . . . . . . . . . . . .  \models

math  relation: much greater (≫). . . . . . . . . . . . . . . . .  \gg
math  relation: much less (≪). . . . . . . . . . . . . . . . . . . .  \ll
math  relation: not equal (≠). . . . . . . . . . . . . . . . . . . . .  \ne
math  relation: not equal (≠). . . . . . . . . . . . . . . . . . . . .  \neq
math  relation: not equal (≠). . . . . . . . . . . . . . . . . . . . .  \not=
define a math  relation operator. . . . . . . . . . . . . . . . . . . . . . . . . . . \mathrel^{*}
math relation: owns (∋). . . . . . . . . . . . . . . . . . . . . . . . . . \owns
math  relation: parallel (∥). . . . . . . . . . . . . . . . . . . . . . . \parallel
math  relation: perpendicular (⊥). . . . . . . . . . . . . . . . .  \perp
math  relation: precedes or equal (⪯). . . . . . . . . . . . .  \preceq
math  relation: precedes (≺). . . . . . . . . . . . . . . . . . . . . .  \prec
math  relation: proportional to (∝). . . . . . . . . . . . . . .  \propto
math  relation: similar or equal (≃). . . . . . . . . . . . . . .  \simeq
math  relation: similar (∼). . . . . . . . . . . . . . . . . . . . . . . .  \sim
math  relation: smile (⌣). . . . . . . . . . . . . . . . . . . . . . . . .  \smile
math  relation: square subset or equal (⊑). . . . . \sqsubseteq
math  relation: square superset or equal (⊒)… \sqsupseteq
math  relation: subset or equal (⊆). . . . . . . . . . . . . . . \subseteq
math  relation: subset (⊂). . . . . . . . . . . . . . . . . . . . . . . . \subset
math  relation: successor or equal (⪰). . . . . . . . . . . . . \succeq
math relation: successor (≻). . . . . . . . . . . . . . . . . . . . . \succ
math  relation: superset or equal (⊇). . . . . . . . . . . . . \supseteq
math  relation: superset (⊃). . . . . . . . . . . . . . . . . . . . . . \supset
math relation: to (→). . . . . . . . . . . . . . . . . . . . . . . . . . . . \to
put symbols over math  relations. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \buildrel
remove a kern just added to the current list.  \unkern^{*}
list.  remove a penalty just added to the current \unpenalty^{*}
remove a skip just added to the current list.  \unskip^{*}
remove last skip on the list. . . . . . . . . \removelastskip
surround a space with a  repeated box or rule. . . . . . . . . . . . . . . . . . . . . . . \cleaders^{*}
fill a space with a  repeated box or rule. . . . . . . . . . . . . . . . . . . . . . . . \leaders^{*}
define a macro with expanded  replacement text. . . . . . . . . . . . . . . . . . . . . . . . . . .  \edef^{*}
define a global macro with expanded  replacement text. . . . . . . . . . . . . . . . . . . . . . . . . . .  \xdef^{*}
require strict line and page breaks. . . . . . . . . .  \fussy^{L}
reset tabs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \cleartabs
math operators.  restore default limit placement on large \displaylimits^{*}
restore tabs stops in tabbing environment. . \poptabs^{L}
meaning.  restores a carriage return to its usual . . . . . \restorecr^{L}
defines a carriage  return as \\. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \obeycr^{L}
restores a carriage  return to its usual meaning. . . . . . . . . . . . . . . \restorecr^{L}
math Greek letter: rho (ρ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \rho
math Greek letter: variant  rho (ϱ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \varrho
add a box to the vertical list shifted  right. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \moveright^{*}
flush lines  right. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \raggedleft^{L}
create a zero-width box with text to the right. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \rlap
math delimiter:  right angle bracket (⟩). . . . . . . . . . . . . . . . . . . . . . \rangle
fill a space with a  right arrow. . . . . . . . . . . . . . . . . . . . . . . . . \rightarrowfill
math symbol: hook  right arrow (↪). . . . . . . . . . . . . . . . . . . . \hookrightarrow
math symbol: long left and  right arrow (⟷). . . . . . . . . . . . . . \longleftrightarrow
math symbol: long  right arrow (⟶). . . . . . . . . . . . . . . . . . . \longrightarrow
right arrow over a math formula. … \overrightarrow
math symbol:  right arrow (→). . . . . . . . . . . . . . . . . . . . . . . . . \rightarrow
math delimiter:  right bracket (]). . . . . . . . . . . . . . . . . . . . . . . . . . . . \rbrack

math delimiter:  right ceiling bracket (⌉). . . . . . . . . . . . . . . . . . . . .  \rceil
math delimiter:  right curly brace (}). . . . . . . . . . . . . . . . . . . . . . . . \rbrace
end a math list with a right delimiter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \right^{*}
math symbol: long  right double arrow (⟹). . . . . . . . . . . \Longrightarrow
math symbol:  right double arrow (⇒). . . . . . . . . . . . . . . . . . \Rightarrow
character placed at the  right end of an input line. . . . . . . . . . . . . . . \endlinechar^{*}
end flush  right environment. . . . . . . . . . . . . . . . . . . . \endflushright^{L}
begin flush  right environment. . . . . . . . . . . . . . . . . . . . . . . \flushright^{L}
math delimiter: right floor (⌋). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \rfloor
math delimiter:  right group (}). . . . . . . . . . . . . . . . . . . . . . . . . . . . \rgroup
math symbol:  right half diamond (⊳). . . . . . . . . . . . . . . . . . . . .  \rhd^{L}
math symbol: underlined  right half diamond (⊵). . . . . . . . . . . . . . . . . . . . .  \unrhd^{L}
math symbol:  right harpoon down (⇁). . . . . . . . . \rightharpoondown
math symbol:  right harpoon up (⇀). . . . . . . . . . . . . . \rightharpoonup
math symbol:  right hook ( ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \rhook
put text flush  right in a column in tabbing environment…  \'^{L}
suppress  right justification of paragraph lines. … \raggedright
lines. suppress right justification of typewriter font . \ttraggedright
math symbol:  right left harpoon (⇀ ↽). . . . . . . . . . \rightleftharpoons
right margin in a list. . . . . . . . . . . . . . . . . . . \rightmargin^{L}
increase left and  right margins by \parindent. . . . . . . . . . . . . . \narrower
2.5-line  right math delimiter. . . . . . . . . . . . . . . . . . . . . . . .  \Biggr
2-line  right math delimiter. . . . . . . . . . . . . . . . . . . . . . . .  \biggr
1.5-line  right math delimiter. . . . . . . . . . . . . . . . . . . . . . . .  \Bigr
1-line  right math delimiter. . . . . . . . . . . . . . . . . . . . . . . .  \bigr
math symbol:  right moustache (}). . . . . . . . . . . . . . . . . . . . \rmoustache
space to the  right of a paragraph. . . . . . . . . . . . . . . . . . . . . . \rightskip^{*}
produces current  right page heading. . . . . . . . . . . . . . . . . . . . . . . . \rightmark^{L}
set headings for left and  right pages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \markboth^{L}
set heading for  right pages. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \markright^{L}
lower  right piece of a horizontal brace. . . . . . . . . . . . \bracerd
upper  right piece of a horizontal brace. . . . . . . . . . . . \braceru
right quote (’). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \rq
flush  right text on a line. . . . . . . . . . . . . . . . . . . . . . . \rightline
math operator:  right triangle (▷). . . . . . . . . . . . . . . . . . . . . \triangleright
flush figures and tables and start a new  right-hand page. . . . . . . . . . . . . . . . . . . \cleardoublepage^{L}
8 point  Roman font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \egtrm^{L}
11 point  Roman font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvrm^{L}
5 point boldface  Roman font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fivebf
5 point  Roman font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fiverm
5 point Roman font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fivrm^{L}
14 point  Roman font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \frtnrm^{L}
9 point  Roman font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ninrm^{L}
select  Roman font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \rm
7 point bold  Roman font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \sevenbf
7 point  Roman font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \sevenrm
7 point  Roman font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sevrm^{L}
6 point  Roman font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sixrm^{L}
17 point  Roman font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \svtnrm^{L}
10 point  Roman font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tenrm
12 point  Roman font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twlrm^{L}
20 point  Roman font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \twtyrm^{L}

display counter as lower-case  Roman numerals. . . . . . . . . . . . . . . . . . . . . . . . . . .  \roman^{L}
display counter as upper-case  Roman numerals. . . . . . . . . . . . . . . . . . . . . . . . . . .  \Roman^{L}
convert a number to lower-case  Roman numerals. . . . . . . . . . . . . . . . . . . . . . \romannumeral^{*}
specified  root of a formula. . . . . . . . . . . . . . . . . . . . . . . . . . .  \root
square  root of a formula. . . . . . . . . . . . . . . . . . . . . . . . . . .  \sqrt
define the page output  routine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \output^{*}
default Plain output  routine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \plainoutput
upper bound on output  routine calls. . . . . . . . . . . . . . . . . . . . . . . . . \maxdeadcycles^{*}
number of output  routine calls since last \shipout. . . . . . . . . \deadcycles^{*}
generate a matrix labeled on  rows and columns. . . . . . . . . . . . . . . . . . . . . \bordermatrix
space between  rows of array or tabular environment. . \arraystretch^{L}
fraction without rule. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \atop^{*}
surround a space with a repeated box or  rule. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \cleaders^{*}
produce a horizontal rule. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \hrule^{*}
fill a space with a  rule. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \hrulefill
fill a space with a repeated box or  rule. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \leaders^{*}
fraction with a  rule. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \over^{*}
produce a vertical  rule. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vrule^{*}
fill a space with an evenly distributed box or  rule. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \xleaders^{*}
fraction with specified  rule and delimiters. . . . . . . . . . . . . . . . \abovewithdelims^{*}
fraction with  rule and given delimiters. . . . . . . . . . . \overwithdelims^{*}
rule between double-column floats and text. \dblfigrule^{L}
place  rule between text and bottom floats. . . . . \botfigrule^{L}
place  rule between top floats and text. . . . . . . . . \topfigrule^{L}
create a rule or line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \rule^{L}
define the  rule separating a page and footnotes. . \footnoterule
width of  rule separating double columns. . . . . . \columnseprule^{L}
macro to draw the  rule separating footnotes from text. … \footnoterule^{L}
environment. double rule separation in array or tabular … \doublerulesep^{L}
thickness of  rule surrounding framed box. . . . . . . . . . . . . . . \fboxrule^{L}
fraction with  rule thickness. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \above^{*}
fraction without a  rule with braces. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \brace
fraction without a  rule with brackets. . . . . . . . . . . . . . . . . . . . . . . . . .  \brack
fraction without  rule with given delimiters. . . . . . . . . . . \atopwithdelims^{*}
fraction without a  rule with parentheses. . . . . . . . . . . . . . . . . . . . . . . \choose
set width of array  rules. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \arrayrulewidth^{L}
width of  rules appended to overfull boxes. . . . . . \overfullrule^{*}
German letter: sharp s (ß). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ss
create a small  sample page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \minipage^{L}
11 point  sans serif font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvsf^{L}
select  sans serif font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sf^{L}
10 point  sans serif font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tensf^{L}
12 point  sans serif font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twlsf^{L}
environment.  save current tab stops in tabbing . . . . . . . . . . \pushtabs^{L}
Scandinavian letter: a with circle (å). . . . . . .  \aa
Scandinavian letter: capital A with circle (Å).  \AA
\dospecials. redefinable scratch control sequence used by . . . . . . . . . . . \do
fonts.  scratch control sequence used in preloading \preloaded
suppress space in  script and small script styles. . . . . . . . . . . . . . \nonscript^{*}
math symbol: small  script L (ℓ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ell
a symbol that will work properly in text and  script sizes.  define . . . . . . . . . . . . . . . . . . \mathpalette
use normal math  script style. . . . . . . . . . . . . . . . . . . . . . . . . . . . \displaystyle^{*}
use very small math  script style. . . . . . . . . . . . . . . . . . . . . . \scriptscriptstyle^{*}

use small math  script style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \scriptstyle^{*}
suppress space in script and small  script styles. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \nonscript^{*}
select font for small math  scripts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \scriptfont^{*}
select font for very small math  scripts. . . . . . . . . . . . . . . . . . . . . . . . . . . \scriptscriptfont^{*}
math function: secant. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \sec
current value of the  second-level item counter. . . . . . . . . . . . . . . . . \theenumii^{L}
second-level item in an index. . . . . . . . . . . . . . . \subitem^{L}
mark denoting  second-level items. . . . . . . . . . . . . . . . . . . . . . \labelitemii^{L}
width of left margin in  second-level list. . . . . . . . . . . . . . . . . . . . . . . \leftmarginii^{L}
start a  section. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \section^{L}
command.  includes the  section number in a \contentsline . . . . . \numberline^{L}
math symbol:  section (§). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \S
set  sectional units to appendix style. . . . . . . . . . . \appendix^{L}
select a font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newfont^{L}
select bold extended font. . . . . . . . . . . . . . . . . . .  \bf
select bold math italic and symbol fonts… \boldmath^{L}
select font for non-math text. . . . . . . . . . . . . . . \textfont^{*}
select font for small math scripts. . . . . . . . \scriptfont^{*}
scripts.  select font for very small math … \scriptscriptfont^{*}
select font four steps larger than normal size.  \huge^{L}
select font one step larger than normal size.  \large^{L}
select font three steps larger than normal size.  \LARGE^{L}
select font two steps larger than normal size.  \Large^{L}
select footnote font size. . . . . . . . . . . . . . . \footnotesize^{L}
select italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \it
select largest available font. . . . . . . . . . . . . . . . .  \Huge^{L}
the next item.  select math spacing of a closing delimiter for \mathclose^{*}
for the next item. select math spacing of an opening delimiter \mathopen^{*}
select normal font size. . . . . . . . . . . . . . . . . . . \normalsize^{L}
select old-style numerals with descenders… \oldstyle
integer. select one of several entries determined by an \ifcase^{*}
select Roman font. . . . . . . . . . . . . . . . . . . . . . . . . .  \rm
select sans serif font. . . . . . . . . . . . . . . . . . . . . . . .  \sf^{L}
select slanted font. . . . . . . . . . . . . . . . . . . . . . . . . .  \sl
select small caps font. . . . . . . . . . . . . . . . . . . . . . .  \sc^{L}
select small font size. . . . . . . . . . . . . . . . . . . . . . . .  \small^{L}
select smallest defined font. . . . . . . . . . . . . . . . .  \tiny^{L}
select subscript or superscript font size. . \scriptsize^{L}
select sub-subscript font size. . . . . \scriptscriptsize^{L}
select typewriter font. . . . . . . . . . . . . . . . . . . . . . .  \tt
give a token list defining the  semantics of a token. . . . . . . . . . . . . . . . . . . . . . . . \meaning^{*}
send a box to the DVI file. . . . . . . . . . . . . . . . . . \shipout^{*}
space between  sentences. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \xspaceskip^{*}
separate cases in an \ifcase. . . . . . . . . . . . . . .  \or^{*}
define the rule  separating a page and footnotes. . . . . . . \footnoterule
environment. half the width separating columns in a tabular . . . . . . . . . . \tabcolsep^{L}
width of rule  separating double columns. . . . . . . . . . . \columnseprule^{L}
macro to draw the rule  separating footnotes from text. . . . . . . . \footnoterule^{L}
ensure footnote line  separation. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footstrut
footnote  separation distance. . . . . . . . . . . . . . . . . . . . . \footnotesep^{L}
column  separation in array environment. . . . . . . . \arraycolsep^{L}
environment. double rule separation in array or tabular . . . . . . . \doublerulesep^{L}
increase line  separation in math mode. . . . . . . . . . . . . . . . . . . \openup

field  separator in \halign or \valign. . . . . . . . . . .  &^{*}
a synonym for the current meaning of a control  sequence.  define . . . . . . . . . . . . . . . . . . . . . . . .  \let^{*}
an undefined control  sequence. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \undefined
assigns the second token to a control  sequence and continues. . . . . . . . . . . . . . . . . . . \futurelet^{*}
expand a control  sequence into character tokens. . . . . . . . . . . . . . \string^{*}
escape character in the output of control  sequence tokens. . . . . . . . . . . . . . . . . . . . . . . . . \escapechar^{*}
redefinable scratch control sequence used by \dospecials. . . . . . . . . . . . . \do
scratch control  sequence used in preloading fonts. . . . . . . . . \preloaded
11 point sans  serif font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvsf^{L}
select sans  serif font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sf^{L}
10 point sans  serif font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tensf^{L}
12 point sans  serif font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twlsf^{L}
end a  \settabs definition. . . . . . . . . . . . . . . . . . . . . . . . \columns
the following macro definition or register setting is global. . . . . . . . . . . . . . . . . . . . . . . . . . . . \global^{*}
define an unusual paragraph  shape. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \parshape^{*}
German letter:  sharp s (ß). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ss
math symbol:  sharp (♯). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sharp
pound, hatch mark,  sharp sign, octothorpe (#). . . . . . . . . . . . . . . . .  \#
shift super accents. . . . . . . . . . . . . . . . . . . . . . . . . .  \skew
add a box to the vertical list  shifted left. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \moveleft^{*}
add a box to the vertical list  shifted right. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \moveright^{*}
number of output routine calls since last  \shipout. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \deadcycles^{*}
show boxes that are  shipped out. . . . . . . . . . . . . . . . . . . . . . . . . . \tracingoutput^{*}
generate a short amount of verbatim text. . . . . . . . . . . . . . \verb^{L}
extra space above displays following  short lines. . . . . . . . . . . . . . . . . . \abovedisplayshortskip^{*}
extra space just below displays following  short lines. . . . . . . . . . . . . . . . . . \belowdisplayshortskip^{*}
show a box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \showbox^{*}
show a token. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \show^{*}
show boxes that are shipped out. . . . . \tracingoutput^{*}
show characters not in the font… \tracinglostchars^{*}
executed.  show commands before they are … \tracingcommands^{*}
show diagnostics on the terminal. . . . . \tracingonline^{*}
show hyphenations of given words. . . . . . \showhyphens
show line-break calculations. . . . . \tracingparagraphs^{*}
show macros as they are expanded… \tracingmacros^{*}
show page-break calculations. . . . . . . . . . \tracingpages^{*}
show statistics about memory usage… \tracingstats^{*}
show the current lists. . . . . . . . . . . . . . . . . . . . . \showlists^{*}
end.  show unassignments when groups . \tracingrestores^{*}
maximum box depth  shown. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \showboxdepth^{*}
maximum boxed items  shown at a given depth. . . . . . . . . . . . . \showboxbreadth^{*}
infinitely stretchable and  shrinkable horizontal space. . . . . . . . . . . . . . . . .  \hss^{*}
infinitely stretchable and  shrinkable vertical space. . . . . . . . . . . . . . . . . . . .  \vss^{*}
amount of glue  shrinkage in current page. . . . . . . . . . . . . . . . \pageshrink^{*}
put marginal notes on normal  side of page. . . . . . . . . . . . . . . . . . . . . . . \normalmarginpar^{L}
put marginal notes on opposite  side of page. . . . . . . . . . . . . . . . . . . . . . \reversemarginpar^{L}
math Greek letter:  sigma (σ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sigma
math Greek letter: capital  sigma (Σ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Sigma
math Greek letter: variant  sigma (ς). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \varsigma
dollar  sign. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \$
math operator: discretionary multiply  sign. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \*
percent  sign (%). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \%
pound, hatch mark, sharp  sign, octothorpe (#). . . . . . . . . . . . . . . . . . . . . . . .  \#

math relation:  similar or equal (≃). . . . . . . . . . . . . . . . . . . . . . . .  \simeq
math relation:  similar (∼). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sim
math function: arc  sine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \arcsin
math function:  sine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sin
math function: hyperbolic  sine. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sinh
create a box with a  single column of items. . . . . . . . . . . . . . . . . . . \shortstack^{L}
start a new page in  single-column format. . . . . . . . . . . . . . . . . . . . . \onecolumn^{L}
current value of the  sixth-level item counter. . . . . . . . . . . . . . . . . . . \theenumvi^{L}
mark denoting  sixth-level items. . . . . . . . . . . . . . . . . . . . . . . . \labelitemvi^{L}
width of left margin in  sixth-level list. . . . . . . . . . . . . . . . . . . . . . . . . \leftmarginvi^{L}
1.5-line math delimiter  size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Big
1-line math delimiter  size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \big
2.5-line math delimiter  size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Bigg
2-line math delimiter  size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \bigg
big font  size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \big^{L}
bigger font  size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Big^{L}
biggest font  size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \BIG^{L}
select footnote font  size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footnotesize^{L}
select font four steps larger than normal  size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \huge^{L}
select font one step larger than normal  size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \large^{L}
select font three steps larger than normal  size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \LARGE^{L}
select font two steps larger than normal  size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Large^{L}
select normal font  size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \normalsize^{L}
select sub-subscript font  size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \scriptscriptsize^{L}
select subscript or superscript font  size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \scriptsize^{L}
select small font size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \small^{L}
size of column that must contain text.  \textfraction^{L}
page.  size of float on double-column \dblfloatpagefraction^{L}
symbol that will work properly in text and script sizes. define a . . . . . . . . . . . . . . . . . . . . . . \mathpalette
breakable small horizontal  skip. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \enskip
large breakable horizontal  skip. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \qquad
breakable medium horizontal  skip. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \quad
vertical  skip a large amount. . . . . . . . . . . . . . . . . . . . . . . . \bigskip^{L}
vertical  skip a medium amount. . . . . . . . . . . . . . . . . . . . . \medskip^{L}
vertical  skip a small amount. . . . . . . . . . . . . . . . . . . . . . \smallskip^{L}
an unoriented skip amount with stretch of 1fill. . . . . . . . . . \fill^{L}
skip horizontal space. . . . . . . . . . . . . . . . . . . . . . . \hspace^{L}
remove a skip just added to the current list. . . . . . . . . . \unskip^{*}
fetch last  skip off the current list. . . . . . . . . . . . . . . . . . . . \lastskip^{*}
remove last  skip on the list. . . . . . . . . . . . . . . . . . . . . \removelastskip
math  skip register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \muskip^{*}
define a name for a math  skip register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \muskipdef^{*}
use a skip register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \skip^{*}
define a name for a  skip register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \skipdef^{*}
allocate a new math  skip register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newmuskip
allocate a new  skip register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newskip
skip vertical space. . . . . . . . . . . . . . . . . . . . . . . . . . \vspace^{L}
11 point  slanted font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvsl^{L}
select  slanted font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sl
10 point  slanted font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tensl
12 point  slanted font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twlsl^{L}
slanted font family. . . . . . . . . . . . . . . . . . . . . . . . . .  \slfam
permit a line break after a  slash. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \slash

Norwegian letter: capital O with  slash (Ø). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \O
Norwegian letter: o with  slash (ø). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \o
math operator: circle  slash (⊘). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \oslash
Polish letter:  slashed L (ł). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \l
Polish letter: upper-case  slashed L (Ł). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \L
vertical skip a  small amount. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \smallskip^{L}
select  small caps font. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sc^{L}
select  small font size. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \small^{L}
breakable  small horizontal skip. . . . . . . . . . . . . . . . . . . . . . .  \enskip
unbreakable  small horizontal space. . . . . . . . . . . . . . . . . . . . . . \enspace
unbreakable  small horizontal space. . . . . . . . . . . . . . . . . . . . \thinspace
math operator:  small integral (∫). . . . . . . . . . . . . . . . . . . . . . . . . . \smallint
use very  small math script style. . . . . . . . . . \scriptscriptstyle^{*}
use  small math script style. . . . . . . . . . . . . . . . . \scriptstyle^{*}
select font for  small math scripts. . . . . . . . . . . . . . . . . . . . . . . \scriptfont^{*}
select font for very  small math scripts. . . . . . . . . . . . . . . \scriptscriptfont^{*}
create a  small sample page. . . . . . . . . . . . . . . . . . . . . . . . . \minipage^{L}
math symbol: small script L (ℓ). . . . . . . . . . . . . . . . . . . . . . . . . . . \ell
suppress space in script and  small script styles. . . . . . . . . . . . . . . . . . . . . . . . \nonscript^{*}
small vertical space. . . . . . . . . . . . . . . . . . . . . . . \smallskip
small vertical space and a good page break. \smallbreak
select  smallest defined font. . . . . . . . . . . . . . . . . . . . . . .  \tiny^{L}

                \smallskip space. . . . . . . . . . . . . . . . . \smallskipamount

math relation:  smile (⌣). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \smile
create  some horizontal space. . . . . . . . . . . . . . . . . . . . . .  \hglue
put a frame around  some text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \frame^{L}
create  some vertical space. . . . . . . . . . . . . . . . . . . . . . . . .  \vglue
math symbol:  southeast arrow (↘). . . . . . . . . . . . . . . . . . . . . . . \searrow
math symbol:  southwest arrow (↙). . . . . . . . . . . . . . . . . . . . . . . \swarrow
math mode negative thin  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \!
unbreakable  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  ~
add extra vertical  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \addvspace^{L}
\bigskip  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigskipamount
unbreakable small horizontal  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \enspace
infinitely stretchable horizontal  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \hfil^{*}
more infinitely stretchable horizontal  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \hfill^{*}
cancel infinitely stretchable horizontal  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \hfilneg^{*}
create some horizontal  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \hglue
add horizontal  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \hskip^{*}
skip horizontal  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \hspace^{L}
infinitely stretchable and shrinkable horizontal space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \hss^{*}
medium vertical  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \medskip
\medskip  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \medskipamount
unbreakable tiny negative math mode horizontal  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \negthinspace
control space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \␣^{*}
small vertical  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \smallskip
\smallskip  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \smallskipamount
a blank space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \space
infinitely stretchable  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \stretch^{L}
unbreakable small horizontal  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \thinspace
infinitely stretchable vertical  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vfil^{*}
more infinitely stretchable vertical  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vfill^{*}
cancel infinitely stretchable vertical  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vfilneg^{*}

create some vertical  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vglue
add vertical  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vskip^{*}
skip vertical  space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vspace^{L}
infinitely stretchable and shrinkable vertical space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vss^{*}
extra  space above displays. . . . . . . . . . . . . \abovedisplayskip^{*}
short lines. extra space above displays following \abovedisplayshortskip^{*}
extra  space added to top of page. . . . . . . . . . . . . . . \topmargin^{L}
create end-of-sentence space after following punctuation. . . . . . . . . . . \@^{L}
extra  space after subscript or superscript. . . . . \scriptspace^{*}
small vertical  space and a good page break. . . . . . . . . . . . \smallbreak
vertical  space around a float in the middle of a page. \intextsep^{L}
paragraph.  space at the end of the last line of a … \parfillskip^{*}
space at the top of a page. . . . . . . . . . . . . . . . . . \topskip^{*}
space at top of a split box. . . . . . . . . . . . \splittopskip^{*}
environment. add extra space before a column in array or tabular \extracolsep^{L}
suppress interline  space before next vertical box. . . . . \nointerlineskip
and \framebox.  space between a box and its contents in \fbox   \fboxsep^{L}
space between a label and text of a list item. \labelsep^{L}
space between aligned tab entries. . . . . . . . . . . \tabskip^{*}
text.  space between columns in double column . \columnsep^{L}
space between double-column floats. … \dblfloatsep^{L}
and text.  space between double-column floats \dbltextfloatsep^{L}
amount of extra  space between entries in an index. . . . . . . . \indexspace^{L}
paragraph.  space between first list item and preceding . \topsep^{L}
space between floats. . . . . . . . . . . . . . . . . . . . . . . \floatsep^{L}
space between floats and the text. . . . . \textfloatsep^{L}
normal  space between lines. . . . . . . . . . . . . . . . . . . \baselineskip^{*}
turn off extra  space between lines. . . . . . . . . . . . . . \offinterlineskip
enough. extra space between lines if \baselineskip isn’t \lineskip^{*}
minimum  space between lines in a vertical list. . \lineskiplimit^{*}
space between main text and footnotes. . . . . \footins^{L}
vertical  space between marginal notes. . . . . . . . \marginparpush^{L}
space between page header and text. . . . . . . . \headsep^{L}
define  space between paragraphs. . . . . . . . . . . . . . . . . . \parskip^{*}
list environment.  space between paragraphs within an item in a  \parsep^{L}
environment.  space between rows of array or tabular  \arraystretch^{L}
space between sentences. . . . . . . . . . . . . . . . . \xspaceskip^{*}
environment.  space between successive items in a list . . . . . \itemsep^{L}
space between text and page footer. . . . . . . . \footskip^{L}
space between words. . . . . . . . . . . . . . . . . . . . . \spaceskip^{*}
space character. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  ␣^{*}
set a character’s  space factor. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \sfcode^{*}
extra  space in badly-stretched lines. . . . . \emergencystretch^{3}
amount of filll  space in current page. . . . . . . . . . . . . \pagefilllstretch^{*}
amount of fill  space in current page. . . . . . . . . . . . . . \pagefillstretch^{*}
amount of fil  space in current page. . . . . . . . . . . . . . . \pagefilstretch^{*}
medium  space in math formulas. . . . . . . . . . . . . . . . . . . \medmuskip^{*}
thick  space in math formulas. . . . . . . . . . . . . . . . . \thickmuskip^{*}
thin  space in math formulas. . . . . . . . . . . . . . . . . . \thinmuskip^{*}
medium  space in math mode. . . . . . . . . . . . . . . . . . . . . . . .  \>
medium  space in math mode. . . . . . . . . . . . . . . . . . . . . . . .  \:^{L}
insert horizontal  space in math mode. . . . . . . . . . . . . . . . . . . . . . . .  \mskip^{*}
suppress  space in script and small script styles. . . . . \nonscript^{*}

extra  space just below displays. . . . . . . . . \belowdisplayskip^{*}
following short lines.  extra  space just below displays … \belowdisplayshortskip^{*}
thick  space (math mode). . . . . . . . . . . . . . . . . . . . . . . . .  \;
thin  space (math mode). . . . . . . . . . . . . . . . . . . . . . . . .  \,
large vertical  space or a good page break. . . . . . . . . . . . . . . . \bigbreak
medium vertical  space or a good page break. . . . . . . . . . . . . . . . \medbreak
use the  space taken by a formula. . . . . . . . . . . . . . . . . . . \phantom
space to the left of a paragraph. . . . . . . . . . . . \leftskip^{*}
space to the right of a paragraph. . . . . . . . . \rightskip^{*}
internal Plain TEX  space used for centering. . . . . . . . . . . . . . . . . . \centering
special Plain TEX  space used in alignment. . . . . . . . . . . . . . . . . . . . \hideskip
extra vertical  space when environment starts a paragraph. \partopsep^{L}
fill a  space with a downward brace. . . . . . . . \downbracefill
fill a  space with a left arrow. . . . . . . . . . . . . . . \leftarrowfill
surround a  space with a repeated box or rule. . . . . . . . . . \cleaders^{*}
fill a  space with a repeated box or rule. . . . . . . . . . . \leaders^{*}
fill a  space with a right arrow. . . . . . . . . . . . \rightarrowfill
fill a  space with a rule. . . . . . . . . . . . . . . . . . . . . . . . . \hrulefill
fill a  space with an evenly distributed box or rule. \xleaders^{*}
fill a  space with an upward brace. . . . . . . . . . . . \upbracefill
fill a  space with dots. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \dotfill
verbatim  spaces in input text. . . . . . . . . . . . . . . . . . . . . \obeyspaces
lower dot math symbol with special  spacing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ldotp
vertical strut to preserve line  spacing. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \strut
suppress special  spacing after punctuation. . . . . . . . . . . . \frenchspacing
enable special  spacing after punctuation. . . . . . . . \nonfrenchspacing
set a character’s  spacing factor. . . . . . . . . . . . . . . . . . . . . . . . . . \spacefactor^{*}
centered dot with special  spacing in math mode. . . . . . . . . . . . . . . . . . . . . .  \cdotp
item.  select math  spacing of a closing delimiter for the next  \mathclose^{*}
item.  select math  spacing of an opening delimiter for the next \mathopen^{*}
set line  spacing to normal values. . . . . . . . . . \normalbaselines
math symbol:  spade suit (♠). . . . . . . . . . . . . . . . . . . . . . . . . . . . \spadesuit
amount by which delimiters can fail to  span included material. . . . . . . . . \delimitershortfall^{*}
span several columns in an alignment. . . . . \multispan
enable only  specific \include commands. . . . . . . . . . . \includeonly^{L}
override \global  specifications. . . . . . . . . . . . . . . . . . . . . . . . . . . . \globaldefs^{*}
specify a character by its numeric code. . . . .  \char^{*}
specify a math character code. . . . . . . . . . . . . . \mathchar^{*}
BibTEX.  specify bibliographic style for … \bibliographystyle^{L}
hyphenation.  specify language to be used for . . . . . . . . . \setlanguage^{3}
specify page numbering style. . . . . . . . . \pagenumbering^{L}
specify penalty for a line or page break. . . . . \penalty^{*}
penalty for insertions that are  split. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \floatingpenalty^{*}
mark text last encountered in a  split box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \splitbotmark^{*}
mark text first encountered in a  split box. . . . . . . . . . . . . . . . . . . . . . . . . . . \splitfirstmark^{*}
maximum depth of boxes in a  split box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \splitmaxdepth^{*}
space at top of a  split box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \splittopskip^{*}
sum of penalties for  split insertions on the page. . . . . . . . \insertpenalties^{*}
split off a specified amount from a vbox. …  \vsplit^{*}
math operator: square cap (⊓). . . . . . . . . . . . . . . . . . . . . . . . . . . . . \sqcap
math operator: large  square cup (⨆). . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigsqcup
math operator: square cup (⊔). . . . . . . . . . . . . . . . . . . . . . . . . . . . . \sqcup
square root of a formula. . . . . . . . . . . . . . . . . . . .  \sqrt

math relation:  square subset or equal (⊑). . . . . . . . . . . . . . \sqsubseteq
math symbol:  square subset (⊏). . . . . . . . . . . . . . . . . . . . . . . . . . \sqsubset^{L}
math relation:  square superset or equal (⊒). . . . . . . . . . . . \sqsupseteq
math symbol:  square superset (⊐). . . . . . . . . . . . . . . . . . . . . . . . \sqsupset^{L}
align a  stack of equations. . . . . . . . . . . . . . . . . . . . . . . . . . \eqalign
align a  stack of equations with equation numbers.  \eqalignno
numbers.  align a  stack of equations with left equation …. \leqalignno
display a  stack of formulas without alignment… \displaylines
stack one equation above another. . . . . . . . . . \stackrel^{L}
math operator: star (⋆). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \star
start a chapter. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \chapter^{L}
start a labeled paragraph. . . . . . . . . . . . . . . . . \paragraph^{L}
start a labeled sub-level paragraph. … \subparagraph^{L}
start a major division of a long document. .  \part^{L}
flush figures and  start a new page. . . . . . . . . . . . . . . . . . . . . . . . . . \clearpage^{L}
start a new page. . . . . . . . . . . . . . . . . . . . . . . . . . . \newpage^{L}
start a new page in double-column format. \twocolumn^{L}
start a new page in single-column format.  \onecolumn^{L}
flush figures and tables and  start a new right-hand page. . . . . . . \cleardoublepage^{L}
start a paragraph indented \parindent. . . . . \indent^{*}
start a paragraph without indentation. . . . . \noindent^{*}
start a section. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \section^{L}
start a subsection. . . . . . . . . . . . . . . . . . . . . . . \subsection^{L}
start a subsubsection. . . . . . . . . . . . . . . . \subsubsection^{L}
start an entry in a list environment. . . . . . . . .  \item^{L}
zero.  start an \halign with \tabskip initialized to  \ialign
number of characters before hyphenation at the  start of a word.  minimum . . \lefthyphenposition^{3}
statement.  character that  starts a new output line in a write . . . . . \newlinechar^{*}
extra vertical space when environment  starts a paragraph. . . . . . . . . . . . . . . . . . . . . . . . \partopsep^{L}
character that starts a new output line in a write  statement. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newlinechar^{*}
show  statistics about memory usage. . . . . . . . \tracingstats^{*}
select font one  step larger than normal size. . . . . . . . . . . . . . . .  \large^{L}
select font four  steps larger than normal size. . . . . . . . . . . . . . .  \huge^{L}
select font three  steps larger than normal size. . . . . . . . . . . . . . .  \LARGE^{L}
select font two  steps larger than normal size. . . . . . . . . . . . . . .  \Large^{L}
permit an alignment entry to  stick out of its column. . . . . . . . . . . . . . . . . . . . \hidewidth
define a tab  stop in tabbing environment. . . . . . . . . . . . . . . .  \=^{L}
unindents left margin one tab  stop in tabbing environment. . . . . . . . . . . . . . . .  \-^{L}
indents left margin one tab  stop in tabbing environment. . . . . . . . . . . . . . . .  \+^{L}
the current line.  stop reading current input file at the end of \endinput^{*}
process input without  stopping for errors. . . . . . . . . . . . . . . . . . . . . \nonstopmode^{*}
restore tabs  stops in tabbing environment. . . . . . . . . . . . . . . \poptabs^{L}
save current tab  stops in tabbing environment. . . . . . . . . . . . . . \pushtabs^{L}
store an hbox or vbox in a box register. . . . . \setbox^{*}
amount of glue  stretch in current page. . . . . . . . . . . . . . . . . \pagestretch^{*}
an unoriented skip amount with  stretch of 1fill. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \fill^{L}
infinitely  stretchable and shrinkable horizontal space.  \hss^{*}
infinitely  stretchable and shrinkable vertical space. …  \vss^{*}
infinitely  stretchable horizontal space. . . . . . . . . . . . . . . .  \hfil^{*}
more infinitely  stretchable horizontal space. . . . . . . . . . . . . . . .  \hfill^{*}
cancel infinitely  stretchable horizontal space. . . . . . . . . . . . . . . . \hfilneg^{*}
infinitely  stretchable space. . . . . . . . . . . . . . . . . . . . . . . . . . . \stretch^{L}
infinitely  stretchable vertical space. . . . . . . . . . . . . . . . . . .  \vfil^{*}

more infinitely  stretchable vertical space. . . . . . . . . . . . . . . . . . .  \vfill^{*}
cancel infinitely  stretchable vertical space. . . . . . . . . . . . . . . . . . . \vfilneg^{*}
require strict line and page breaks. . . . . . . . . . . . . . . . . . \fussy^{L}
convert a number to a token  string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \number^{*}
box containing a  strut. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \strutbox
vertical strut to preserve line spacing. . . . . . . . . . . . . . . \strut
math  strut with height and depth of parentheses. \mathstrut
set sectional units to appendix  style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \appendix^{L}
bold font style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bf^{L}
use normal math script  style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \displaystyle^{*}
set a formula flush left in display  style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \lefteqn^{L}
choose a math formula based on the current  style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathchoice^{*}
specify page numbering  style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pagenumbering^{L}
set global page  style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \pagestyle^{L}
use very small math script  style. . . . . . . . . . . . . . . . . . . . . . . . . . . . \scriptscriptstyle^{*}
use small math script  style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \scriptstyle^{*}
set current page  style. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \thispagestyle^{L}
declare document  style and options. . . . . . . . . . . . . . . . . . . . \documentstyle^{L}
specify bibliographic  style for BibTEX. . . . . . . . . . . . . . . . \bibliographystyle^{L}
use in-text  style for math. . . . . . . . . . . . . . . . . . . . . . . . . . . . \textstyle^{*}
suppress space in script and small script  styles. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \nonscript^{*}
mark the beginning of a major  subdivision. . . . . . . . . . . . . . . . . . . . . . . . . . . \beginsection
define an inner math  subformula. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathinner^{*}
display a  sub-item. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \itemitem
start a labeled  sub-level paragraph. . . . . . . . . . . . . . . . . . . \subparagraph^{L}
subscript, equivalent to _. . . . . . . . . . . . . . . . . .  \sb
extra space after  subscript or superscript. . . . . . . . . . . . . . . . \scriptspace^{*}
select  subscript or superscript font size. . . . . . . . \scriptsize^{L}
place superscripts and  subscripts after math operators. . . . . . . . . . . . \nolimits^{*}
start a  subsection. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \subsection^{L}
\parindent.  indent second and  subsequent lines in a paragraph by . . . . . . . . .  \hang
environment.  indent second and  subsequent paragraphs in a list . . . . . . \listparindent^{L}
math relation: square  subset or equal (⊑). . . . . . . . . . . . . . . . . . . . . . \sqsubseteq
math relation:  subset or equal (⊆). . . . . . . . . . . . . . . . . . . . . . . . \subseteq
math symbol: square  subset (⊏). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \sqsubset^{L}
math relation: subset (⊂). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \subset
select  sub-subscript font size. . . . . . . . . . . . \scriptscriptsize^{L}
start a  subsubsection. . . . . . . . . . . . . . . . . . . . . . . . \subsubsection^{L}
test always  succeeds. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \iftrue^{*}
space between  successive items in a list environment. . . . . . . \itemsep^{L}
math relation:  successor or equal (⪰). . . . . . . . . . . . . . . . . . . . . . \succeq
math relation: successor (≻). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \succ
math symbol: club  suit (♣). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \clubsuit
math symbol: diamond  suit (♢). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \diamondsuit
math symbol: heart  suit (♡). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \heartsuit
math symbol: spade  suit (♠). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \spadesuit
math operator: direct  sum, circle plus (⊕). . . . . . . . . . . . . . . . . . . . . . . .  \oplus
math operator: amalgamated  sum, co-product (⨿). . . . . . . . . . . . . . . . . . . . . . . .  \amalg
on the page.  sum of penalties for split insertions  \insertpenalties^{*}
math operator: large  sum (∑). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sum
math function:  sup. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sup
shift  super accents. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \skew
extra space after subscript or  superscript. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \scriptspace^{*}

superscript, equivalent to ^. . . . . . . . . . . . . . . . .  \sp
select subscript or  superscript font size. . . . . . . . . . . . . . . . . . . . . \scriptsize^{L}
operators. place superscripts and subscripts after math . . . . . \nolimits^{*}
math relation: square  superset or equal (⊒). . . . . . . . . . . . . . . . . . . . \sqsupseteq
math relation:  superset or equal (⊇). . . . . . . . . . . . . . . . . . . . . . \supseteq
math symbol: square  superset (⊐). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \sqsupset^{L}
math relation:  superset (⊃). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \supset
suppress a line break. . . . . . . . . . . . . . . . . . . \nolinebreak^{L}
suppress a page break. . . . . . . . . . . . . . . . . . \nopagebreak^{L}
text.  suppress bottom justification of page . \raggedbottom
suppress expansion of a character. . . . . . . . . . \noexpand^{*}
vertical box.  suppress interline space before next \nointerlineskip
suppress numbering of displayed equations.  \nonumber^{L}
suppress page numbers. . . . . . . . . . . . . . \nopagenumbers
lines. suppress right justification of paragraph \raggedright
font lines.  suppress right justification of typewriter \ttraggedright
styles.  suppress space in script and small script . . \nonscript^{*}
punctuation.  suppress special spacing after . . . . . . . . \frenchspacing
preamble for this entry. suppress the template in the alignment . . . . . \omit^{*}
suppress writing all auxiliary files. . . . . . . . . . . \nofiles^{L}
math function: limit  supremum (limsup). . . . . . . . . . . . . . . . . . . . . . . . \limsup
math symbol: surd (√). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \surd
surround a space with a repeated box or rule. \cleaders^{*}
thickness of rule  surrounding framed box. . . . . . . . . . . . . . . . . . . \fboxrule^{L}
mode.  switch to horizontal mode from vertical . \leavevmode
underscore symbol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \_
display counter as footnote  symbol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fnsymbol^{L}
close brace  symbol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \}
open brace  symbol. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \{
ligature digraph  symbol ae (æ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ae
math  symbol: aleph (ℵ). . . . . . . . . . . . . . . . . . . . . . . . . .  \aleph
angle  symbol (^{}). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \angle
math  symbol: bottom (⊥). . . . . . . . . . . . . . . . . . . . . . . .  \bot
math  symbol: box (□). . . . . . . . . . . . . . . . . . . . . . . . . . .  \Box^{L}
ligature digraph  symbol capital AE (Æ). . . . . . . . . . . . . . . . . . . . .  \AE
ligature digraph  symbol capital OE (Œ). . . . . . . . . . . . . . . . . . . . .  \OE
math  symbol: club suit (♣). . . . . . . . . . . . . . . . . . . . . . \clubsuit
copyright  symbol (\begin{matrix}c \\ ◯ \end{matrix}). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \copyright
dagger  symbol (†). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \dag
double dagger  symbol (‡). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ddag
math  symbol: diamond (◇). . . . . . . . . . . . . . . . . . . . . . \Diamond^{L}
math  symbol: diamond suit (♢). . . . . . . . . . . . . . \diamondsuit
math  symbol: dotless i (ı). . . . . . . . . . . . . . . . . . . . . . . .  \imath
math  symbol: dotless j (ȷ). . . . . . . . . . . . . . . . . . . . . . . .  \jmath
math  symbol: double relation bar (=). . . . . . . . . . . . \Relbar
math  symbol: double vertical bar (‖). . . . . . . . . . \Arrowvert
math  symbol: down arrow (↓). . . . . . . . . . . . . . . . . . \downarrow
math  symbol: downward double arrow (⇓). . . . . . \Downarrow
math  symbol: empty set (∅). . . . . . . . . . . . . . . . . . . . . \emptyset
math  symbol: exists quantifier (∃). . . . . . . . . . . . . . . . \exists
math  symbol: flat (♭). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \flat
8 point L^{A}TEX  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \egtly^{L}

8 point math  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \egtsy^{L}
11 point L^{A}TEX  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvly^{L}
11 point math  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvsy^{L}
5 point math  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \fivesy
5 point L^{A}TEX  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \fivly^{L}
5 point math  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \fivsy^{L}
14 point L^{A}TEX  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \frtnly^{L}
14 point math  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \frtnsy^{L}
9 point L^{A}TEX  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ninly^{L}
9 point math  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ninsy^{L}
7 point math  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \sevensy
7 point L^{A}TEX  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sevly^{L}
7 point math  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sevsy^{L}
6 point L^{A}TEX  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sixly^{L}
6 point math  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \sixsy^{L}
17 point L^{A}TEX  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \svtnly^{L}
17 point math  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \svtnsy^{L}
10 point math extension  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tenex
10 point L^{A}TEX  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tenly^{L}
10 point math  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tensy
12 point L^{A}TEX  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twlly^{L}
12 point math  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twlsy^{L}
20 point L^{A}TEX  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \twtyly^{L}
20 point math  symbol font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \twtysy^{L}
select bold math italic and  symbol fonts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \boldmath^{L}
unselect bold math italic and  symbol fonts. . . . . . . . . . . . . . . . . . . . . . . . . . . . \unboldmath^{L}
math symbol: for-all quantifier (∀). . . . . . . . . . . . . . . \forall
display a  symbol from a font. . . . . . . . . . . . . . . . . . . . . . . . . \symbol^{L}
math  symbol: heart suit (♡). . . . . . . . . . . . . . . . . . . . \heartsuit
math  symbol: hook left arrow (↩). . . . . . . . \hookleftarrow
math  symbol: hook right arrow (↪). . . . . . \hookrightarrow
math  symbol: imaginary, Fraktur I (ℑ). . . . . . . . . . .  \Im
math  symbol: infinity (∞). . . . . . . . . . . . . . . . . . . . . . .  \infty
math  symbol: iota (ι). . . . . . . . . . . . . . . . . . . . . . . . . . . .  \iota
math  symbol: join (⋈). . . . . . . . . . . . . . . . . . . . . . . . . . .  \Join^{L}
math  symbol: leads to (⇝). . . . . . . . . . . . . . . . . . . . . . . \leadsto^{L}
math  symbol: left arrow (←). . . . . . . . . . . . . . . . . . . \leftarrow
math  symbol: left double arrow (⇐). . . . . . . . . . . . \Leftarrow
math symbol: left half diamond (⊲). . . . . . . . . . . . . . \lhd^{L}
math  symbol: left harpoon down (↽)… \leftharpoondown
math  symbol: left harpoon up (↼). . . . . . . . \leftharpoonup
math symbol: left hook ( ). . . . . . . . . . . . . . . . . . . . . . . \lhook
math  symbol: left moustache ({). . . . . . . . . . . . . \lmoustache
math symbol: left quote (‘). . . . . . . . . . . . . . . . . . . . . . . \lq
math  symbol: left-right arrow (↔). . . . . . . . \leftrightarrow
math  symbol: left-right double arrow (⇔). \Leftrightarrow
math  symbol: logical or (∨). . . . . . . . . . . . . . . . . . . . . .  \lor
(⟷).  math  symbol: long left and right arrow \longleftrightarrow
math  symbol: long left arrow (⟵). . . . . . . . \longleftarrow
math  symbol: long left double arrow (⟸).  \Longleftarrow
arrow (⟺).  math  symbol: long left-right double . \Longleftrightarrow
math  symbol: long maps to (⟼). . . . . . . . . . . . . \longmapsto

math  symbol: long right arrow (⟶). . . . . \longrightarrow
math  symbol: long right double arrow (⟹). \Longrightarrow
math  symbol: maps to char ( ). . . . . . . . . . . . . . . . \mapstochar
math  symbol: maps to (↦). . . . . . . . . . . . . . . . . . . . . . . \mapsto
math  symbol: mho (℧). . . . . . . . . . . . . . . . . . . . . . . . . . .  \mho^{L}
math  symbol: nabla (∇). . . . . . . . . . . . . . . . . . . . . . . . .  \nabla
math  symbol: natural (♮). . . . . . . . . . . . . . . . . . . . . . . . . \natural
math  symbol: northeast arrow (↗). . . . . . . . . . . . . . . \nearrow
math  symbol: northwest arrow (↖). . . . . . . . . . . . . . \nwarrow
ligature digraph  symbol oe (œ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \oe
math  symbol: paragraph (¶). . . . . . . . . . . . . . . . . . . . .  \P
math  symbol: partial (∂). . . . . . . . . . . . . . . . . . . . . . . . . \partial
math  symbol: Planck’s constant or h-bar (h̄). . . . .  \hbar
British pound  symbol (£). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \pounds^{L}
math  symbol: prime (x′). . . . . . . . . . . . . . . . . . . . . . . . .  \prime
math  symbol: real, Fraktur R (ℜ). . . . . . . . . . . . . . . .  \Re
math  symbol: relation bar (−). . . . . . . . . . . . . . . . . . .  \relbar
math  symbol: right arrow (→). . . . . . . . . . . . . . . . \rightarrow
math  symbol: right double arrow (⇒). . . . . . . . . \Rightarrow
math  symbol: right half diamond (⊳). . . . . . . . . . . .  \rhd^{L}
math  symbol: right harpoon down (⇁).  \rightharpoondown
math  symbol: right harpoon up (⇀). . . . . . \rightharpoonup
math  symbol: right hook ( ). . . . . . . . . . . . . . . . . . . . . .  \rhook
math  symbol: right left harpoon (⇀ ↽).  \rightleftharpoons
math  symbol: right moustache (}). . . . . . . . . . . . \rmoustache
math  symbol: section (§). . . . . . . . . . . . . . . . . . . . . . . . .  \S
math  symbol: sharp (♯). . . . . . . . . . . . . . . . . . . . . . . . . .  \sharp
math  symbol: small script L (ℓ). . . . . . . . . . . . . . . . . .  \ell
math  symbol: southeast arrow (↘). . . . . . . . . . . . . . . \searrow
math  symbol: southwest arrow (↙). . . . . . . . . . . . . . \swarrow
math  symbol: spade suit (♠). . . . . . . . . . . . . . . . . . . \spadesuit
math  symbol: square subset (⊏). . . . . . . . . . . . . . . . . \sqsubset^{L}
math  symbol: square superset (⊐). . . . . . . . . . . . . . . \sqsupset^{L}
math  symbol: surd (√). . . . . . . . . . . . . . . . . . . . . . . . . . .  \surd
and script sizes.  define a  symbol that will work properly in text . \mathpalette
math  symbol: top (⊤). . . . . . . . . . . . . . . . . . . . . . . . . . . .  \top
math  symbol: triangle (△). . . . . . . . . . . . . . . . . . . . . . \triangle
math  symbol: underlined left half diamond (⊴)…  \unlhd^{L}
math  symbol: underlined right half diamond (⊵).  \unrhd^{L}
math  symbol: up-and-down double arrow (⇕). \Updownarrow
math  symbol: upward double arrow (⇑). . . . . . . . . . \Uparrow
math  symbol: vertical bar (|). . . . . . . . . . . . . . . . . . \arrowvert
math  symbol: vertical bar (∥). . . . . . . . . . . . . . . . . . . .  \|
math  symbol: vertical dots (⋮). . . . . . . . . . . . . . . . . . . .  \vdots
math  symbol: Weierstrass p (℘). . . . . . . . . . . . . . . . . .  \wp
lower dot math  symbol with special spacing. . . . . . . . . . . . . . . .  \ldotp
internal Plain TEX operation to define math text  symbols. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathhexbox
put  symbols over math relations. . . . . . . . . . . . . . . \buildrel
sequence. define a synonym for the current meaning of a control \let^{*}
space between aligned  tab entries. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tabskip^{*}
move to next  tab position in tabbing environment. . . . . . . .  \>^{L}
define a  tab stop in tabbing environment. . . . . . . . . . . .  \=^{L}

unindents left margin one  tab stop in tabbing environment. . . . . . . . . . . .  \-^{L}
indents left margin one  tab stop in tabbing environment. . . . . . . . . . . .  \+^{L}
save current  tab stops in tabbing environment. . . . . . . . . . \pushtabs^{L}
begin a  tabbed line in an inner environment. . . . . . . \tabalign
begin a  tabbed line in an outer environment. . . . . . . .  \+
produce TEX accents in  tabbing environment. . . . . . . . . . . . . . . . . . . . . . .  \a^{L}
discard current line in  tabbing environment. . . . . . . . . . . . . . . . . . . . . . .  \kill^{L}
cancel effect of one \+ command in  tabbing environment. . . . . . . . . . . . . . . . . . . . . . .  \<^{L}
define a tab stop in  tabbing environment. . . . . . . . . . . . . . . . . . . . . . .  \=^{L}
move to next tab position in  tabbing environment. . . . . . . . . . . . . . . . . . . . . . .  \>^{L}
put text flush right in a column in  tabbing environment. . . . . . . . . . . . . . . . . . . . . . .  \'^{L}
unindents left margin one tab stop in  tabbing environment. . . . . . . . . . . . . . . . . . . . . . .  \-^{L}
restore tabs stops in  tabbing environment. . . . . . . . . . . . . . . . . . . . . . . \poptabs^{L}
save current tab stops in  tabbing environment. . . . . . . . . . . . . . . . . . . . . . . \pushtabs^{L}
indents left margin one tab stop in tabbing environment. . . . . . . . . . . . . . . . . . . . . . . \+^{L}
distance left by \' command between  tabbing fields. . . . . . . . . . . . . . . . . . . . . . . . . . . . \tabbingsep^{L}
adds an entry to the specified list or  table. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \addcontentsline^{L}
create an aligned  table. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \halign^{*}
create a multicolumn entry in an aligned table. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \span^{*}
add text to  table contents, figures, or tables. . . . . . \addtocontents^{L}
generate a  table of contents. . . . . . . . . . . . . . . . . . \tableofcontents^{L}
add text to table contents, figures, or  tables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \addtocontents^{L}
generate a list of  tables. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \listoftables^{L}
page. flush figures and tables and start a new right-hand . \cleardoublepage^{L}
reset  tabs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \cleartabs
define horizontal  tabs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \settabs
box containing current tabs. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tabs
restore  tabs stops in tabbing environment. . . . . . . . . . \poptabs^{L}
start an \halign with \tabskip initialized to zero. . . . . . . . . . . . . . . . \ialign
space between rows of array or  tabular environment. . . . . . . . . . . . . . . . . . \arraystretch^{L}
multicolumn line in array or tabular environment. . . . . . . . . . . . . . . . . . . . . . . . \cline^{L}
double rule separation in array or  tabular environment. . . . . . . . . . . . . . . . . \doublerulesep^{L}
add extra space before a column in array or  tabular environment. . . . . . . . . . . . . . . . . . . \extracolsep^{L}
multicolumn entry in array or  tabular environment. . . . . . . . . . . . . . . . . . . \multicolumn^{L}
half the width separating columns in a  tabular environment. . . . . . . . . . . . . . . . . . . . . . \tabcolsep^{L}
horizontal line in array and  tabular environments. . . . . . . . . . . . . . . . . . . . . . .  \hline^{L}
vertical line in array and  tabular environments. . . . . . . . . . . . . . . . . . . . . . .  \vline^{L}
math function: arc  tangent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \arctan
math function:  tangent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tan
math function: hyperbolic  tangent. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tanh
math Greek letter:  tau (τ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tau
entry.  suppress the  template in the alignment preamble for this  \omit^{*}
math operator:  tensor product, circle times (⊗). . . . . . . . . . . .  \otimes
write balanced error message to the  terminal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \errmessage^{*}
write balanced text to  terminal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \message^{*}
show diagnostics on the  terminal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tracingonline^{*}
execute a command from the  terminal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \typein^{L}
write a message on the  terminal. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \typeout^{L}
terminate a \csname token list. . . . . . . . . . . . \endcsname^{*}
terminate L^{A}TEX and flush the final page. …  \stop^{L}
terminate TEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \end^{*}
INITEX only. terminate TEX and write a format file: . . . . . \dump^{*}
en entry in a file of index  terms. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \indexentry^{L}

test always false. . . . . . . . . . . . . . . . . . . . . . . . . . . . \iffalse^{*}
test always succeeds. . . . . . . . . . . . . . . . . . . . . . . . \iftrue^{*}
test category codes. . . . . . . . . . . . . . . . . . . . . . . . .  \ifcat^{*}
test for a horizontal box. . . . . . . . . . . . . . . . . . . . \ifhbox^{*}
test for a vertical box. . . . . . . . . . . . . . . . . . . . . .  \ifvbox^{*}
test for an empty box. . . . . . . . . . . . . . . . . . . . . .  \ifvoid^{*}
test for an internal mode. . . . . . . . . . . . . . . . . . . \ifinner^{*}
test for end of file. . . . . . . . . . . . . . . . . . . . . . . . . .  \ifeof^{*}
test for horizontal mode. . . . . . . . . . . . . . . . . . . . \ifhmode^{*}
test for math mode. . . . . . . . . . . . . . . . . . . . . . . . . \ifmmode^{*}
test for odd integer. . . . . . . . . . . . . . . . . . . . . . . . .  \ifodd^{*}
test for vertical mode. . . . . . . . . . . . . . . . . . . . . . . \ifvmode^{*}
test the next character. . . . . . . . . . . . . . . . . \@ifnextchar^{L}
test two dimensions. . . . . . . . . . . . . . . . . . . . . . . .  \ifdim^{*}
test two integers. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ifnum^{*}
terminate  TEX. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \end^{*}
produce  TEX accents in tabbing environment. . . . . . . .  \a^{L}
terminate  TEX and write a format file: INITEX only. .  \dump^{*}
name of current  TEX format package. . . . . . . . . . . . . . . . . . . . . . . . \fmtname
version of current  TEX format package. . . . . . . . . . . . . . . . . . . . . \fmtversion
pause for normal errors while processing  TEX input. . . . . . . . . . . . . . . . . . . . . . . . . . . \errorstopmode^{*}
errors. process TEX input without pausing for normal . . \scrollmode^{*}
TEX primitive \par. . . . . . . . . . . . . . . . . . . . . . . .  \@@par^{L}
space between columns in double column  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \columnsep^{L}
end a line in \halign aligned text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \cr^{*}
rule between double-column floats and  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \dblfigrule^{L}
space between double-column floats and  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \dbltextfloatsep^{L}
define a macro with expanded replacement text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \edef^{*}
equivalent to \cr, end of aligned  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \endline
insert a footnote mark without  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footnotemark^{L}
macro to draw the rule separating footnotes from  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \footnoterule^{L}
put a frame around some text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \frame^{L}
space between page header and  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \headsep^{L}
distance between marginal note and  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \marginparsep^{L}
define kerning around math in  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathsurround^{*}
verbatim line breaks in input  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \obeylines
verbatim spaces in input  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \obeyspaces
suppress bottom justification of page  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \raggedbottom
set a length command to width of  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \settowidth^{L}
space between floats and the  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \textfloatsep^{L}
select font for non-math  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \textfont^{*}
size of column that must contain  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \textfraction^{L}
place rule between top floats and  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \topfigrule^{L}
generate a short amount of verbatim  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \verb^{L}
define a global macro with expanded replacement  text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \xdef^{*}
place rule between  text and bottom floats. . . . . . . . . . . . . . . . . . \botfigrule^{L}
space between main  text and footnotes. . . . . . . . . . . . . . . . . . . . . . . . . . \footins^{L}
space between  text and page footer. . . . . . . . . . . . . . . . . . . . . . . \footskip^{L}
define a symbol that will work properly in  text and script sizes. . . . . . . . . . . . . . . . . . . . \mathpalette
insert  text at the top of the page. . . . . . . . . . . . . . . . \topinsert
tokens to insert when math in  text begins. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \everymath^{*}
height of  text, excluding head and foot. . . . . . . . . . . . \textheight^{L}
mark  text first encountered in a split box.  \splitfirstmark^{*}

mark  text first encountered on page just boxed.  \firstmark^{*}
environment.  put  text flush right in a column in tabbing . . . . .  \'^{L}
center  text in a line. . . . . . . . . . . . . . . . . . . . . . . . . . . . \centerline
create an underlined box with  text in it. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \underline^{L}
include balanced  text in DVI file for post-processing. . . . . . . . . \special^{*}
place balanced  text into a mark item on the current list. …  \mark^{*}
10 point  text italic font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tenit
mark  text last encountered in a split box. … \splitbotmark^{*}
mark  text last encountered on page just boxed… \botmark^{*}
flush  text left on a line. . . . . . . . . . . . . . . . . . . . . . . . . . \leftline
space between a label and  text of a list item. . . . . . . . . . . . . . . . . . . . . . . . . . \labelsep^{L}
flush right  text on a line. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \rightline
height of text on a page or \vbox. . . . . . . . . . . . . . . . . . . . \vsize^{*}
length of  text preceding a display. . . . . . . . . . . . \predisplaysize^{*}
internal Plain TEX operation to define math  text symbols. . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathhexbox
translate balanced  text to lower-case. . . . . . . . . . . . . . . . . . . . . . . . . \lowercase^{*}
add  text to table contents, figures, or tables. \addtocontents^{L}
write balanced  text to terminal. . . . . . . . . . . . . . . . . . . . . . . . . . . . \message^{*}
create a zero-width box with  text to the left. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \llap
create a zero-width box with  text to the right. . . . . . . . . . . . . . . . . . . . . . . . . . . .  \rlap
translate balanced  text to upper-case. . . . . . . . . . . . . . . . . . . . . . . . \uppercase^{*}
formatted. define where text will be inserted when the page is . . . . . . \insert^{*}
produce footnote  text without a mark. . . . . . . . . . . . . . . . . . \footnotetext^{L}
select font four steps larger  than normal size. . . . . . . . . . . . . . . . . . . . . . . . . . .  \huge^{L}
select font one step larger  than normal size. . . . . . . . . . . . . . . . . . . . . . . . . . .  \large^{L}
select font three steps larger  than normal size. . . . . . . . . . . . . . . . . . . . . . . . . . .  \LARGE^{L}
select font two steps larger  than normal size. . . . . . . . . . . . . . . . . . . . . . . . . . .  \Large^{L}
define a new  theorem environment. . . . . . . . . . . . . . . . . . . . \newtheorem^{L}
environment.  begin  theorem with special format in math . . . . . . \proclaim
math Greek letter:  theta (θ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \theta
math Greek letter: capital  theta (Θ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Theta
math Greek letter: variant  theta (ϑ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vartheta
thick lines for lines and circles. . . . . . . . . . . \thicklines^{L}
thick space in math formulas. . . . . . . . . . . \thickmuskip^{*}
thick space (math mode). . . . . . . . . . . . . . . . . . .  \;
fraction with rule thickness. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \above^{*}
thickness of rule surrounding framed box… \fboxrule^{L}
thin lines for lines and circles. . . . . . . . . . . . . \thinlines^{L}
math mode negative thin space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \!
thin space in math formulas. . . . . . . . . . . . . \thinmuskip^{*}
thin space (math mode). . . . . . . . . . . . . . . . . . . .  \,
current value of the  third-level item counter. . . . . . . . . . . . . . . . . \theenumiii^{L}
third-level item in an index. . . . . . . . . . . . . . \subsubitem^{L}
mark denoting  third-level items. . . . . . . . . . . . . . . . . . . . . . \labelitemiii^{L}
width of left margin in  third-level list. . . . . . . . . . . . . . . . . . . . . . . \leftmarginiii^{L}
tie-after accent (oo). . . . . . . . . . . . . . . . . . . . . . . .  \t
tilde accent (ñ). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \~
math accent: tilde (x̃). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tilde
math accent: wide  tilde (x̃). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \widetilde
space. unbreakable tiny negative math mode horizontal … \negthinspace
produce the  title. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \maketitle^{L}
declare the  title. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \title^{L}
additional authors on  title page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \and^{L}

date on  title page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \date^{L}
add footnote to  title page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \thanks^{L}
title page author. . . . . . . . . . . . . . . . . . . . . . . . . . .  \author^{L}
display today’s date. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \today^{L}
internal Plain TEX command to piece  together long arrows. . . . . . . . . . . . . . . . . . . . . . . . \joinrel
expand the token following the next  token. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \expandafter^{*}
give a token list defining the semantics of a  token. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \meaning^{*}
show a token. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \show^{*}
insert a  token after the current group is completed. \aftergroup^{*}
command. insert a token after the next assignment … \afterassignment^{*}
expand the  token following the next token. . . . . . . . . \expandafter^{*}
terminate a \csname  token list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \endcsname^{*}
give a  token list defining the semantics of a token.  \meaning^{*}
use a  token list register. . . . . . . . . . . . . . . . . . . . . . . . . . .  \toks^{*}
define a name for a  token list register. . . . . . . . . . . . . . . . . . . . . . . . . . . \toksdef^{*}
write a token list to a file. . . . . . . . . . . . . . . . . . . . . . . . . . \write^{*}
allocate a new  token register. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newtoks
convert a number to a  token string. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \number^{*}
assigns the second  token to a control sequence and continues.  \futurelet^{*}
an empty list of  tokens. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \empty
character in the output of control sequence  tokens.  escape . . . . . . . . . . . . . . . . . . . . . . . \escapechar^{*}
compare  tokens. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ifx^{*}
expand a control sequence into character  tokens. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \string^{*}
begin expanding  tokens to construct a command name. . . . . . . \csname^{*}
nonredundant \crcr.  tokens to insert after every \cr or . . . . . . . . . . \everycr^{*}
tokens to insert when a paragraph begins. . \everypar^{*}
tokens to insert when a vbox begins. . . . . . \everyvbox^{*}
tokens to insert when an hbox begins. . . . . \everyhbox^{*}
begins.  tokens to insert when display math … \everydisplay^{*}
tokens to insert when math in text begins. \everymath^{*}
tokens to insert when the job begins. . . . . . . \everyjob^{*}
read, expand, then ignore  tokens until a non-space is found. . . . . . \ignorespaces^{*}
badness  tolerance after hyphenation. . . . . . . . . . . . . . . \tolerance^{*}
attempted.  badness  tolerance before hyphenation is . . . . . . . \pretolerance^{*}
vertical mode material with the baseline at the  top.  create a box with . . . . . . . . . . . . . . . . . .  \vtop^{*}
fraction of two-column page for  top floats. . . . . . . . . . . . . . . . . . . . . . . . . . . \dbltopfraction^{L}
fraction of column for  top floats. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \topfraction^{L}
place rule between  top floats and text. . . . . . . . . . . . . . . . . . . . . . \topfigrule^{L}
space at the  top of a page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \topskip^{*}
insertion class for inserts at the  top of a page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \topins
space at  top of a split box. . . . . . . . . . . . . . . . . . . . . \splittopskip^{*}
extra space added to  top of page. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \topmargin^{L}
penalty for creating a widow line at  top of page. . . . . . . . . . . . . . . . . . . . . . . . . . . \widowpenalty^{*}
insert text at the  top of the page. . . . . . . . . . . . . . . . . . . . . . . . . . . \topinsert
math symbol: top (⊤). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \top
trace output processing. . . . . . . . . . . . . . . . . . \showoutput^{L}
translate balanced text to lower-case. . . . . . \lowercase^{*}
translate balanced text to upper-case. . . . . \uppercase^{*}
math operator: large down  triangle (▽). . . . . . . . . . . . . . . . . . . . . . . \bigtriangledown
math operator: large up  triangle (△). . . . . . . . . . . . . . . . . . . . . . . . . \bigtriangleup
math symbol:  triangle (△). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \triangle
math operator: left  triangle (◁). . . . . . . . . . . . . . . . . . . . . . . . . . . \triangleleft
math operator: right  triangle (▷). . . . . . . . . . . . . . . . . . . . . . . . . . \triangleright

fraction of  two-column page for top floats. . . . . . \dbltopfraction^{L}
set a character’s processing category  type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \catcode^{*}
set a math character’s processing category  type. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \mathcode^{*}
typeset page number. . . . . . . . . . . . . . . . . . . . . . .  \folio
11 point  typewriter font. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \elvtt^{L}
9 point  typewriter font. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \nintt^{L}
10 point  typewriter font. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tentt
select  typewriter font. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \tt
12 point  typewriter font. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \twltt^{L}
typewriter font family. . . . . . . . . . . . . . . . . . . . . .  \ttfam
suppress right justification of  typewriter font lines. . . . . . . . . . . . . . . . . \ttraggedright
math operator: large  U plus (⨄). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \biguplus
math operator:  U plus (⊎). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \uplus
diæresis or  umlaut accent (ẍ). . . . . . . . . . . . . . . . . . . . . . . . . .  \"
long Hungarian  umlaut accent (ő). . . . . . . . . . . . . . . . . . . . . . . . . .  \H
insert  unaligned material in \halign or \valign. . \noalign^{*}
show  unassignments when groups end… \tracingrestores^{*}
vertical list.  un-box a copy of a vbox and add it to the . . \unvcopy^{*}
horizontal list. un-box a copy of an hbox and add it to the \unhcopy^{*}
un-box a vbox and add it to the vertical list.  \unvbox^{*}
list. un-box an hbox and add it to the horizontal \unhbox^{*}
unbreakable small horizontal space. . . . . . . . . \enspace
unbreakable small horizontal space. . . . . . . . \thinspace
unbreakable space. . . . . . . . . . . . . . . . . . . . . . . . . .  ~
horizontal space.  unbreakable tiny negative math mode . \negthinspace
an  undefined control sequence. . . . . . . . . . . . . . . \undefined
draw a line  under a formula. . . . . . . . . . . . . . . . . . . . . . . . . . \underline^{*}
horizontal brace  under a math formula. . . . . . . . . . . . . . . . . . . \underbrace
put a math field  under a radical. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \radical^{*}
dot  under accent (\begin{matrix}x \\ . \end{matrix}). . . . . . . . . . . . . . . . . . . . . . . . . . . .  \d
math accent: bar  under (x̄). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \b
create a box with an  underline. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \underbar
create an  underlined box with text in it. . . . . . . . . . . . \underline^{L}
math symbol:  underlined left half diamond (⊴). . . . . . . . . . .  \unlhd^{L}
math symbol:  underlined right half diamond (⊵). . . . . . . . . .  \unrhd^{L}
underscore symbol. . . . . . . . . . . . . . . . . . . . . . . . . .  \_
environment.  unindents left margin one tab stop in tabbing  \-^{L}
math operator:  union or cup (∪). . . . . . . . . . . . . . . . . . . . . . . . . . .  \cup
unit of distance in picture environment. . \unitlength^{L}
unit of measure for opening up displays. . . . .  \jot
set sectional  units to appendix style. . . . . . . . . . . . . . . . . . . . \appendix^{L}
an unoriented skip amount with stretch of 1fill. \fill^{L}
unselect bold math italic and symbol fonts. \unboldmath^{L}
define an  unusual paragraph shape. . . . . . . . . . . . . . . . . . \parshape^{*}
math delimiter:  up-and-down arrow (↕). . . . . . . . . . . . . . . . . \updownarrow
math symbol:  up-and-down double arrow (⇕). . . . . . . . . \Updownarrow
upper bound on output routine calls.  \maxdeadcycles^{*}
upper left piece of a horizontal brace. . . . . . . \bracelu
upper right piece of a horizontal brace. . . . . . \braceru
translate balanced text to  upper-case. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \uppercase^{*}
upper-case code for a character. . . . . . . . . . . . . \uccode^{*}
display counter as  upper-case letter. . . . . . . . . . . . . . . . . . . . . . . . . . .  \Alph^{L}
display counter as  upper-case Roman numerals. . . . . . . . . . . . . . . .  \Roman^{L}

Polish letter:  upper-case slashed L (Ł). . . . . . . . . . . . . . . . . . .  \L
math Greek letter:  upsilon (υ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \upsilon
math Greek letter: capital  upsilon (Υ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \Upsilon
math delimiter:  upward arrow (↑). . . . . . . . . . . . . . . . . . . . . . . . . . . \uparrow
fill a space with an  upward brace. . . . . . . . . . . . . . . . . . . . . . . . . . \upbracefill
math symbol:  upward double arrow (⇑). . . . . . . . . . . . . . . . . . . \Uparrow
show statistics about memory  usage. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tracingstats^{*}
redefinable scratch control sequence used by \dospecials. . . . . . . . . . . . . . . . . . . . . . . \do
internal Plain TEX space  used for centering. . . . . . . . . . . . . . . . . . . . . . . . \centering
current language  used for hyphenation. . . . . . . . . . . . . . . . . . . . . . \language^{3}
define a new language to be  used for hyphenation. . . . . . . . . . . . . . . . . . . \newlanguage^{3}
specify language to be  used for hyphenation. . . . . . . . . . . . . . . . . . . \setlanguage^{3}
special Plain TEX space  used in alignment. . . . . . . . . . . . . . . . . . . . . . . . . . \hideskip
scratch control sequence  used in preloading fonts. . . . . . . . . . . . . . . . . . \preloaded
help message to display if  user asks for help. . . . . . . . . . . . . . . . . . . . . . . . . . . \errhelp^{*}
restores a carriage return to its  usual meaning. . . . . . . . . . . . . . . . . . . . . . . . . . . . \restorecr^{L}
math join operator: large  V (⋁). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigvee
math relation: dash  V (⊣). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \dashv
math binary operator: join or  V (∨). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vee
field separator in \halign or  \valign. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  &^{*}
insert unaligned material in \halign or  \valign. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \noalign^{*}
divide a register by a  value. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \divide^{*}
multiply a register by a  value. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \multiply^{*}
set  value of a counter. . . . . . . . . . . . . . . . . . . . . . . \setcounter^{L}
produce the value of a counter. . . . . . . . . . . . . . . . . . . . . . . . . . \value^{L}
increment the  value of a length command. . . . . . . . . . . . . \addtolength^{L}
was boxed.  value of \botmark just before current page . \topmark^{*}

                    value of normal \lineskip. . . . . . . . . \normallineskip
\lineskiplimit.  value of normal . . . . . . . . . . . . . . \normallineskiplimit

current  value of the fifth-level item counter. . . . . . . . \theenumv^{L}
current  value of the figure counter. . . . . . . . . . . . . . . . \thefigure^{L}
current  value of the first-level item counter. . . . . . . . \theenumi^{L}
current  value of the fourth-level item counter. . . . . \theenumiv^{L}
current  value of the second-level item counter. . . . . \theenumii^{L}
current  value of the sixth-level item counter. . . . . . \theenumvi^{L}
current  value of the third-level item counter. . . . . \theenumiii^{L}
set line spacing to normal  values. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \normalbaselines
define a new counter  variable. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \newcounter^{L}
ratio for  variable delimiters times 1000. . . . . \delimiterfactor^{*}
math Greek letter:  variant epsilon (ε). . . . . . . . . . . . . . . . . . . . . . . \varepsilon
math Greek letter:  variant phi (φ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . \varphi
math Greek letter: variant pi (ϖ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \varpi
math Greek letter:  variant rho (ϱ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . \varrho
math Greek letter:  variant sigma (ς). . . . . . . . . . . . . . . . . . . . . . . . . . \varsigma
math Greek letter:  variant theta (ϑ). . . . . . . . . . . . . . . . . . . . . . . . . . . \vartheta
height of text on a page or \vbox. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vsize^{*}
split off a specified amount from a  vbox. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vsplit^{*}
un-box a vbox and add it to the vertical list. . . . . . . . . \unvbox^{*}
un-box a copy of a  vbox and add it to the vertical list. . . . . . . . . \unvcopy^{*}
tokens to insert when a  vbox begins. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \everyvbox^{*}
maximum overrun before overfull vbox error. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vfuzz^{*}
limit for bad  vbox errors. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vbadness^{*}
store an hbox or  vbox in a box register. . . . . . . . . . . . . . . . . . . . . . \setbox^{*}

create a centered  vbox in a math list. . . . . . . . . . . . . . . . . . . . . . . . . \vcenter^{*}
create a zero-width  vbox the height and depth of a formula. … \vphantom
math accent: vector (x⃗). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vec
vector in a picture environment. . . . . . . . . . . . . \vector^{L}
verbatim line breaks in input text. . . . . . . . \obeylines
verbatim spaces in input text. . . . . . . . . . . . \obeyspaces
generate a short amount of verbatim text. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \verb^{L}
version of current TEX format package… \fmtversion
math symbol:  vertical bar (|). . . . . . . . . . . . . . . . . . . . . . . . . . . \arrowvert
math symbol: double  vertical bar (‖). . . . . . . . . . . . . . . . . . . . . . . . . . \Arrowvert
math symbol:  vertical bar (∥). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \|
math delimiter:  vertical bar (|). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vert
math delimiter: double  vertical bar (∥). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Vert
test for a  vertical box. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \ifvbox^{*}
suppress interline space before next  vertical box. . . . . . . . . . . . . . . . . . . . . . . \nointerlineskip
piece of a  vertical brace (|). . . . . . . . . . . . . . . . . . . . . . . . \bracevert
large  vertical break. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigskip
math symbol:  vertical dots (⋮). . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vdots
environments.  vertical line in array and tabular . . . . . . . . . . .  \vline^{L}
minimum space between lines in a  vertical list. . . . . . . . . . . . . . . . . . . . . . . . . . \lineskiplimit^{*}
depth of the last box on the  vertical list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \prevdepth^{*}
un-box a vbox and add it to the  vertical list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \unvbox^{*}
un-box a copy of a vbox and add it to the  vertical list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \unvcopy^{*}
place a footnote in a caption or other  vertical list. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vfootnote
produce a  vertical list of the body of a page. . . . . \pagecontents
add a box to the  vertical list shifted left. . . . . . . . . . . . . . . . . . . . . \moveleft^{*}
add a box to the  vertical list shifted right. . . . . . . . . . . . . . . . . . \moveright^{*}
insert  vertical material into a paragraph. . . . . . . . . . \vadjust^{*}
test for  vertical mode. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ifvmode^{*}
switch to horizontal mode from  vertical mode. . . . . . . . . . . . . . . . . . . . . . . . . . . \leavevmode
create a box with  vertical mode material. . . . . . . . . . . . . . . . . . . . . .  \vbox^{*}
the top.  create a box with  vertical mode material with the baseline at .  \vtop^{*}
vertical offset of a page. . . . . . . . . . . . . . . . . . . . . \voffset^{*}
produce a vertical rule. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vrule^{*}
vertical skip a large amount. . . . . . . . . . . . . . . . \bigskip^{L}
vertical skip a medium amount. . . . . . . . . . . . . \medskip^{L}
vertical skip a small amount. . . . . . . . . . . . . . \smallskip^{L}
add extra  vertical space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \addvspace^{L}
medium  vertical space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \medskip
small  vertical space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . \smallskip
infinitely stretchable  vertical space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vfil^{*}
more infinitely stretchable  vertical space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vfill^{*}
cancel infinitely stretchable  vertical space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vfilneg^{*}
create some  vertical space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vglue
add  vertical space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \vskip^{*}
skip  vertical space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vspace^{L}
infinitely stretchable and shrinkable vertical space. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \vss^{*}
small  vertical space and a good page break. … \smallbreak
a page.  vertical space around a float in the middle of \intextsep^{L}
vertical space between marginal notes. \marginparpush^{L}
large  vertical space or a good page break. . . . . . . . \bigbreak

medium  vertical space or a good page break. . . . . . . . \medbreak
paragraph. extra vertical space when environment starts a . \partopsep^{L}
vertical strut to preserve line spacing. . . . . . .  \strut
math binary operator: meet or wedge (∧). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \wedge
math meet operator: large  wedge (⋀). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \bigwedge
math symbol:  Weierstrass p (℘). . . . . . . . . . . . . . . . . . . . . . . . . . .  \wp
formatted.  define  where text will be inserted when the page is  \insert^{*}
10 point  wide circle font. . . . . . . . . . . . . . . . . . . . . . . . . . . . \tencircw^{L}
math accent:  wide hat (x̂). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \widehat
10 point wide line font. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \tenlnw^{L}
math accent:  wide tilde (x̃). . . . . . . . . . . . . . . . . . . . . . . . . . . . . \widetilde
penalty for creating a  widow line at top of page. . . . . . . . . . . . . \widowpenalty^{*}
penalty for creating a  widow line before a display… \displaywidowpenalty^{*}
create a box of current line width. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \line
label  width in a list environment. . . . . . . . . . . . . . \labelwidth^{L}
width of a box. . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \wd^{*}
width of a column. . . . . . . . . . . . . . . . . . . . . . \columnwidth^{L}
width of a displayed equation. . . . . . . . . \displaywidth^{*}
create a zero-height hbox with the  width of a formula. . . . . . . . . . . . . . . . . . . . . . . . . \hphantom
width of a null delimiter. . . . . . . \nulldelimiterspace^{*}
width of a paragraph or \hbox. . . . . . . . . . . . .  \hsize^{*}
set  width of array rules. . . . . . . . . . . . . . . . \arrayrulewidth^{L}
width of left margin in fifth-level list… \leftmarginv^{L}
width of left margin in first-level list… \leftmargini^{L}
width of left margin in fourth-level list. \leftmarginiv^{L}
width of left margin in second-level list. \leftmarginii^{L}
width of left margin in sixth-level list.  \leftmarginvi^{L}
width of left margin in third-level list. \leftmarginiii^{L}
set  width of lines in picture environment. \linethickness^{L}
width of marginal notes. . . . . . . . . . . . \marginparwidth^{L}
width of printing on page. . . . . . . . . . . . . . . . . \textwidth^{L}
columns.  width of rule separating double . . . . . . \columnseprule^{L}
boxes.  width of rules appended to overfull … \overfullrule^{*}
set a length command to  width of text. . . . . . . . . . . . . . . . . . . . . . . . . . . . \settowidth^{L}
width of the current line. . . . . . . . . . . . . . . . . . \linewidth^{L}
environment.  half the  width separating columns in a tabular …. \tabcolsep^{L}
of characters before hyphenation at the start of a  word.  minimum number … \lefthyphenposition^{3}
of characters after hyphenation at the end of a  word.  minimum number . . \righthyphenposition^{3}
word delimiter character. . . . . . . . . . . . . . \boundarychar^{3}
show hyphenations of given  words. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \showhyphens
space between  words. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \spaceskip^{*}
define a symbol that will  work properly in text and script sizes. . \mathpalette
math binary operator:  wreath product (≀). . . . . . . . . . . . . . . . . . . . . . . . .  \wr
terminate TEX and  write a format file: INITEX only. . . . . . . . . . .  \dump^{*}
write a glossary entry. . . . . . . . . . . . . . . . . . . . . . \glossary^{L}
write a message on the terminal. . . . . . . . . . . . \typeout^{L}
write a token list to a file. . . . . . . . . . . . . . . . . . .  \write^{*}
terminal.  write balanced error message to the . . . . . \errmessage^{*}
write balanced text to terminal. . . . . . . . . . . . . \message^{*}
perform a \read or  \write immediately. . . . . . . . . . . . . . . . . . . . . . \immediate^{*}
character that starts a new output line in a  write statement. . . . . . . . . . . . . . . . . . . . . . . . \newlinechar^{*}
write to the log file only. . . . . . . . . . . . . . . . . . . .  \wlog
suppress  writing all auxiliary files. . . . . . . . . . . . . . . . . . . . \nofiles^{L}

math Greek letter:  xi (ξ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \xi
math Greek letter: capital  xi (Ξ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \Xi
current month of the  year. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \month^{*}
start an \halign with \tabskip initialized to  zero. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \ialign
create a formula with zero height. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . \smash
align  zero or more columns. . . . . . . . . . . . . . . . . . . . . . . \valign^{*}
create a  zero-height hbox with the width of a formula. \hphantom
create a  zero-width box with text to the left. . . . . . . . .  \llap
create a  zero-width box with text to the right. . . . . . .  \rlap
formula.  create a  zero-width vbox the height and depth of a . \vphantom
math Greek letter:  zeta (ζ). . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . .  \zeta

Primitive and Plain TEX commands

This is a list of the TEX primitives and the Plain TEX commands. Primitives are marked with an asterisk. TEX version 3 commands are marked with a ‘3’.

*␣ space character. ~ unbreakable space. *$ enter math mode. *$$ enter display math mode. % begin comment. *& field separator in \halign or \valign. *\␣ control space. \! math mode negative thin space. \# pound, hatch mark, sharp sign,

octothorpe (#).

\$ dollar sign. \% percent sign (%). \& ampersand. \' acute accent (é). \* math operator: discretionary multiply

sign.

\+ begin a tabbed line in an outer

environment.

\, thin space (math mode). *\- discretionary hyphen. \. dot accent (ẋ). *\/ italic correction. \; thick space (math mode). \= macron or bar accent (ō). \> medium space in math mode. \{ open brace symbol. \} close brace symbol. \^circumflex accent (ô). \_underscore symbol. \` grave accent (è). \| math symbol: vertical bar (∥). \~ tilde accent (ñ). \" diæresis or umlaut accent (ẍ). \aa symbol å. \AA aymcol Å.

*\above fraction with rule thickness.

                                              *\abovedisplayshortskip extra space above

displays following short lines.

*\abovedisplayskip extra space above displays. *\abovewithdelims fraction with specified rule and delimiters. *\accent put an accent over the next character. \active category code for active characters. \acute math accent: acute (x́). *\adjdemerits demerits for adjacent incompatible

lines. *\advance perform arithmetic on a register.

                                               \advancepageno advance \pageno by one.

\ae ligature digraph symbol ae (æ). \AE ligature digraph symbol capital AE

(Æ).

*\afterassignment insert a token after the next

assignment command.

*\aftergroup insert a token after the current

group is completed.

\aleph math symbol: aleph (ℵ).

\allocationnumber most recently allocated register number.

\allowbreak allow a line break.

\alpha math Greek letter: alpha (α). \amalg math operator: amalgamated sum, co-product (⨿).

\angle angle symbol (^{}).

\approx math relation: approximately equal

(≈). \arccos math function: arc cosine. \arcsin math function: arc sine. \arctan math function: arc tangent. \arg math function: arg. \arrowvert math symbol: vertical bar (|). \Arrowvert math symbol: double vertical bar

(‖). \ast math operator: asterisk (∗).

\asymp math relation: asymptote (≍). *\atop fraction without rule. *\atopwithdelims fraction without rule with

given delimiters.

\b math accent: bar under (x̄). \backslash math delimiter: backslash (\). 3\badness badness of a box. \bar math accent: bar (x̄). *\baselineskip normal space between lines. *\batchmode process input without displaying

errors.

*\begingroup begin a group. \beginsection mark the beginning of a major

subdivision.

*\belowdisplayshortskip extra space just below

displays following short lines.

*\belowdisplayskip extra space just below

displays.

\beta math Greek letter: beta (β). \bf select bold extended font. \bffam bold font family. \bgroup begin a group. \big 1-line math delimiter size. \Big 1.5-line math delimiter size. \bigbreak large vertical space or a good page

break.

\bigcap math operator: large cap (⋂). \bigcirc math operator: large circle (◯). \bigcup math operator: large cup (⋃). \bigg 2-line math delimiter size. \Bigg 2.5-line math delimiter size. \biggl 2-line left math delimiter. \Biggl 2.5-line left math delimiter. \biggm 2-line middle math delimiter. \Biggm 2.5-line middle math delimiter. \biggr 2-line right math delimiter. \Biggr 2.5-line right math delimiter. \bigl 1-line left math delimiter. \Bigl 1.5-line left math delimiter. \bigm 1-line middle math delimiter. \Bigm 1.5-line middle math delimiter. \bigodot math operator: large circle with dot

(⨀).

\bigoplus math operator: large circle with plus

(⨁).

\bigotimes math operator: large circle with

times (⨂).

\bigr 1-line right math delimiter. \Bigr 1.5-line right math delimiter. \bigskip large vertical break.

\bigskipamount  \bigskip space.

\bigsqcup math operator: large square cup (⨆).

\bigtriangledown math operator: large down triangle (▽). \bigtriangleup math operator: large up triangle

(△).

\biguplus math operator: large U plus (⨄). \bigvee math join operator: large V (⋁). \bigwedge math meet operator: large wedge (⋀). *\binoppenalty penalty for line break after

binary operation.

\bmod math operator: binary modulo (mod). \bordermatrix generate a matrix labeled on

rows and columns.

\bot math symbol: bottom (⊥).

*\botmark mark text last encountered on page just boxed. 3\boundarychar word delimiter character. \bowtie math relation: bowtie (▷◁). *\box use a box. *\boxmaxdepth maximum depth of boxes on explicit pages. \brace fraction without a rule with braces. \braceld lower left piece of a horizontal brace. \bracelu upper left piece of a horizontal brace. \bracerd lower right piece of a horizontal brace. \braceru upper right piece of a horizontal brace. \bracevert piece of a vertical brace (|). \brack fraction without a rule with brackets. \break force a line break. \breve math accent: breve (x̆). *\brokenpenalty penalty if page break after

hyphenated line. \buildrel put symbols over math relations. \bullet math operator: bullet (•). \bye finish processing input. \c cedilla accent (ç). \cal math mode calligraphic letters font. \cap math operator: intersection or cap

(∩).

\cases generate a matrix with a left brace delimiter. *\catcode set a character’s processing category

type.

\cdot math operator: centered dot (·). \cdotp centered dot with special spacing in math mode. \cdots three centered dots in math mode

(⋯).

\centering internal Plain TEX space used for centering.

\centerline center text in a line.

*\char specify a character by its numeric

code.

*\chardef define a name for a character. \check math accent: check (x̌). \chi math Greek letter: chi (χ). \choose fraction without a rule with

parentheses.

\circ math operator: circle (∘). *\cleaders surround a space with a repeated box

or rule.

\cleartabs reset tabs.

*\closein close an input file. *\closeout close an output file. *\clubpenalty penalty if page break after first

line of paragraph.

\clubsuit math symbol: club suit (♣). \colon colon in a math formula.

\columns end a \settabs definition.

\cong math relation: congruent (≅). \coprod math operator: large co-product (∐). *\copy use a copy of a box. \copyright copyright symbol (\begin{matrix}c \\ ◯ \end{matrix}). \cos math function: cosine. \cosh math function: hyperbolic cosine. \cot math function: cotangent. \coth math function: hyperbolic cotangent. *\count use a count register. *\countdef define a name for a count register. *\cr end a line in \halign aligned text.

*\crcr  ensure a \cr in \halign.

\csc math function: cosecant. *\csname begin expanding tokens to construct a

command name.

\cup math operator: union or cup (∪). \d dot under accent (\begin{matrix}x \\ . \end{matrix}). \dag dagger symbol (†). \dagger math operator: dagger (†). \dashv math relation: dash V (⊣). *\day current day of the month. \ddag double dagger symbol (‡). \ddagger math operator: double dagger (‡). \ddot math accent: double dot (ẍ). \ddots three diagonal dots in math mode

(...).

*\deadcycles number of output routine calls since

              last \shipout.

*\def define a macro. *\defaulthyphenchar default hyphen when a font

is loaded.

*\defaultskewchar default horizontal kern

character to position accents.

\deg math function: degree of a polynomial (deg).

*\delcode define a character as a delimiter. *\delimiter define a delimiter for math mode. *\delimiterfactor ratio for variable delimiters times 1000.

                                              *\delimitershortfall amount by which

delimiters can fail to span included material.

\delta math Greek letter: delta (δ).

\Delta math Greek letter: capital delta (Δ). \det math function: det. \diamond math operator: diamond (⋄). \diamondsuit math symbol: diamond suit (♢). \dim math function: dim. *\dimen use a dimension register. *\dimendef define a name for a dimension register. *\discretionary append a discretionary item to the current list. *\displayindent indentation of line for displayed

equation. *\displaylimits restore default limit placement

on large math operators.

\displaylines display a stack of formulas

without alignment.

*\displaystyle use normal math script style.

                                              *\displaywidowpenalty penalty for creating a

widow line before a display.

*\displaywidth width of a displayed equation. \div math operator: div. *\divide divide a register by a value. \do redefinable scratch control sequence

                                                             used by \dospecials.

\dospecials change the character code for a set of special characters.

\dosupereject flush insertions and eject to a new page.

\dot math accent: dot (ẋ).

\doteq math relation: dotted equal (≐).

\dotfill fill a space with dots.

\dots an ellipsis, equivalent to \ldots in math mode. *\doublehyphendemerits demerits for consecutive broken lines. \downarrow math symbol: down arrow (↓). \Downarrow math symbol: downward double arrow (⇓). \downbracefill fill a space with a downward

brace.

*\dp depth of a box.

*\dump terminate TEX and write a format

file: INITEX only.

*\edef define a macro with expanded

replacement text.

\egroup end a group.

\eject force a page break. \ell math symbol: small script L (ℓ). *\else alternative to an \if.

3\emergencystretch extra space in

badly-stretched lines.

\empty an empty list of tokens. \emptyset math symbol: empty set (∅). *\end terminate TEX. *\endcsname terminate a \csname token list. \endgraf equivalent to \par. *\endgroup end a group. *\endinput stop reading current input file at the

end of the current line.

\endinsert end of an insert. \endline equivalent to \cr, end of aligned text. *\endlinechar character placed at the right end

of an input line.

\enskip breakable small horizontal skip. \enspace unbreakable small horizontal space. \epsilon math Greek letter: epsilon (ϵ). \eqalign align a stack of equations. \eqalignno align a stack of equations with

equation numbers.

*\eqno equation number. \equiv math relation: equivalence (≡). *\errhelp help message to display if user asks

for help.

*\errmessage write balanced error message to the

terminal.

3\errorcontextlines number of lines of context to be displayed in an error message.

*\errorstopmode pause for normal errors while

processing TEX input.

*\escapechar escape character in the output of

control sequence tokens.

\eta math Greek letter: eta (η). *\everycr tokens to insert after every \cr or

nonredundant \crcr.

*\everydisplay tokens to insert when display

math begins.

*\everyhbox tokens to insert when an hbox

begins.

*\everyjob tokens to insert when the job begins. *\everymath tokens to insert when math in text

begins.

*\everypar tokens to insert when a paragraph

begins.

*\everyvbox tokens to insert when a vbox begins. *\exhyphenpenalty penalty for line break after

explicit hyphen.

\exists math symbol: exists quantifier (∃). \exp math function: exp.

*\expandafter expand the token following the

next token.

*\fam current family number. *\fi end of an \if clause. \filbreak break a page unless there is a better

              \filbreak below.
*\finalhyphendemerits demerits for a

penultimate broken line.

*\firstmark mark text first encountered on page

just boxed.

\fivebf 5 point boldface Roman font. \fivei 5 point math italic font. \fiverm 5 point Roman font. \fivesy 5 point math symbol font. \flat math symbol: flat (♭). *\floatingpenalty penalty for insertions that are

split.

\fmtname name of current TEX format package. \fmtversion version of current TEX format

package.

\folio typeset page number. *\font load a font. *\fontdimen set a font-related parameter. *\fontname gives the external file name for the

given font.

\footins insertion class for footnote inserts. \footline page foot line. \footnote create a footnote. \footnoterule define the rule separating a page

and footnotes.

\footstrut ensure footnote line separation. \forall math symbol: for-all quantifier (∀). \frenchspacing suppress special spacing after

punctuation.

\frown math relation: frown (⌢). *\futurelet assigns the second token to a control

sequence and continues.

\gamma math Greek letter: gamma (γ). \Gamma math Greek letter: capital gamma

(Γ).

\gcd math function: gcd. *\gdef define a macro globally. \ge math relation: greater or equal (≥). \geq math relation: greater or equal (≥). \gets math relation: gets (←). \gg math relation: much greater (≫). *\global the following macro definition or

register setting is global.

*\globaldefs override \global specifications. \goodbreak good page break. \grave math accent: grave (x̀). \H long Hungarian umlaut accent (ő). *\halign create an aligned table.

\hang indent second and subsequent lines in

a paragraph by \parindent.

*\hangafter hanging indentation changes after

specified number of lines.

*\hangindent set hanging indentation. \hat math accent: hat (x̂). *\hbadness limit for bad hbox errors. \hbar math symbol: Planck’s constant or

h-bar (h̄).

*\hbox create a box with horizontal mode

material.

\headline page head line. \heartsuit math symbol: heart suit (♡). *\hfil infinitely stretchable horizontal space. *\hfill more infinitely stretchable horizontal

space.

*\hfilneg cancel infinitely stretchable horizontal

space.

*\hfuzz maximum overrun before overfull

hbox messages occur.

\hglue create some horizontal space. \hideskip special Plain TEX space used in

alignment.

\hidewidth permit an alignment entry to stick

out of its column.

*\hoffset horizontal offset of a page. \hom math function: hom. \hookleftarrow math symbol: hook left arrow

(↩).

\hookrightarrow math symbol: hook right

arrow (↪).

\hphantom create a zero-height hbox with the

width of a formula.

*\hrule produce a horizontal rule. \hrulefill fill a space with a rule. *\hsize width of a paragraph or \hbox. *\hskip add horizontal space. *\hss infinitely stretchable and shrinkable

horizontal space.

*\ht height of a box. *\hyphenation insert words into hyphenation

dictionary.

*\hyphenchar hyphenation character for this font. *\hyphenpenalty penalty for line break after

discretionary hyphen.

\i dotless i letter (ı).

 \ialign  start an \halign with \tabskip

initialized to zero.

*\if compare two character codes. *\ifcase select one of several entries

determined by an integer.

*\ifcat test category codes. *\ifdim test two dimensions.

*\ifeof test for end of file.

\iff math relation: if and only if ( ⟺ ). *\iffalse test always false. *\ifhbox test for a horizontal box. *\ifhmode test for horizontal mode. *\ifinner test for an internal mode. *\ifmmode test for math mode. *\ifnum test two integers. *\ifodd test for odd integer. *\iftrue test always succeeds. *\ifvbox test for a vertical box. *\ifvmode test for vertical mode. *\ifvoid test for an empty box. *\ifx compare tokens. *\ignorespaces read, expand, then ignore tokens

until a non-space is found.

\Im math symbol: imaginary, Fraktur I

(ℑ). \imath math symbol: dotless i (ı).

                                              *\immediate perform a \read or \write

immediately.

\in math relation: in (∈). *\indent start a paragraph indented

                                                              \parindent.

\inf math function: inf.

\infty math symbol: infinity (∞).

*\input read a file.

*\insert define where text will be inserted

when the page is formatted.

*\insertpenalties sum of penalties for split

insertions on the page.

\int math operator: integral (∫).

                                               \interdisplaylinepenalty penalty for breaking

between lines of a display.

                                               \interfootnotelinepenalty penalty for

breaking between lines of a footnote. *\interlinepenalty additional penalty for page

break between lines.

\iota math symbol: iota (ι).

\it select italic font. \item display an item. \itemitem display a sub-item. \itfam italic font family. \j dotless j letter (ȷ). \jmath math symbol: dotless j (ȷ). *\jobname name of the principal input file. \joinrel internal Plain TEX command to piece together long arrows.

\jot unit of measure for opening up displays. \kappa math Greek letter: kappa (κ). \ker math function: ker.

*\kern kern a given distance. \l Polish letter: slashed L (ł). \L Polish letter: upper-case slashed L

(Ł).

\lambda math Greek letter: lambda (λ). \Lambda math Greek letter: capital lambda

(Λ).

\land math operator: logical and (∧). \langle math delimiter: left angle bracket (⟨). 3\language current language used for

hyphenation.

*\lastbox fetch last box off the current list. *\lastkern fetch last kern off the current list. *\lastpenalty fetch last penalty off the current

list.

*\lastskip fetch last skip off the current list. \lbrace math delimiter: left curly brace ({). \lbrack math delimiter: left bracket ([). *\lccode lower-case code for a character. \lceil math delimiter: left ceiling bracket

(⌈).

\ldotp lower dot math symbol with special

spacing.

\ldots three low dots in math mode (…). \le math relation: less or equal (≤). *\leaders fill a space with a repeated box or

rule.

\leavevmode switch to horizontal mode from

vertical mode.

*\left begin a new math list with a left

delimiter.

\leftarrow math symbol: left arrow (←). \Leftarrow math symbol: left double arrow (⇐). \leftarrowfill fill a space with a left arrow. \leftharpoondown math symbol: left harpoon

down (↽).

\leftharpoonup math symbol: left harpoon up

(↼).

3\lefthyphenposition minimum number of

characters before hyphenation at the start of a word.

\leftline flush text left on a line. \leftrightarrow math symbol: left-right arrow

(↔).

\Leftrightarrow math symbol: left-right double

arrow (⇔).

*\leftskip space to the left of a paragraph. \leq math relation: less or equal (≤). \leqalignno align a stack of equations with left

equation numbers.

*\leqno left equation number. *\let define a synonym for the current

meaning of a control sequence.

\lfloor math delimiter: left floor bracket (⌊). \lg math function: lg. \lgroup math delimiter: left group ({). \lhook math symbol: left hook ( ). \lim math function: lim. \liminf math function: limit infimum

(liminf).

*\limits place math limits above and below

math operators.

\limsup math function: limit supremum

(limsup).

\line create a box of current line width. *\linepenalty amount added to badness of every

line in a paragraph.

*\lineskip extra space between lines if

              \baselineskip isn’t enough.

*\lineskiplimit minimum space between lines in

a vertical list.

\ll math relation: much less (≪). \llap create a zero-width box with text to

the left.

\lmoustache math symbol: left moustache ({). \ln math function: ln. \lnot math operator: logical not, hook (¬). \log math function: log. *\long the following macro may have

multiple paragraphs as

parameters.

\longleftarrow math symbol: long left arrow

(⟵).

\Longleftarrow math symbol: long left double

arrow (⟸).

\longleftrightarrow math symbol: long left

and right arrow (⟷).

\Longleftrightarrow math symbol: long

left-right double arrow (⟺).

\longmapsto math symbol: long maps to (⟼). \longrightarrow math symbol: long right arrow

(⟶).

\Longrightarrow math symbol: long right

double arrow (⟹).

\loop begin a loop. *\looseness force a change to the number of lines

in a paragraph.

\lor math symbol: logical or (∨). *\lower lower a box a given distance. *\lowercase translate balanced text to lower-case. \lq math symbol: left quote (‘). *\mag magnification ratio times 1000. \magnification set the magnification for the

document.

\magstep define font magnification.

\magstephalf use 11 point magnification. \makefootline constructs a box with the page

foot.

\makeheadline constructs a box with the page

header.

\mapsto math symbol: maps to (↦). \mapstochar math symbol: maps to char ( ). *\mark place balanced text into a mark item

on the current list.

*\mathaccent place an accent over the next math

field.

*\mathbin define a binary math operator. *\mathchar specify a math character code. *\mathchardef define a name for a math

character.

*\mathchoice choose a math formula based on the

current style.

*\mathclose select math spacing of a closing

delimiter for the next item.

*\mathcode set a math character’s processing

category type.

\mathhexbox internal Plain TEX operation to

define math text symbols.

*\mathinner define an inner math subformula. *\mathop define a large math operator. *\mathopen select math spacing of an opening

delimiter for the next item.

*\mathord define an ordinary math operator. \mathpalette define a symbol that will work

properly in text and script sizes.

*\mathpunct define a math punctuation operator. *\mathrel define a math relation operator. \mathstrut math strut with height and depth of

parentheses.

*\mathsurround define kerning around math in

text.

\matrix generate a matrix without delimiters. \max math function: max. *\maxdeadcycles upper bound on output routine

calls.

*\maxdepth maximum depth of boxes on main

pages.

\maxdimen largest permissible dimension. *\meaning give a token list defining the

semantics of a token.

\medbreak medium vertical space or a good page

break.

*\medmuskip medium space in math formulas. \medskip medium vertical space.

\medskipamount  \medskip space.

*\message write balanced text to terminal. \mid math relation: mid (|). \midinsert insert at current position in page.

\min math function: min. \mit math mode italic font. *\mkern kern in math mode. \models math relation: models (⊧). *\month current month of the year. *\moveleft add a box to the vertical list shifted

left.

*\moveright add a box to the vertical list shifted

right.

\mp math operator: minus plus (∓). *\mskip insert horizontal space in math mode. \mu math Greek letter: mu (μ). *\multiply multiply a register by a value. \multispan span several columns in an

alignment.

*\muskip math skip register. *\muskipdef define a name for a math skip

register.

\nabla math symbol: nabla (∇). \narrower increase left and right margins by

              \parindent.

\natural math symbol: natural (♮). \ne math relation: not equal (≠). \nearrow math symbol: northeast arrow (↗). \neg math operator: negate (¬). \negthinspace unbreakable tiny negative math

mode horizontal space.

\neq math relation: not equal (≠). \newbox allocate a new box register. \newcount allocate a new count register. \newdimen allocate a new dimension register. \newfam define a new font family. \newhelp define a new help message. \newif define a new \if command. \newinsert allocate a new insert register. 3\newlanguage define a new language to be used

for hyphenation.

*\newlinechar character that starts a new output

line in a write statement.

\newmuskip allocate a new math skip register. \newread allocate a new input file. \newskip allocate a new skip register. \newtoks allocate a new token register. \newwrite allocate a new output file. \ni math relation: contains (∋). *\noalign insert unaligned material in \halign

              or \valign.

\nobreak prohibit a line or page break. *\noexpand suppress expansion of a character. *\noindent start a paragraph without

indentation.

\nointerlineskip suppress interline space before next vertical box.

*\nolimits place superscripts and subscripts after

math operators.

\nonfrenchspacing enable special spacing after

punctuation.

*\nonscript suppress space in script and small

script styles.

*\nonstopmode process input without stopping for

errors.

\nopagenumbers suppress page numbers. \normalbaselines set line spacing to normal

values.

\normalbaselineskip normal \baselineskip.

\normalbottom justify page bottoms to the same

height.

\normallineskip value of normal \lineskip.
\normallineskiplimit value of normal
              \lineskiplimit.

\not math operator: not ( ). \not= math relation: not equal (≠). \notin math operator: not in (∉). \nu math Greek letter: nu (ν). \null an empty hbox. *\nulldelimiterspace width of a null delimiter. *\nullfont a font with no characters. *\number convert a number to a token string. \nwarrow math symbol: northwest arrow (↖). \o Norwegian letter: o with slash (ø). \O Norwegian letter: capital O with slash

(Ø).

\obeylines verbatim line breaks in input text. \obeyspaces verbatim spaces in input text. \odot math operator: circle dot (⊙). \oe ligature digraph symbol oe (œ). \OE ligature digraph symbol capital OE

(Œ).

\offinterlineskip turn off extra space between

lines.

\oint math operator: contour integral (∮). \oldstyle select old-style numerals with

descenders.

\omega math Greek letter: omega (ω). \Omega math Greek letter: capital omega (Ω). \ominus math operator: circle minus (⊖). *\omit suppress the template in the

alignment preamble for this entry.

*\openin open a file for input. *\openout open a file for output. \openup increase line separation in math mode. \oplus math operator: direct sum, circle plus

(⊕).

*\or separate cases in an \ifcase. \oslash math operator: circle slash (⊘).

\otimes math operator: tensor product, circle

times (⊗). *\outer the following macro must not be

called from another macro.

*\output define the page output routine. *\outputpenalty penalty at the current page break.

*\over fraction with a rule.

\overbrace horizontal brace over a math formula. *\overfullrule width of rules appended to overfull boxes. \overleftarrow left arrow over a math formula. *\overline draw a line over a formula. \overrightarrow right arrow over a math formula. *\overwithdelims fraction with rule and given delimiters.

\owns math relation: owns (∋). \P math symbol: paragraph (¶). \pagebody put page contents in a box of the proper height. \pagecontents produce a vertical list of the body of a page.

*\pagedepth depth of the current page. *\pagefilllstretch amount of filll space in current page. *\pagefillstretch amount of fill space in current

page.

*\pagefilstretch amount of fil space in current

page.

*\pagegoal desired page height. \pageinsert insert a whole page. \pageno current page number. *\pageshrink amount of glue shrinkage in current

page.

*\pagestretch amount of glue stretch in current

page. *\pagetotal natural height of page so far. *\par end a paragraph. \parallel math relation: parallel (∥). *\parfillskip space at the end of the last line of a paragraph. *\parindent define paragraph indentation. *\parshape define an unusual paragraph shape. *\parskip define space between paragraphs. \partial math symbol: partial (∂). *\patterns define a set of hyphenation patterns. *\pausing pause after each line is read from a

file.

*\penalty specify penalty for a line or page break.

\perp math relation: perpendicular (⊥).

\phantom use the space taken by a formula. \phi math Greek letter: phi (ϕ). \Phi math Greek letter: capital phi (Φ). \pi math Greek letter: pi (π). \Pi math Greek letter: capital pi (Π). \plainoutput default Plain output routine. \pm math operator: plus or minus (±). \pmatrix generate a matrix with parentheses

delimiters.

\pmod math function: mod within

parentheses ( (mod )).

*\postdisplaypenalty penalty for page break

just after a display.

\Pr math function: probability (Pr). \prec math relation: precedes (≺). \preceq math relation: precedes or equal (⪯). *\predisplaypenalty penalty for page break just

before a display.

*\predisplaysize length of text preceding a

display.

\preloaded scratch control sequence used in

preloading fonts.

*\pretolerance badness tolerance before

hyphenation is attempted.

*\prevdepth depth of the last box on the vertical

list.

*\prevgraf number of lines in the last paragraph. \prime math symbol: prime (x′). \proclaim begin theorem with special format in

math environment.

\prod math operator: large product (∏). \propto math relation: proportional to (∝). \psi math Greek letter: psi (ψ). \Psi math Greek letter: capital psi (Ψ). \qquad large breakable horizontal skip. \quad breakable medium horizontal skip. *\radical put a math field under a radical. \raggedbottom suppress bottom justification of

page text.

\raggedright suppress right justification of

paragraph lines.

*\raise raise a box a given distance. \rangle math delimiter: right angle bracket

(⟩).

\rbrace math delimiter: right curly brace (}). \rbrack math delimiter: right bracket (]). \rceil math delimiter: right ceiling bracket

(⌉).

\Re math symbol: real, Fraktur R (ℜ). *\read read a line from a file. *\relax do nothing. \relbar math symbol: relation bar (−). \Relbar math symbol: double relation bar (=).

*\relpenalty penalty for line break after math relation. \removelastskip remove last skip on the list.

                                               \repeat  end of a \loop body.

\rfloor math delimiter: right floor (⌋). \rgroup math delimiter: right group (}). \rho math Greek letter: rho (ρ). \rhook math symbol: right hook ( ). *\right end a math list with a right delimiter. \rightarrow math symbol: right arrow (→). \Rightarrow math symbol: right double arrow

(⇒).

\rightarrowfill fill a space with a right arrow. \rightharpoondown math symbol: right harpoon

down (⇁). \rightharpoonup math symbol: right harpoon

up (⇀).

                                              3\righthyphenposition minimum number of

characters after hyphenation at the end of a word.

\rightleftharpoons math symbol: right left harpoon (⇀ ↽).

\rightline flush right text on a line.

*\rightskip space to the right of a paragraph. \rlap create a zero-width box with text to

the right.

\rm select Roman font.

\rmoustache math symbol: right moustache (}).

*\romannumeral convert a number to lower-case Roman numerals.

\root specified root of a formula.

\rq right quote (’).

\S math symbol: section (§). \sb subscript, equivalent to _. *\scriptfont select font for small math scripts. *\scriptscriptfont select font for very small math scripts. *\scriptscriptstyle use very small math script

style. *\scriptspace extra space after subscript or superscript. *\scriptstyle use small math script style. *\scrollmode process TEX input without pausing for normal errors. \searrow math symbol: southeast arrow (↘). \sec math function: secant. *\setbox store an hbox or vbox in a box register. 3\setlanguage specify language to be used for hyphenation. \setminus math operator: set minus (\). \settabs define horizontal tabs.

\sevenbf 7 point bold Roman font. \seveni 7 point math italic font. \sevenrm 7 point Roman font. \sevensy 7 point math symbol font. *\sfcode set a character’s space factor. \sharp math symbol: sharp (♯). *\shipout send a box to the DVI file. *\show show a token. *\showbox show a box. *\showboxbreadth maximum boxed items shown

at a given depth.

*\showboxdepth maximum box depth shown. \showhyphens show hyphenations of given words. *\showlists show the current lists. *\showthe display the contents of a register. \sigma math Greek letter: sigma (σ). \Sigma math Greek letter: capital sigma (Σ). \sim math relation: similar (∼). \simeq math relation: similar or equal (≃). \sin math function: sine. \sinh math function: hyperbolic sine. \skew shift super accents. *\skewchar horizontal kern to position accents. *\skip use a skip register. *\skipdef define a name for a skip register. \sl select slanted font. \slash permit a line break after a slash. \slfam slanted font family. \smallbreak small vertical space and a good

page break.

\smallint math operator: small integral (∫). \smallskip small vertical space.

\smallskipamount  \smallskip space.

\smash create a formula with zero height. \smile math relation: smile (⌣). \sp superscript, equivalent to ^. \space a blank space. *\spacefactor set a character’s spacing factor. *\spaceskip space between words. \spadesuit math symbol: spade suit (♠). *\span create a multicolumn entry in an

aligned table.

*\special include balanced text in DVI file for

post-processing.

*\splitbotmark mark text last encountered in a

split box.

*\splitfirstmark mark text first encountered in

a split box.

*\splitmaxdepth maximum depth of boxes in a

split box.

*\splittopskip space at top of a split box. \sqcap math operator: square cap (⊓). \sqcup math operator: square cup (⊔).

\sqrt square root of a formula. \sqsubseteq math relation: square subset or

equal (⊑).

\sqsupseteq math relation: square superset or

equal (⊒).

\ss German letter: sharp s (ß). \star math operator: star (⋆). *\string expand a control sequence into

character tokens.

\strut vertical strut to preserve line spacing. \strutbox box containing a strut. \subset math relation: subset (⊂). \subseteq math relation: subset or equal (⊆). \succ math relation: successor (≻). \succeq math relation: successor or equal (⪰). \sum math operator: large sum (∑). \sup math function: sup. \supereject flush all insertions and eject to a

new page.

\supset math relation: superset (⊃). \supseteq math relation: superset or equal (⊇). \surd math symbol: surd (√). \swarrow math symbol: southwest arrow (↙). \t tie-after accent (oo). \tabalign begin a tabbed line in an inner

environment.

\tabs box containing current tabs. *\tabskip space between aligned tab entries. \tan math function: tangent. \tanh math function: hyperbolic tangent. \tau math Greek letter: tau (τ). \tenbf 10 point bold font. \tenex 10 point math extension symbol font. \teni 10 point math italic font. \tenit 10 point text italic font. \tenrm 10 point Roman font. \tensl 10 point slanted font. \tensy 10 point math symbol font. \tentt 10 point typewriter font. \TeX “TEX” logo. *\textfont select font for non-math text. \textindent display an item without hanging

indentation.

*\textstyle use in-text style for math. *\the convert a numeric register to

displayable form.

\theta math Greek letter: theta (θ). \Theta math Greek letter: capital theta (Θ). *\thickmuskip thick space in math formulas. *\thinmuskip thin space in math formulas. \thinspace unbreakable small horizontal space. \tilde math accent: tilde (x̃). *\time current time of day.

\times math operator: times (×). \to math relation: to (→). *\toks use a token list register. *\toksdef define a name for a token list register. *\tolerance badness tolerance after hyphenation. \top math symbol: top (⊤). \topins insertion class for inserts at the top of

a page.

\topinsert insert text at the top of the page. *\topmark value of \botmark just before current

page was boxed.

*\topskip space at the top of a page. \tracingall turn on all debugging commands. *\tracingcommands show commands before they

are executed.

*\tracinglostchars show characters not in the

font.

*\tracingmacros show macros as they are

expanded.

*\tracingonline show diagnostics on the

terminal.

*\tracingoutput show boxes that are shipped

out.

*\tracingpages show page-break calculations.

*\tracingparagraphs show line-break

calculations.

*\tracingrestores show unassignments when

groups end.

*\tracingstats show statistics about memory

usage.

\triangle math symbol: triangle (△). \triangleleft math operator: left triangle (◁). \triangleright math operator: right triangle

(▷).

\tt select typewriter font. \ttfam typewriter font family. \ttraggedright suppress right justification of

typewriter font lines.

\u breve accent (x̆). *\uccode upper-case code for a character. *\uchyph positive if hyphenating words

beginning with capital letters.

\undefined an undefined control sequence. \underbar create a box with an underline. \underbrace horizontal brace under a math

formula.

*\underline draw a line under a formula. *\unhbox un-box an hbox and add it to the

horizontal list.

*\unhcopy un-box a copy of an hbox and add it

to the horizontal list.

*\unkern remove a kern just added to the current list.

*\unpenalty remove a penalty just added to the

current list.

*\unskip remove a skip just added to the

current list.

*\unvbox un-box a vbox and add it to the

vertical list.

*\unvcopy un-box a copy of a vbox and add it to

the vertical list.

\uparrow math delimiter: upward arrow (↑). \Uparrow math symbol: upward double arrow

(⇑).

\upbracefill fill a space with an upward brace. \updownarrow math delimiter: up-and-down

arrow (↕).

\Updownarrow math symbol: up-and-down

double arrow (⇕).

\uplus math operator: U plus (⊎). *\uppercase translate balanced text to

upper-case.

\upsilon math Greek letter: upsilon (υ). \Upsilon math Greek letter: capital upsilon

(Υ).

\v check accent (x̌). *\vadjust insert vertical material into a

paragraph.

*\valign align zero or more columns. \varepsilon math Greek letter: variant epsilon

(ε).

\varphi math Greek letter: variant phi (φ). \varpi math Greek letter: variant pi (ϖ). \varrho math Greek letter: variant rho (ϱ). \varsigma math Greek letter: variant sigma (ς). \vartheta math Greek letter: variant theta (ϑ). *\vbadness limit for bad vbox errors. *\vbox create a box with vertical mode

material.

*\vcenter create a centered vbox in a math list. \vdash math relation: dash (⊢). \vdots math symbol: vertical dots (⋮). \vec math accent: vector (x⃗). \vee math binary operator: join or V (∨). \vert math delimiter: vertical bar (|). \Vert math delimiter: double vertical bar

(∥).

*\vfil infinitely stretchable vertical space. *\vfill more infinitely stretchable vertical

space.

*\vfilneg cancel infinitely stretchable vertical

space.

\vfootnote place a footnote in a caption or other

vertical list.

*\vfuzz maximum overrun before overfull vbox error.

\vglue create some vertical space. *\voffset vertical offset of a page. \vphantom create a zero-width vbox the height

and depth of a formula.

*\vrule produce a vertical rule. *\vsize height of text on a page or \vbox. *\vskip add vertical space. *\vsplit split off a specified amount from a

vbox.

*\vss infinitely stretchable and shrinkable

vertical space.

*\vtop create a box with vertical mode material with the baseline at the top.

*\wd width of a box. \wedge math binary operator: meet or wedge

(∧).

\widehat math accent: wide hat (x̂). \widetilde math accent: wide tilde (x̃). *\widowpenalty penalty for creating a widow line

at top of page.

\wlog write to the log file only. \wp math symbol: Weierstrass p (℘). \wr math binary operator: wreath

product (≀).

*\write write a token list to a file. *\xdef define a global macro with expanded

replacement text.

\xi math Greek letter: xi (ξ). \Xi math Greek letter: capital xi (Ξ). *\xleaders fill a space with an evenly distributed

box or rule.

*\xspaceskip space between sentences. *\year current year of our Lord. \zeta math Greek letter: zeta (ζ).

Number of TEX primitives and commands: 864.

L^{A}TEX Commands

\' put text flush right in a column in

tabbing environment.

\( begin math mode. \) end math mode. \- unindents left margin one tab stop in

tabbing environment.

\: medium space in math mode. \< cancel effect of one \+ command in

tabbing environment.

\= define a tab stop in tabbing

environment.

\> move to next tab position in tabbing

environment.

\@ create end-of-sentence space after

following punctuation.

\[ begin display math mode. \\ force a line break in a paragraph. \] end display math mode. \+ indents left margin one tab stop in

tabbing environment.

\@@par TEX primitive \par. \@beginparpenalty penalty at the beginning of

a list.

\@endparpenalty penalty at the end of a list. \@hangfrom hanging indent of a paragraph. \@ifnextchar test the next character. \@itempenalty penalty between list items. \a produce TEX accents in tabbing

environment.

\addcontentsline adds an entry to the specified

list or table.

\addtocontents add text to table contents,

figures, or tables.

\addtocounter increment a counter. \addtolength increment the value of a length

command.

\addvspace add extra vertical space. \alph display counter as lower-case letter. \Alph display counter as upper-case letter.

\and additional authors on title page. \appendix set sectional units to appendix style. \arabic display counter as Arabic numerals. \array begin array environment. \arraycolsep column separation in array

environment.

\arrayrulewidth set width of array rules. \arraystretch space between rows of array or

tabular environment.

\author title page author.

\baselinestretch multiple of normal
              \baselineskip.

\begin beginning of an environment. \bf bold font style. \bibitem create a bibliography entry. \bibliography enable BibTEX bibliography

interface.

\bibliographystyle specify bibliographic style

for BibTEX.

\big big font size. \Big bigger font size. \BIG biggest font size. \bigskip vertical skip a large amount. \boldmath select bold math italic and symbol

fonts.

\botfigrule place rule between text and bottom

floats.

\bottomfraction fraction of column for bottom

floats.

\Box math symbol: box (□). \caption produce a numbered caption. \center begin centering environment. \centering declaration to center lines. \chapter start a chapter. \circle circle in a picture environment. \cite generate an in-text citation of a

reference.

\cleardoublepage flush figures and tables and start a new right-hand page.

\clearpage flush figures and start a new page. \cline multicolumn line in array or tabular

environment.

\columnsep space between columns in double

column text.

\columnseprule width of rule separating double

columns.

\columnwidth width of a column. \dashbox create a box in dashes, with

positioning.

\date date on title page. \dblfigrule rule between double-column floats

and text.

\dblfloatpagefraction size of float on

double-column page.

\dblfloatsep space between double-column

floats.

\dbltextfloatsep space between double-column

floats and text.

\dbltopfraction fraction of two-column page for

top floats.

\Diamond math symbol: diamond (◇). \document read in the .AUX files and disable

alpha @ processing.

\documentstyle declare document style and

options.

\doublerulesep double rule separation in array

or tabular environment.

\egtit 8 point italic font. \egtly 8 point L^{A}TEX symbol font. \egtmi 8 point math italic font. \egtrm 8 point Roman font. \egtsy 8 point math symbol font. \elvbf 11 point bold extended font. \elvit 11 point italic font. \elvly 11 point L^{A}TEX symbol font. \elvmi 11 point math italic font. \elvrm 11 point Roman font. \elvsf 11 point sans serif font. \elvsl 11 point slanted font. \elvsy 11 point math symbol font. \elvtt 11 point typewriter font. \em emphasis font. \end end of an environment. \endarray end array environment. \endcenter end centering environment. \endequation end an equation. \endfigure end a floating figure. \endflushleft end flush left environment. \endflushright end flush right environment. \endpicture end picture environment.

\endthebibliography end bibliography

environment.

\equation begin and cross reference an equation. \evensidemargin left hand margin on even pages.

\extracolsep add extra space before a column in

array or tabular environment.

\fbox create and frame a box.

\fboxrule thickness of rule surrounding framed

box.

\fboxsep space between a box and its contents

                                                             in \fbox and \framebox.

\figure begin a floating figure.

\fill an unoriented skip amount with

stretch of 1fill.

\fivly 5 point L^{A}TEX symbol font. \fivmi 5 point math italic font. \fivrm 5 point Roman font. \fivsy 5 point math symbol font. \floatpagefraction portion of page that may

be occupied by floats.

\floatsep space between floats.

\flushbottom justify page bottoms to the same

height. \flushleft begin flush left environment. \flushright begin flush right environment. \fnsymbol display counter as footnote symbol. \footheight height of page footer. \footins space between main text and footnotes. \footnotemark insert a footnote mark without

text. \footnoterule macro to draw the rule

separating footnotes from text.

\footnotesep footnote separation distance. \footnotesize select footnote font size. \footnotetext produce footnote text without a

mark.

\footskip space between text and page footer. \frac generate a fraction. \frame put a frame around some text. \framebox create and frame a box, with positioning. \frtnbf 14 point bold extended font. \frtnly 14 point L^{A}TEX symbol font. \frtnmi 14 point math italic font. \frtnrm 14 point Roman font. \frtnsy 14 point math symbol font. \fussy require strict line and page breaks. \glossary write a glossary entry. \headheight height of page header. \headsep space between page header and text. \hline horizontal line in array and tabular environments.

\hspace skip horizontal space.

\huge select font four steps larger than

normal size.

\Huge select largest available font. \include read a file unless disabled by

\includeonly.
\includeonly enable only specific \include

commands.

\index create an index entry. \indexentry en entry in a file of index terms. \indexspace amount of extra space between

entries in an index.

\interfootnotelinepenalty interline penalty

for footnotes.

\intextsep vertical space around a float in the

middle of a page.

\item start an entry in a list environment. \itemindent indent before the label in a list

environment.

\itemsep space between successive items in a

list environment.

\Join math symbol: join (⋈). \kill discard current line in tabbing

environment.

\label define a cross reference label. \labelitemi mark denoting first-level items. \labelitemii mark denoting second-level items. \labelitemiii mark denoting third-level items. \labelitemiv mark denoting fourth-level items. \labelitemv mark denoting fifth-level items. \labelitemvi mark denoting sixth-level items. \labelsep space between a label and text of a

list item.

\labelwidth label width in a list environment. \large select font one step larger than

normal size.

\Large select font two steps larger than

normal size.

\LARGE select font three steps larger than

normal size.

\LaTeX L^{A}TEX logo. \leadsto math symbol: leads to (⇝). \lefteqn set a formula flush left in display style. \leftmargin left margin of a list environment. \leftmargini width of left margin in first-level

list.

\leftmarginii width of left margin in

second-level list.

\leftmarginiii width of left margin in

third-level list.

\leftmarginiv width of left margin in

fourth-level list.

\leftmarginv width of left margin in fifth-level list.

\leftmarginvi width of left margin in sixth-level

list.

\leftmark produces current left page heading. \lhd math symbol: left half diamond (⊲). \line line in a picture environment. \linebreak encourage a line break. \linethickness set width of lines in picture

environment.

\linewidth width of the current line. \listoffigures generate a list of figures. \listoftables generate a list of tables. \listparindent indent second and subsequent

paragraphs in a list environment.

\load load a font. \makeatletter set @ alphabetic to access internal

commands.

\makeatother set @ non-alphabetic to hide

internal commands.

\makebox create a box, with positioning. \makeglossary enable glossary processing. \makeindex enable index creation. \makelabel create item label for a list

environment.

\maketitle produce the title. \marginpar produce a marginal note. \marginparpush vertical space between marginal

notes.

\marginparsep distance between marginal note

and text.

\marginparwidth width of marginal notes. \markboth set headings for left and right pages. \markright set heading for right pages. \mathindent indentation of display equations. \mbox create a box. \medskip vertical skip a medium amount. \mho math symbol: mho (℧). \minipage create a small sample page. \multicolumn multicolumn entry in array or

tabular environment.

\multiput place several copies of a picture

object.

\newcommand define a new command. \newcounter define a new counter variable. \newenvironment define a new environment. \newfont select a font. \newlength define a new length command. \newline force a line break in a paragraph. \newpage start a new page. \newsavebox define a new box register. \newtheorem define a new theorem environment. \ninbf 9 point bold extended font. \ninit 9 point italic font. \ninly 9 point L^{A}TEX symbol font.

\ninmi 9 point math italic font. \ninrm 9 point Roman font. \ninsy 9 point math symbol font. \nintt 9 point typewriter font. \nocite reference a BibTEX item without

citation.

\nofiles suppress writing all auxiliary files. \nolinebreak suppress a line break. \nonumber suppress numbering of displayed

equations.

\nopagebreak suppress a page break. \normalmarginpar put marginal notes on default

side of page.

\normalsize select normal font size. \numberline includes the section number in a

\contentsline command.

\obeycr defines a carriage return as \\. \oddsidemargin left hand margin on odd pages. \onecolumn start a new page in single-column

format.

\oval oval in a picture environment. \pagebreak encourage a page break. \pagenumbering specify page numbering style. \pageref page number of a cross reference label. \pagestyle set global page style. \paragraph start a labeled paragraph. \parbox put a paragraph in a box. \parsep space between paragraphs within an

item in a list environment.

\part start a major division of a long

document.

\partopsep extra vertical space when

environment starts a paragraph.

\picture begin picture environment. \poptabs restore tabs stops in tabbing

environment.

\pounds British pound symbol (£). \protect protect fragile commands and moving

arguments.

\pushtabs save current tab stops in tabbing

environment.

\put place a picture object. \raggedbottom justify page bottoms to their

natural height.

\raggedleft flush lines right. \raggedright flush lines left. \raisebox raise a box a distance. \ref refer to a cross reference label. \refstepcounter increment and reference a

counter.

\renewcommand redefine a command. \renewenvironment redefine an environment.

\restorecr restores a carriage return to its usual

meaning.

\reversemarginpar put marginal notes on

opposite side of page.

\rhd math symbol: right half diamond (⊳). \rightmargin right margin in a list. \rightmark produces current right page heading. \roman display counter as lower-case Roman

numerals.

\Roman display counter as upper-case Roman

numerals.

\rule create a rule or line. \samepage inhibit a page break over a region. \savebox create and name a box, with

positioning.

\sbox create and name a box. \sc select small caps font. \scriptscriptsize select sub-subscript font size. \scriptsize select subscript or superscript font

size.

\section start a section. \setcounter set value of a counter. \setlength set a length command. \settowidth set a length command to width of

text.

\sevit 7 point italic font. \sevly 7 point L^{A}TEX symbol font. \sevmi 7 point math italic font. \sevrm 7 point Roman font. \sevsy 7 point math symbol font. \sf select sans serif font. \shortstack create a box with a single column of

items.

\showoutput trace output processing. \sixly 6 point L^{A}TEX symbol font. \sixmi 6 point math italic font. \sixrm 6 point Roman font. \sixsy 6 point math symbol font. \sloppy accept looser line and page breaks. \small select small font size. \smallskip vertical skip a small amount. \sqsubset math symbol: square subset (⊏). \sqsupset math symbol: square superset (⊐). \stackrel stack one equation above another. \stepcounter increment a counter. \stop terminate L^{A}TEX and flush the final

page.

\stretch infinitely stretchable space. \subitem second-level item in an index. \subparagraph start a labeled sub-level

paragraph.

\subsection start a subsection. \subsubitem third-level item in an index.

\subsubsection start a subsubsection. \svtnbf 17 point bold extended font. \svtnly 17 point L^{A}TEX symbol font. \svtnmi 17 point math italic font. \svtnrm 17 point Roman font. \svtnsy 17 point math symbol font. \symbol display a symbol from a font. \tabbingsep distance left by \' command

between tabbing fields.

\tabcolsep half the width separating columns in

a tabular environment.

\tableofcontents generate a table of contents. \tencirc 10 point circle font. \tencircw 10 point wide circle font. \tenln 10 point line font. \tenlnw 10 point wide line font. \tenly 10 point L^{A}TEX symbol font. \tenmi 10 point math italic font. \tensf 10 point sans serif font. \textfloatsep space between floats and the text. \textfraction size of column that must contain

text.

\textheight height of text, excluding head and

foot.

\textwidth width of printing on page. \thanks add footnote to title page. \theenumi current value of the first-level item

counter.

\theenumii current value of the second-level item

counter.

\theenumiii current value of the third-level item

counter.

\theenumiv current value of the fourth-level item

counter.

\theenumv current value of the fifth-level item

counter.

\theenumvi current value of the sixth-level item

counter.

\theequation number of the current displayed

equation.

\thefigure current value of the figure counter. \thefootnote produces the footnote number. \thepage current page number. \thicklines thick lines for lines and circles. \thinlines thin lines for lines and circles. \thispagestyle set current page style. \tiny select smallest defined font. \title declare the title. \today display today’s date. \topfigrule place rule between top floats and

text.

\topfraction fraction of column for top floats. \topmargin extra space added to top of page.

\topsep space between first list item and

preceding paragraph.

\twlbf 12 point bold extended font.

\twlit 12 point italic font.

\twlly 12 point L^{A}TEX symbol font. \twlmi 12 point math italic font. \twlrm 12 point Roman font. \twlsf 12 point sans serif font. \twlsl 12 point slanted font. \twlsy 12 point math symbol font. \twltt 12 point typewriter font. \twocolumn start a new page in double-column

format. \twtybf 20 point bold extended font. \twtyly 20 point L^{A}TEX symbol font. \twtymi 20 point math italic font. \twtyrm 20 point Roman font. \twtysy 20 point math symbol font. \typein execute a command from the terminal. \typeout write a message on the terminal. \unboldmath unselect bold math italic and symbol fonts. \underline create an underlined box with text in

it. \unitlength unit of distance in picture

environment. \unlhd math symbol: underlined left half diamond (⊴). \unrhd math symbol: underlined right half diamond (⊵).

\usebox use a named box.

\usecounter associate a counter with an

item-type command.

\value produce the value of a counter. \vector vector in a picture environment. \verb generate a short amount of verbatim

text.

\vline vertical line in array and tabular environments.

\vspace skip vertical space.

Number of L^{A}TEX commands: 367.